### v1

# Reflection Theory of Nichols Algebras over Coquasi-Hopf Algebras with Bijective Antipode

Abstract
We investigate the reflection theory of Nichols algebras over arbitrary coquasi-Hopf algebras with bijective antipode, generalizing previous results restricted to the pointed cosemisimple setting Li and Liu (2025b). By establishing a braided monoidal equivalence between categories of rational Yetter-Drinfeld modules via a dual pair, we demonstrate that a tuple of finite-dimensional irreducible Yetter-Drinfeld modules admitting all reflections gives rise to a semi-Cartan graph. As an application, we consider an explicit example of a rank three Nichols algebra from Huang et al. (2026). We show that it yields a standard Cartan graph and prove that it is, in fact, an affine Nichols algebra.
1 Introduction
1.1 The classical setting: Nichols algebras over Hopf algebras
Nichols algebras play an important role in the theory of Hopf algebras and quantum groups, serving as key building blocks for the classification of pointed Hopf algebras and the structural analysis of braided tensor categories. Their combinatorial and representation-theoretic properties are intrinsically linked to the underlying Yetter–Drinfeld modules, and understanding their behavior under natural maps, such as reflections, has become a central theme in the research of Nichols algebra.
Heckenberger introduced the concepts of semi-Cartan graphs and Weyl groupoids [Heckenberger, 2006,Heckenberger, 2009,Heckenberger, 2010], which provide powerful combinatorial tools for studying the structure and classification of Nichols algebras. A semi-Cartan graph is essentially a quadruple $\mathcal{G}=\mathcal{G}(\mathbb{I},\mathcal{X},r,A)$ consisting of a finite index set $\mathbb{I}$, a set of points $\mathcal{X}$, a set of reflection maps $r$, and a generalized Cartan matrix $A^{X}$ associated with each point, satisfying specific compatibility axioms (see Definition 5.14 for details) The most celebrated theorem states that a tuple of simple Yetter-Drinfeld modules over a Hopf algebra with bijective antipode gives rise to a semi-Cartan graph, provided that it admits all reflections [Andruskiewitsch et al., 2010,Heckenberger and Schneider, 2010b,Heckenberger and Schneider, 2013a].
For Hopf algebras, the reflection theory of Nichols algebras, has been extensively developed and successfully applied to classify finite-dimensional Nichols algebras of diagonal type [Heckenberger, 2006,Heckenberger, 2010,Angiono, 2013,Angiono, 2015]. Moreover, the classification of finite-dimensional pointed Hopf algebras over abelian groups has been completed with the aid of the classification of Nichols algebras of diagonal type [Andruskiewitsch and Schneider, 1998,Andruskiewitsch and Schneider, 2000,Andruskiewitsch and Schneider, 2010,Andruskiewitsch et al., 2014,Angiono and García Iglesias, 2019,Andruskiewitsch and Angiono, 2017]. Significant progress has also been made in classifying finite-dimensional Nichols algebras over non-abelian groups [Andruskiewitsch and Graña, 1999,Andruskiewitsch and Graña, 2003,Andruskiewitsch et al., 2010,Milinski and Schneider, 2000,Fomin and Kirillov, 1999,Graña, 2000,Graña et al., 2011,Dello and Zhang, 2016,Heckenberger and Schneider, 2010b,Heckenberger and Schneider, 2013b,Heckenberger and Schneider, 2010a,Heckenberger and Vendramin, 2014,Heckenberger and Vendramin, 2015,Heckenberger and Vendramin, 2017].
1.2 From Hopf algebras to coquasi‑Hopf algebras
In this work, we focus on the more general setting of coquasi-Hopf algebras. The classification of coradically graded pointed finite-dimensional coquasi-Hopf algebras over abelian groups has been completed in [Huang et al., 2011,Huang and Liu, 2012,Huang et al., 2015,Huang et al., 2021,Huang et al., 2020,Huang et al., 2026,Li and Liu, 2025a]. While the theory of Nichols algebras over Hopf algebras is mature, the parallel theory for coquasi-Hopf algebras remains less developed due to the lack of associativity. Recent progress has been made for pointed cosemisimple coquasi-Hopf algebras under mild conditions Li and Liu (2025b), which requires the finite‑dimensionality of certain Nichols algebras appearing in reflection sequences. This ensures the following braided monoidal equivalence, which is crucial to the main theorem:
${{}^{\mathcal{B}(V)}_{\mathcal{B}(V)}\mathcal{YD}(\mathcal{C})}\longrightarrow{{}^{\mathcal{B}(V^{*})}_{\mathcal{B}(V^{*})}\mathcal{YD}(\mathcal{C})}.$
Here $\mathcal{C}={{}^{G}_{G}\mathcal{YD}^{\Phi}}$ and $\mathcal{B}(V)$ is a finite-dimensional Nichols algebra in $\mathcal{C}$.
The primary obstacle in extending the theory to the arbitrary setting lies in the duality. We now consider two arbitrary Hopf algebras $A$, $B$ in $\mathcal{C}$ with bijective antipodes related by a non-degenerate Hopf pairing. Unfortunately, the monoidal categories $^{A}_{A}\mathcal{YD}(\mathcal{C})$} and $^{B}_{B}\mathcal{YD}(\mathcal{C})$} are not monoidal equivalence in general. More precisely, when $A$, $B$ are infinite-dimensional, a module of $A$ in $\mathcal{C}$ can not give rise to a comodule of $B$ in $\mathcal{C}$.
1.3 Braided monoidal equivalence induced by a dual pair and its application to reflection theory
In the present work, we overcome several new difficulties that do not appear in pointed cosemisimple settings and remove the assumption on finite-dimensionality. As a result, we extend the reflection theory to arbitrary coquasi‑Hopf algebras $H$ with bijective antipode.
Inspired by methods in [Heckenberger and Schneider, 2013a,Heckenberger and Schneider, [2020] ©2020], we can define rational modules of $A$ in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$. This allows us to define the category ${{}^{R}_{R}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$, which is proven to be a monoidal subcategory of $^{R}_{R}\mathcal{YD}(\mathcal{C})$}. Now for two locally finite Hopf algebras $A$, $B$ in $\mathcal{C}$ with bijective antipodes related by a non-degenerate Hopf pairing compatible with the grading, called a dual pair. We prove the following result:
Theorem 1.1 (Theorem 3.16).
There is a braided monoidal equivalence:
$\Omega:{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}\longrightarrow{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}.$
Furthermore, we show that if $V$ is a finite-dimensional Yetter-Drinfeld module in $^{H}_{H}\mathcal{YD}$}, there is a dual pair between $\mathcal{B}(V)$ and $\mathcal{B}(V^{*})$. Thus the above theorem can be applied to Nichols algebras to get the following braided monoidal equivalence:
$\Omega_{V}:{{}^{\mathcal{B}(V)}_{\mathcal{B}(V)}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}\longrightarrow{{}^{\mathcal{B}(V^{*})}_{\mathcal{B}(V^{*})}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}.$
Now we are going to state our main result. Let $\theta\geq 1$ be a positive integer, $\mathbb{I}=\{1,2,..,\theta\}$ and $M=(M_{1},\ldots,M_{\theta})$, where $M_{1},\ldots,M_{\theta}\in{{}^{H}_{H}\mathcal{YD}}$ are finite-dimensional irreducible Yetter-Drinfeld modules. We say $M$ admits $i$-th reflection for some $1\leq i\leq\theta$ if for all $j\neq i$ there is a natural number $m_{ij}^{M}\geq 0$ such that $(\mathrm{ad}\,M_{i})^{m_{ij}^{M}}(M_{j})$ is a non-zero finite-dimensional subspace of $\mathcal{B}(M)$, and $(\mathrm{ad}\,M_{i})^{m_{ij}^{M}+1}(M_{j})=0$. Assume $M$ admits the $i$-th reflection. Then we set $R_{i}(M)=(V_{1},\ldots,V_{\theta})$, where
$V_{j}=\begin{cases}M_{i}^{*},&\text{if }j=i,\\ (\mathrm{ad}\,M_{i})^{m_{ij}^{M}}(M_{j}),&\text{if }j\neq i.\end{cases}$
It is still a tuple of finite-dimensional irreducible Yetter-Drinfeld modules. By above theorem, we have a braided monoidal equivalence:
$\Omega_{i}:{{}^{\mathcal{B}(M_{i})}_{\mathcal{B}(M_{i})}\mathcal{YD}(\mathcal{\mathcal{C}})}_{\operatorname{rat}}\longrightarrow{{}^{\mathcal{B}(M_{i}^{*})}_{\mathcal{B}(M_{i}^{*})}\mathcal{YD}(\mathcal{\mathcal{C}})}_{\operatorname{rat}}.$
Then we prove the following theorem.
Theorem 1.2.
(Theorem 5.12) With above assumptions on $M$. Suppose $M$ admits the $i$-th reflection for $1\leq i\leq\theta$. Then there is an isomorphism of Hopf algebras in $^{H}_{H}\mathcal{YD}$}:
$\Theta_{i}:\mathcal{B}(R_{i}(M))\cong\Omega_{i}\left(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\right)\#\mathcal{B}(M_{i}^{*}).$ (1.1)
Although our main conclusion appears consistent with the Hopf case, the specific details differ significantly. Some details require adjustment or reformulation in this more general setting. Some identities that are obviously valid in Hopf algebras require extensive computations to verify in our context.
As a corollary, suppose $M$ admits all reflections and let $\mathcal{X}=\{[X]\mid X\in\mathcal{F}_{\theta}(M)\}$, see Definition 5.13 for related notations. We denote the generalized Cartan matrix $A^{[X]}=(a_{ij}^{X})_{i,j\in\mathbb{I}}$ for each $X]\in\mathcal{X}$, where $(a_{ij}^{X})_{i,j\in\mathbb{I}}$ is defined in Lemma 5.9. Let $r$ be the following map,
$r:\mathbb{I}\times\mathcal{X}\rightarrow\mathcal{X},\ i\times[X]\mapsto[R_{i}(X)].$
Then the quadruple
$\mathcal{G}(M)=(\mathbb{I},\mathcal{X},r,(A^{X})_{X\in\mathcal{X}})$
is a semi-Cartan graph.
Finally, we apply this machinery to a crucial test case: a rank-three Nichols algebra $\mathcal{B}(M)=\mathcal{B}(M_{1}\oplus M_{2}\oplus M_{3})$ of non-diagonal type, originally identified in [Huang et al., 2026, Proposition 4.1], which is proved to be infinite-dimensional via three different approachs [Huang et al., 2026,Li and Liu, 2025a,Li and Liu, 2025b]. We compute the Weyl groupoid of $\mathcal{W}(\mathcal{G}(M))$, and show that the set of real roots of $\mathcal{G}(M)$ at each $X\in\mathcal{X}$ is just the set of real roots of Kac-Moody Lie algebra of $\mathcal{A}_{2}^{(1)}$. Furthermore, we prove that $\mathcal{G}(M)$ is a Cartan graph. Finally, utilizing the correspondence between Cartan graphs and Tits cones[Cuntz, 2018,Cuntz et al., 2019], we demonstrate that this algebra is affine, meaning its Tits cone is a half-plane. Our observations shows that affine Nichols algebras can be realized over coquasi-Hopf algebras.
1.4 Organization of this paper
The paper is organized as follows. After reviewing preliminaries on coquasi-Hopf algebras, Yetter–Drinfeld modules, and bosonization in Section 2, we introduce rational modules and establish equivalences of braided monoidal categories via dual pair in Section 3. Section 4 specializes to Nichols algebras and their pairings. Section 5 is devoted to the general reflection theory: we study the projection of a Nichols algebra and reflections of Nichols algebras. Finally, in Section 6 we work out a detailed example of an affine Nichols algebra and showing that it gives rise to a Cartan graph.
2 Preliminaries
In this section, we briefly recall the fundamental definitions and fix the notation for coquasi-Hopf algebras and Yetter-Drinfeld modules. For a more detailed exposition and proofs of standard properties, we refer the reader to Li and Liu (2025b). Throughout the paper, $\mathbbm{k}$ denotes an algebraically closed field of characteristic zero.
2.1 Coquasi-Hopf algebras
We follow the standard definition of a coquasi-Hopf algebra as introduced in Drinfeld (1989). A coquasi-Hopf algebra is a coalgebra $(H,\Delta,\epsilon)$ equipped with a compatible quasi-algebra structure $(m,\mu,\Phi)$ and a quasi-antipode $(\mathcal{S},\alpha,\beta)$. The associator $\Phi$ is a convolution-invertible map satisfying the 3-cocycle condition, while $\alpha,\beta$ enforce the quasi-antipode axioms. We omit the full list of axioms here; see [Li and Liu, 2025b, Definition 2.1] for the complete list of identities.
Throughout this paper, we use the Sweedler notation $\Delta(a)=a_{1}\otimes a_{2}$ for the coproduct and $a_{1}\otimes a_{2}\otimes\cdots\otimes a_{n+1}$ for the result of the $n$-iterated application of $\Delta$ on $a$. We say $H$ has a bijective antipode if $\mathcal{S}$ is bijective.
Following the definition in [Ardizzoni and Pavarin, 2012], we recall the notion of a preantipode, which plays an important role in further calculations.
Definition 2.1.
[Ardizzoni and Pavarin, 2012, Definition 3.6] Let $(H,\Phi)$ be a coquasi-bialgebra, a preantipode for $H$ is a $\mathbbm{k}$-linear map $\mathbb{S}:H\rightarrow H$ such that, for all $h\in H$,
$\mathbb{S}(h_{1})_{1}h_{2}\otimes\mathbb{S}(h_{1})_{2}$ $=1_{H}\otimes\mathbb{S}(h),$
$\mathbb{S}(h_{2})_{1}\otimes h_{1}\mathbb{S}(h_{2})_{2}$ $=\mathbb{S}(h)\otimes 1_{H},$
$\Phi(h_{1},\mathbb{S}(h_{2}),h_{3})$ $=\varepsilon(h).$
When $H$ is equipped with the structure of a coquasi-Hopf algebra with antipode $(\mathcal{S},\alpha,\beta)$, a preantipode of $H$ always exists.
Lemma 2.2.
[Ardizzoni and Pavarin, 2012, Theorem 3.10] Let $H$ be a coquasi-Hopf algebra with antipode $(\mathcal{S},\alpha,\beta)$, then
$\mathbb{S}:=\beta*\mathcal{S}*\alpha$ (2.1)
is a preantipode for $H$, where * denotes the convolution product.
2.2 Yetter-Drinfeld module categories over coquasi-Hopf algebras
We now turn our attention to the Yetter-Drinfeld module category structure. The definition of Yetter-Drinfeld module category $^{H}_{H}\mathcal{YD}$} over an arbitrary coquasi Hopf-algebra $H$ was already given in [Ardizzoni and Pavarin, 2013]. Since it plays a crucial role in our paper, we recall this definition.
Definition 2.3.
[Ardizzoni and Pavarin, 2013, Definition 3.1] Let $H$ be a coquasi-Hopf algebra with associator $\Phi$. A left-left Yetter–Drinfeld module over $H$ is a triple $(V,\delta_{V},\rhd)$ such that:
  • $(V,\delta_{V})$ is a left comodule of $H$ and we denote $\delta_{V}(v)$ by $v_{-1}\otimes v_{0}$ as usual;
  • $\rhd:H\otimes V\to V$ is a $\mathbbm{k}$-linear map satisfying for all $h,l\in H$ and $v\in V$:
    $(hl)\rhd v=\frac{\Phi(h_{2},(l_{2}\rhd v_{0})_{-1},l_{3})}{\Phi(h_{1},l_{1},v_{-1})\Phi((h_{3}\rhd(l_{2}\rhd v_{0})_{0})_{-1},h_{4},l_{4})}(h_{3}\rhd(l_{2}\rhd v_{0})_{0})_{0},$ (2.2)
    $1_{H}\rhd v=v,$ (2.3)
    $(h_{1}\rhd v)_{-1}h_{2}\otimes(h_{1}\rhd v)_{0}=h_{1}v_{-1}\otimes(h_{2}\rhd v_{0}).$ (2.4)
A morphism $f:(V,\delta_{V},\rhd)\rightarrow(V^{\prime},\delta_{V^{\prime}},\rhd^{\prime})$ is a colinear map $f:(V,\delta_{V})\rightarrow(V^{\prime},\delta_{V^{\prime}})$ such that $f(h\rhd v)=h\rhd^{\prime}f(v)$ for all $h\in H$.
The category $^{H}_{H}\mathcal{YD}$} is a $\mathbbm{k}$-linear braided monoidal abelian category over the field $\mathbbm{k}$. The unit object of $^{H}_{H}\mathcal{YD}$} is $\mathbbm{k}$, which is regarded as an object in $^{H}_{H}\mathcal{YD}$} via trivial structures. For $V,W\in{{}^{H}_{H}\mathcal{YD}}$, the tensor product of Yetter-Drinfeld modules is defined by:
$(V,\delta_{V},\rhd)\otimes(W,\delta_{W},\rhd)=(V\otimes W,\delta_{V\otimes W},\rhd),$
where $\delta_{V\otimes W}$ is given by
$\delta_{V\otimes W}(v\otimes w)=v_{-1}w_{-1}\otimes v_{0}\otimes w_{0},$
and
$h\rhd(v\otimes w)=\frac{\Phi(h_{1},v_{-1},w_{-2})\Phi((h_{2}\rhd v_{0})_{-1},(h_{4}\rhd w_{0})_{-1},h_{5})}{\Phi((h_{2}\rhd v_{0})_{-2},h_{3},w_{-1})}(h_{2}\rhd v_{0})_{0}\otimes(h_{4}\rhd w_{0})_{0}.$ (2.5)
The braiding $c_{V,W}:V\otimes W\rightarrow W\otimes V$ is given by:
$c_{V,W}(v\otimes w)=(v_{-1}\rhd w)\otimes v_{0}.$ (2.6)
In Balan (2009), the authors proved that the category $\mathcal{YD}^{H}_{H}$ of right-right Yetter-Drinfeld module is braided equivalent to $\mathcal{Z}_{l}(\mathcal{M}^{H})$, which is the left center of the category of right comodules of $H$. The proof for $^{H}_{H}\mathcal{YD}$} is completely dual. We need recall the definition of right center here for completeness.
Definition 2.4.
For a coquasi-Hopf algebra $H$ with bijective antipode, the right center of the category of left $H$-comodules $\mathcal{Z}_{r}(^{H}\mathcal{M})$ has objects $(V,c_{-,V})$, where $V\in{{}^{H}\mathcal{M}}$ and $c_{W,V}:W\otimes V\longrightarrow V\otimes W$ is a natural transformation in $^{H}\mathcal{M}$ that satisfies the commutative diagram
$(X\otimes W)\otimes V$$V\otimes(X\otimes W)$$X\otimes(W\otimes V)$$X\otimes(V\otimes W)$$(X\otimes V)\otimes W$$(V\otimes X)\otimes W$$c_{X\otimes W,V}$$a_{X,W,V}$$\operatorname{id}_{X}\otimes c_{W,V}$$a^{-1}_{X,V,W}$$c_{X,V}\otimes\operatorname{id}_{W}$$a_{V,X,W}$
for all $W,X\in^{H}\mathcal{M}$ and $c_{\mathbbm{k},V}=\mathrm{id}_{V}$. A morphism $f:(V,c_{-,V})\longrightarrow(V^{\prime},c_{-,V^{\prime}})$ in the center is a right $H$-comodule map satisfying
$c_{W,V^{\prime}}\circ(\operatorname{id}_{W}\otimes f)=(f\otimes\operatorname{id}_{W})\circ c_{W,V}.$
for all $W\in{{}^{H}\mathcal{M}}$.
The category $\mathcal{Z}_{r}(^{H}\mathcal{M})$ is braided monoidal with tensor product $(V,c_{-,V})\otimes(W,c_{-,W})=(V\otimes W,c_{-,V\otimes W})$.
Lemma 2.5.
Let $H$ be a coquasi-Hopf algebra with bijective antipode. The category $^{H}_{H}\mathcal{YD}$} is braided monoidal isomorphic to the right center of the category of left $H$-comodules $\mathcal{Z}_{r}(^{H}\mathcal{M})$.
Proof.
We construct mutually inverse braided monoidal functors between the two categories.
Let $(V,c_{V,-})$ be an object in $\mathcal{Z}(^{H}\mathcal{M})$, we can define a left linear map on $V$ as follows:
$\rhd:H\otimes V\longrightarrow V,\qquad h\rhd v=(\mathrm{id}_{V}\otimes\varepsilon)c_{H,V}(h\otimes v)$ (2.7)
Let $W\in{{}^{H}\mathcal{M}}$ and $w^{*}\in W^{*}$. Define the map $f_{w^{*}}:W\longrightarrow H,f_{w^{*}}(w)=w_{-1}w^{*}(w_{0})$. Then $f_{w^{*}}$ is $H$-comodule map. By naturality of $c_{-,V}$, we have
$c_{H,V}\circ(f_{w^{*}}\otimes\operatorname{id}_{V})$ $=(\operatorname{id}_{V}\otimes f_{w^{*}})\circ c_{W,V}$
$=c_{H,V}(w_{-1}w^{*}(w_{0})\otimes v).$
Applying $\operatorname{id}_{V}\otimes\varepsilon$ to both sides leads to
$(\operatorname{id}_{V}\otimes w^{*})=w*(w_{0})(w_{-1}\rhd v).$
This implies
$c_{W,V}(w\otimes v)=(w_{-1}\rhd v)\otimes w_{0}.$
for all objects $W\in{{}^{H}\mathcal{M}}$.
Since $c_{\mathbbm{k},V}=\operatorname{id}_{V}$ and the unit map $u:\mathbbm{k}\longrightarrow H$ is a left comodule map, where $\mathbbm{k}$ has trivial comodule structure and $H$ is a $H$-comodule via $\Delta$. We have (2.3) holds:
$1_{H}\rhd v=\varepsilon(1_{H})v=v.$
Next, for $h\in H$, $v\in V$, since $c_{H,V}$ is a left $H$-comodule map,
$(h_{1}\rhd v)_{-1}h_{2}\otimes(h_{1}\rhd v)_{0}\otimes h_{3}=\delta((h_{1}\rhd v)\otimes h_{2})=h_{1}v_{-1}\otimes(h_{2}\rhd v_{0})\otimes h_{3}.$
Applying $\varepsilon$ to last the factor we have (2.3). While relation (2.2) is a consequence of hexagon axiom applied to $W=X=H$. Note that $c_{H\otimes H,V}((h\otimes l)\otimes v)=(h_{1}l_{1}\rhd v)\otimes(h_{2}\otimes l_{2})$. By hexagon axiom it equals to
$\frac{\Phi(h_{2},(l_{2}\rhd v_{0})_{-1},l_{3})}{\Phi(h_{1},l_{1},v_{-1})\Phi((h_{3}\rhd(l_{2}\rhd v_{0})_{0})_{-1},h_{4},l_{4})}(h_{3}\rhd(l_{2}\rhd v_{0})_{0})_{0}\otimes(h_{5}\otimes l_{5}).$
By applying $\varepsilon$ to the second and third factors we get (2.2). Then $(V,\delta_{,}\rhd)$ becomes a left-left $H$ Yetter-Drinfeld module. It is easy to see for any morphism $(V,c_{-,V})\longrightarrow(V^{\prime},c_{-,V^{\prime}}), ff$ will be a morphism in $^{H}_{H}\mathcal{YD}$}.
[…]
The two constructions are clearly inverse to each other. Furthermore, the braiding in $^{H}_{H}\mathcal{YD}$} given by (2.6) coincides with the braiding of the right center under the above correspondence. Hence we obtain a braided monoidal isomorphism. ∎
In order to simplify notation, we recall the following linear maps $p_{R}$, $q_{R}$, $p_{L}$, $q_{L}$$\in\operatorname{Hom}(H\otimes H,\mathbbm{k})$, which is introduced in Balan (2009),
$p_{R}(h,g)$ $=\Phi^{-1}(h,g_{1}\beta(g_{2}),\mathcal{S}(g_{3})),$
$q_{R}(h,g)$ $=\Phi(h,g_{3},\alpha(\mathcal{S}^{-1}(g_{2}))\mathcal{S}^{-1}(g_{1})),$
$p_{L}(h,g)$ $=\Phi(\mathcal{S}^{-1}(h_{3})\beta(\mathcal{S}^{-1}(h_{2})),h_{1},g),$
$q_{L}(h,g)$ $=\Phi^{-1}(\mathcal{S}(h_{1})\alpha(h_{2}),h_{3},g),$
for any $h,g\in H$. These maps satisfy various compatibility relations with the antipode and associator, see Balan (2009) for the complete list.
Remark 2.6.
Let $H$ be a coquasi-Hopf algebra with bijective antipode.
(1) [Li and Liu, 2025b, Lemma 2.6] The equation (2.2) is equivalent to
$(h\rhd v)_{-1}\otimes(h\rhd v)_{0}=p_{R}((h_{3}\rhd v_{0})_{-1},h_{4})q_{R}(h_{1}v_{-2},\mathcal{S}(h_{6}))(h_{2}v_{-1})\mathcal{S}(h_{5})\otimes(h_{3}\rhd v_{0})_{0}.$ (2.8)
(2) [Li and Liu, 2025b, Lemma 2.6] The equation (2.2) implies:
$v_{-1}\otimes(h\rhd v_{0})=q_{L}(h_{3},v_{-1})p_{L}(\mathcal{S}(h_{1}),(h_{4}\rhd v_{0})_{-1}h_{6})\mathcal{S}(h_{2})((h_{4}\rhd v_{0})_{-2}h_{5})\otimes(h_{4}\rhd v_{0})_{0}.$ (2.9)
(3)[Lu et al., , Theorem 2.8] The inverse braiding is given by:
$c_{V,W}^{-1}(w\otimes v)$ $=\Phi^{-1}(\mathcal{S}(v_{-2}),w_{-1},v_{-5})q_{L}(\mathcal{S}^{-1}(v_{-1}),w_{-2}v_{-6})p_{R}((\mathcal{S}^{-1}(v_{-3}\rhd w_{0})_{-1}),\mathcal{S}^{-1}(v_{-4}))$ (2.10)
$v_{0}\otimes(\mathcal{S}^{-1}(v_{-3})\rhd w_{0})_{0}.$
(4) [Lu et al., , Proposition 2.5] The coquasi-Hopf algebra $H$ itself can be viewed as an object in $^{H}_{H}\mathcal{YD}$} via the structures
$\delta_{H}(h)=h_{1}\mathcal{S}(h_{3})\otimes h_{2},$ (2.11)
$h\rhd h^{\prime}=\Phi(h_{1},h_{1}^{\prime},\mathcal{S}(h_{7}^{\prime}))\Phi(h_{2}h_{2}^{\prime},\mathcal{S}(h_{4}h_{4}^{\prime}),h_{7})g(h_{5},h_{5}^{\prime})q_{R}(\mathcal{S}(h_{6}^{\prime}),\mathcal{S}(h_{6}))h_{3}h_{3}^{\prime}.$
Here
$g(h,h^{\prime}):=\Phi^{-1}(\mathcal{S}(h_{1}h_{1}^{\prime}),h_{3}h_{3}^{\prime},\mathcal{S}(h_{5}^{\prime})\mathcal{S}(h_{5}))\chi(h_{4},h_{4}^{\prime})\alpha(h_{2}h_{2}^{\prime}),$
and
$\chi(h,h^{\prime})=\Phi(h_{1}h^{\prime}_{1},\mathcal{S}(h_{5}^{\prime}),\mathcal{S}(h_{4}))\Phi^{-1}(h_{2},h_{2}^{\prime},\mathcal{S}(h_{4}^{\prime}))\beta(h_{3})\beta(h_{3}^{\prime}).$
2.3 Bosonization for coquasi-Hopf algebras
Bosonization is a fundamental construction that builds a new Hopf algebra from a Hopf algebra in the Yetter–Drinfeld category. In [Bulacu and Nauwelaerts, 2002], the authors introduced bosonization for quasi-Hopf algebras. Dually, bosonization for coquasi-Hopf algebras is given in Ardizzoni and Pavarin (2013). We recall this construction, as it provides the setting for our later study of reflections via projection and coinvariants.
Let $H$ be a coquasi-Hopf algebra with associator $\Phi$. Suppose $R$ is a Hopf algebra in $^{H}_{H}\mathcal{YD}$}, we denote
$r^{1}\otimes r^{2}:=\Delta_{R}(r).$
Lemma 2.7.
[Ardizzoni and Pavarin, 2013, Theorem 5.2] Let us consider on $M=R\otimes H$ the following structures:
$(r\otimes h)(s\otimes k)$ $=\frac{\Phi(h_{2},s_{-1},k_{2})\Phi(r_{-1},(h_{3}\rhd s_{0})_{-1},h_{5}k_{4})}{\Phi(r_{-2},h_{1},s_{-2}k_{1})\Phi((h_{3}\rhd s_{0})_{-2},h_{4},k_{3})}r_{0}(h_{3}\rhd s_{0})_{0}\otimes h_{6}k_{5},$
$u_{M}(k)$ $=k1_{R}\otimes 1_{M},$
$\Delta_{M}(r\otimes h)$ $=\Phi^{-1}(r_{-1}^{1},r_{-2}^{2},h_{1})r_{0}^{1}\otimes r_{-1}^{2}h_{2}\otimes r_{0}^{2}\otimes h_{3},$
$\varepsilon_{M}(r\otimes h)$ $=\varepsilon_{R}(r)\varepsilon_{M}(h),$
$\Phi_{M}((r\otimes h),(s\otimes k),(t\otimes l))$ $=\varepsilon_{R}(r)\varepsilon_{R}(s)\varepsilon_{R}(t)\Phi(h,k,l),$
where $r,s,t\in R$, $h,k,l\in H$. With above operations, $R\otimes H$ is a coquasi-bialgebra, we denote by R#HR\#H.
Remark 2.8.
The bosonization R#HR\#H can be further equipped with an antipode, making it a coquasi-Hopf algebra. The antipode $(\mathcal{S},\alpha,\beta)$ is given by:
$\alpha(r\#h)=0,\ \ \alpha(1\#h)=\alpha(h),$
$\beta(r\#h)=0,\ \ \beta(1\#h)=\beta(h),$
$\mathcal{S}(r\#h)=m_{R\#H}\circ(\mathcal{S}_{H}\otimes\mathcal{S}_{R})\circ c_{R,H}(r\otimes h)=\mathcal{S}_{H}(r_{-1}\rhd h)\mathcal{S}_{R}(r_{0}).$
In particular, if $r$ is a primitive element in $R$, then
$\mathcal{S}(r\#1)=\mathcal{S}_{H}(r_{-2})\alpha(r_{-1})\mathcal{S}_{R}(y_{0})=-\mathcal{S}_{H}(r_{-2})\alpha(r_{-1})y_{0}.$
The antipode $\mathcal{S}$ is bijective if and only if $\mathcal{S}_{H}$ and $\mathcal{S}_{R}$ are bijective.
Now suppose $N$ are $H$ are both coquasi-Hopf algebras. Furthermore, assume there exists morphisms of coquasi-Hopf algebras
$\pi:N\rightarrow H\ \text{ and }\sigma:H\rightarrow N$
such that $\pi\sigma=\operatorname{id}_{N}$.
Lemma 2.9.
[Ardizzoni and Pavarin, 2013, Theorem 5.8] Under above assumptions, $L:=N^{\operatorname{co}H}$ is a Hopf algebra in $^{H}_{H}\mathcal{YD}$}. For all $a\in N$, we define the following linear map
$\tau(a)=\Phi_{N}(a_{1},(\sigma\circ\mathbb{S}_{H}\circ\pi(a_{3}))_{1},a_{4})a_{2}(\sigma\circ\mathbb{S}_{H}\circ\pi(a_{3}))_{2}.$ (2.12)
Then $\tau:N\rightarrow L$ is well-defined. For all $r,s\in L$, $h\in H$, $k\in\mathbbm{k}$, the Yetter-Drinfeld module structure of $L$ is given by
$\operatorname{ad}(h)(r)$ $:=\tau(\sigma(h)r)=\Phi_{H}(h_{1}r_{-1},\mathbb{S}(h_{3})_{1},h_{4})(h_{2}r_{0})\mathbb{S}(h_{3})_{2},$
$r_{-1}\otimes r_{0}$ $:=\rho_{R}(r):=\pi(r_{1})\otimes r_{2}.$
Moreover, $L$ is a Hopf algebra in $^{H}_{H}\mathcal{YD}$} via
$m_{L}(r\otimes s):=rs,\ \ \ \ \ \ \ u_{L}(k)=k1_{N},$
$\Delta_{L}(r):=\tau(r_{1})\otimes\tau(r_{2}),\ \ \ \ \varepsilon_{L}(r):=\varepsilon_{N}(r),$
$\mathcal{S}_{R}=\operatorname{ad}\circ(\operatorname{id}_{H}\otimes\tau\circ\mathcal{S}_{N}\circ i)\circ\rho_{R},$
where $i:L\rightarrow N$ is the embedding map. Furthermore, there is an isomorphism of coquasi-Hopf algebra $\phi:L\#H\rightarrow N$ given by
$\phi(r\otimes h)=r\sigma(h),\ \ \phi^{-1}(a)=\tau(a_{1})\otimes\pi(a_{2}).$
The following condition occurs frequently in this paper. We give an example here.
Example 2.10.
Given $R,R^{\prime}$ be two Hopf algebras in $^{H}_{H}\mathcal{YD}$} together with a surjective Hopf algebra map $\pi^{\prime}:R^{\prime}\rightarrow R$ and an injective map $\sigma^{\prime}:R\rightarrow R^{\prime}$ such that $\pi^{\prime}\circ\sigma^{\prime}=\operatorname{id}_{R}$. Let $M^{\prime}=R^{\prime}\#H$, and M=R#HM=R\#H. Obviously, these maps naturally extend to yield a surjective coquasi-Hopf algebra map $\pi:M^{\prime}\rightarrow M$ and an injective coquasi-Hopf algebra map $\sigma:M\rightarrow M^{\prime}$, such that $\pi\circ\sigma=\operatorname{id}_{M}$.
Let $K:=M^{\prime\operatorname{co}M}$, then $L$ is an object in $^{M}_{M}\mathcal{YD}$}, with the following structures:
$\operatorname{ad}:M\otimes K\rightarrow K,h\otimes k\mapsto\Phi_{H}(h_{1}k_{-1},\mathbb{S}_{H}(h_{3})_{1},h_{4})(h_{2}k_{0})\mathbb{S}_{H}(h_{3})_{2},$
$\operatorname{\rho}:K\rightarrow H\otimes K,k\mapsto\pi(k_{1})\otimes k_{2}.$
To analyze the linear map $\operatorname{ad}$ in concrete calculations, the following computation is useful. Let $y\in R$ be a primitive element in $^{H}_{H}\mathcal{YD}$}, One can compute the preantipode $\mathbb{S}(y)$ in $M$.
$\mathbb{S}(y)=\beta*\mathcal{S}*\alpha(y)=\beta(y_{1})\mathcal{S}_{M}(y_{2})\alpha(y_{3})=\beta(y_{-1})\mathcal{S}_{M}(y_{0})=-\beta(y_{-3})\mathcal{S}_{H}(y_{-2})\alpha(y_{-1})y_{0}.$
Then we can proceed to calculate $\operatorname{ad}(y)(x)$, where $x\in K$,
$\operatorname{ad}(y)(x)$ $=yx+(y_{-1}x)\mathbb{S}(y_{0})$ (2.13)
$=yx+(y_{-4}x)\beta(y_{-3})(\mathcal{S}_{H}(y_{-2})\alpha(y_{-1})y_{0})$
$=yx-\Phi(y_{-7}x_{-1},\mathcal{S}(y_{-4}),y_{-1})\beta(y_{-6})((y_{-5}x_{0})\mathcal{S}(y_{-3}))\alpha(y_{-2})y_{0}.$
On the other hand, note that for all $h\in H$, $\Delta_{H}(\mathbb{S}(h))=\beta(h_{1})\mathcal{S}(h_{3})\otimes\mathcal{S}(h_{2})\alpha(h_{4})$. Therefore, we have
$yx-\operatorname{ad}(y_{-1})(x)y_{0}$ $=yx-\Phi(y_{-4}x_{-1},\mathbb{S}(y_{-2})_{1},y_{-1}))((y_{-3}x_{0})\mathbb{S}(y_{-2})_{2})y_{0}$
$=yx-\Phi(y_{-7}x_{-1},\mathcal{S}(y_{-4}),y_{-1})\beta(y_{-6})((y_{-5}x_{0})\mathcal{S}(y_{-3}))\alpha(y_{-2})y_{0}.$
This implies
$\operatorname{ad}(y)(x)=yx-\operatorname{ad}(y_{-1})(x)y_{0}.$ (2.14)
2.4 Yetter-Drinfeld modules in arbitrary braided monoidal category
In Section 3, we will investigate Yetter-Drinfeld modules over a Hopf algebra $A$ in the category $_{H}^{H}\mathcal{YD}$. Since $^{H}_{H}\mathcal{YD}$} is braided monoidal, we require the general framework of Yetter-Drinfeld modules in arbitrary braided monoidal categories. We recall the definition from Bespalov (1997), one may refer to [Heckenberger and Schneider, [2020] ©2020, Section 3.4] for more details. Now we let $\mathcal{C}$ be a braided monoidal category and $R$ a Hopf algebra in $\mathcal{C}$. The category of Yetter-Drinfeld modules $^{R}_{R}\mathcal{YD}(\mathcal{C})$} is well-defined.
Definition 2.11.
Let $R$ be a Hopf algebra with bijective antipode in a braided monoidal category $\mathcal{C}$. Suppose that $X$ is a left module and left comodule over $R$ with structure map $\rho_{X}$ and $\delta_{X}$. The triple $(X,\rho_{X},\delta_{X})$ is a Yetter-Drinfeld module if in addition, in $\operatorname{Hom}_{\mathcal{C}}(R\otimes X,R\otimes X)$,
$(\mu_{X}\otimes\operatorname{id})$ $\circ a_{R,R,X}^{-1}\circ(\operatorname{id}\otimes c_{X,R})\circ a_{R,X,R}\circ(\delta_{X}\circ\rho_{X}\otimes\operatorname{id})\circ a^{-1}_{R,X,R}\circ(\operatorname{id}\otimes c_{R,X})\circ a_{R,R,X}\circ(\Delta_{R}\otimes\operatorname{id})$
$=(\mu_{R}\otimes\rho_{X})\circ a^{-1}_{R,R,R\otimes X}\circ(\operatorname{id}\otimes a_{R,R,X}\circ(\operatorname{id}\otimes c_{R,R})\otimes\operatorname{id})\circ(\operatorname{id}\otimes a^{-1}_{R,R,X})\circ a_{R,R,R\otimes X}\circ(\Delta_{R}\otimes\delta_{X}).$
It is known that the category $^{R}_{R}\mathcal{YD}(\mathcal{C})$} is a braided monoidal category. For $X,Y\in{{}^{R}_{R}\mathcal{YD}(\mathcal{C})}$ the braiding isomorphism in $^{R}_{R}\mathcal{YD}(\mathcal{C})$} is given by
$c_{X,Y}^{\mathcal{YD}}$ $:X\otimes Y\rightarrow Y\otimes X,$
$c_{X,Y}^{\mathcal{YD}}$ $:=(\rho_{Y}\otimes\operatorname{id}_{X})\circ a^{-1}_{R,Y,X}\circ(\operatorname{id}_{R}\otimes c_{X,Y})\circ a_{R,X,Y}\circ(\delta_{X}\otimes\operatorname{id}_{Y}).$
Remark 2.12.
(1) According to the definition of $c_{X,Y}^{\mathcal{YD}}$, it is not hard to see, if $X\in{{}^{R}\mathcal{C}}$ and $Y\in{{}_{R}\mathcal{C}}$, then $c_{X,Y}^{\mathcal{YD}}$ is a morphism in $\mathcal{C}$.
(2) [Heckenberger and Schneider, [2020] ©2020, Proposition 3.4.5] Let $V$ be an object in $\mathcal{C}$, $(V,\lambda)\in{{}_{R}\mathcal{C}}$, and $(V,\delta)\in{{}^{R}\mathcal{C}}$. Then the following are equivalent.
  1. (i)
    $(V,\lambda,\delta)$ is a left Yetter-Drinfeld module over $R$.
  2. (ii)
    For all $X\in{{}_{R}\mathcal{C}}$, $c_{V,X}^{\mathcal{YD}}$ is a morphism in $_{R}\mathcal{C}$}.
  3. (iii)
    $c_{V,R}^{\mathcal{YD}}$ is a morphism in $_{R}\mathcal{C}$, where $R$ is a left $R$-module by the multiplication in $R$.
  4. (iv)
    For all $X\in{{}^{R}\mathcal{C}}$, $c_{X,V}^{\mathcal{YD}}$ is a morphism in $^{R}\mathcal{C}$}.
(3) In general, let $(\mathcal{C},\otimes,\textbf{1},c)$ be a braided monoidal category, we can define the reverse monoidal category $\overline{\mathcal{C}}:=(\mathcal{C},\otimes,\textbf{1},\overline{c})$, where for $X,Y\in\mathcal{C}$,
$\overline{c}_{X,Y}:=c^{-1}_{Y,X}:X\otimes Y\rightarrow Y\otimes X.$
For $X\in{{}_{R}\mathcal{C}}$ and $Y\in{{}^{R}\mathcal{C}}$, let
$\overline{c}^{\mathcal{YD}}_{X,Y}=\overline{c}_{X,Y}\circ(\lambda_{X}\otimes\operatorname{id}_{Y})\circ((\mathcal{S}^{-1}_{R}\otimes\operatorname{id}_{X})\otimes\operatorname{id}_{Y})\circ(\overline{c}_{X,R}\otimes\operatorname{id}_{Y})\circ a^{-1}_{X,R,Y}\circ(\operatorname{id}_{X}\otimes\delta_{Y}).$ (2.15)
By [Heckenberger and Schneider, [2020] ©2020, Proposition 3.4.8], $c^{YD}_{Y,X}$ is an isomorphism in $\mathcal{C}$ with inverse $\overline{c}_{X,Y}^{\mathcal{YD}}$.
The definition of the bosonization naturally extends to the case that $K$ is a Hopf algebra in the braided monoidal category $^{R}_{R}\mathcal{YD}(\mathcal{C})$}, where $\mathcal{C}$ is now an arbitrary braided monoidal category. We collect several key results from [Bespalov, 1997,Bespalov, 1995,Alonso Alvarez and Fernández Vilaboa, 2000] that will be important for the construction.
Lemma 2.13.
[Bespalov, 1997, Theorem 3.9.5] There is a braided monoidal isomorphism from the Hopf bimodule category to the Yetter-Drinfeld module category:
${{}^{R}_{R}\mathcal{C}^{R}_{R}}\cong{{}^{R}_{R}\mathcal{YD}(\mathcal{C})}.$
Lemma 2.14.
[Bespalov, 1995, Proposition 4.2.3] Let $A$ be a Hopf algebra in $\mathcal{C}$ and $K$ a Hopf algebra in $^{A}_{A}\mathcal{YD}(\mathcal{C})$}. There is an obvious isomorphism of braided monoidal categories
${{}^{K\#A}_{K\#A}\mathcal{YD}}(\mathcal{C})\cong{{}^{K}_{K}\mathcal{YD}}({{}^{A}_{A}\mathcal{YD}(\mathcal{C})}).$ (2.16)
Lemma 2.15.
[Alonso Alvarez and Fernández Vilaboa, 2000, Theorem 3.2] Let $H$ and $A$ be Hopf algebras in a braided monoidal category $\mathcal{C}$. Let $\pi:H\to A$ and $\iota:A\to H$ be Hopf algebra morphisms such that $\pi\circ\iota=\mathrm{id}_{A}$. If $\mathcal{C}$ has equalizers and $A\otimes(-)$ preserves equalizers, there is a Hopf algebra $R$ in the braided monoidal category $^{A}_{A}\mathcal{YD}(\mathcal{C})$}, such that
$H\cong R\#A.$
Remark 2.16.
Let $H$ be a coquasi-Hopf algebra with bijective antipode.
(1) By [Bulacu and Nauwelaerts, 2002, Proposition 3.9, Lemma 4.5] and Lemma 2.13, we have a braided monoidal equivalence
${{}^{R\#H}_{R\#H}\mathcal{YD}}\cong{{}^{R}_{R}\mathcal{YD}(\mathcal{C})},$
where $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$.
(2) Since $^{H}_{H}\mathcal{YD}$} is an abelian category, the conditions in Lemma 2.15 hold automatically. That is, we can talk about bosonization and taking coinvariants in $^{H}_{H}\mathcal{YD}$} without any restrictions.
3 Rational Yetter-Drinfeld module over coquasi-Hopf algebras
In this section, $H$ always represents a coquasi-Hopf algebra with bijective antipode, and $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$ denotes the Yetter-Drinfeld module category of $H$ . We investigate rational Yetter-Drinfeld modules in $^{H}_{H}\mathcal{YD}$}, a notion that will be important for our later study of Nichols algebras and their reflections.
3.1 Rational modules in $^{H}_{H}\mathcal{YD}$}
Since $H$ is a non-associative algebra, we can not talk about modules over $H$ in the usual sense. However, if $R$ is an algebra in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$, then the category $_{R}\mathcal{C}$ of left $R$-modules in $\mathcal{C}$ is well-defined. Motivated by [Heckenberger and Schneider, [2020] ©2020, Definition 12.2.3], we now introduce the concept of rationality under the setting of coquasi-Hopf algebras.
Definition 3.1.
Let $R=\bigoplus_{n\geq 0}R(n)$ be an $\mathbb{N}_{0}$-graded Hopf algebra in $\mathcal{C}$.
(1) A left module $(X,\lambda_{X})$ over $R$ in $\mathcal{C}$ is called rational if for any $x\in X$, there exists a natural number $n_{0}$ such that $\lambda(R(n)\otimes x)=0$ for all $n\geq n_{0}$. We denote the category of left rational module over $R$ in $\mathcal{C}$ by ${{}_{R}\mathcal{C}}_{\operatorname{rat}}$.
(2) A left Yetter-Drinfeld module $X$ over $^{R}_{R}\mathcal{YD}(\mathcal{C})$} is called rational if $X$ is rational as a module of $R$ in $\mathcal{C}$. We denote the corresponding category by ${{}^{R}_{R}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$.
Remark 3.2.
It is worth noting that for a locally finite $\mathbb{N}_{0}$-graded Hopf algebra $R$, this graded definition of rationality coincides with the classical one. Intuitively, the condition $\lambda(R(n)\otimes x)=0$ for large $n$ implies that the action on any element $x$ factors through a finite-dimensional truncation of $R$, which recovers the standard notion of rationality in the non-graded setting.
A natural question arises naturally, are rational modules closed under tensor products. The following lemma ensures this.
Lemma 3.3.
Let $R=\bigoplus_{n\geq 0}R(n)$ be an $\mathbb{N}_{0}$-graded Hopf algebra in $\mathcal{C}$.
(1) The category ${{}_{R}\mathcal{C}}_{\operatorname{rat}}$ is a monoidal subcategory of $_{R}\mathcal{C}$}, which is closed under direct sums, subobjects and quotient objects.
(2) The category ${{}^{R}_{R}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$ is a monoidal subcategory of $^{R}_{R}\mathcal{YD}(\mathcal{C})$}, which is closed under direct sums, subobjects and quotient objects.
Proof.
(1) Let $V,W\in{{}_{R}\mathcal{C}}_{\operatorname{rat}}$ with $R$-action $\lambda_{V}$, $\lambda_{W}$ respectively. It is standard that the tensor product $V\otimes W$ is an object in $_{R}\mathcal{C}$, with $R$-action
$\lambda_{V\otimes W}$ $=(\lambda_{V}\otimes\lambda_{W})\circ(a_{R,V,R}^{-1}\otimes\operatorname{id}_{W})\circ a_{R\otimes V,R,W}\circ((\operatorname{id}_{R}\otimes c_{R,V})\otimes\operatorname{id}_{W})$
$\circ(a_{R,R,V}\otimes\operatorname{id}_{W})\circ a^{-1}_{R\otimes R,V,W}\circ(\Delta_{R}\circ\operatorname{id}_{V\otimes W}).$
Now take $r\in R(m)$, we have
$\lambda_{V\otimes W}(r\otimes(v\otimes w))=\frac{\Phi(r^{2}_{-4},v_{-1},w_{-3})\Phi(r^{1}_{-1},(r^{2}_{-3}\rhd v_{0})_{-1},r^{2}_{-1}w_{-1})}{\Phi(r^{1}_{-2},r^{2}_{-5},v_{-2}w_{-4})\Phi((r^{2}_{-3}\rhd v_{0})_{-2},r^{2}_{-2},w_{-2})}\lambda_{V}(r^{1}_{0}\otimes(r^{2}_{-3}\rhd v_{0})_{0})\otimes\lambda_{W}(r^{2}_{0}\otimes w_{0}).$
Since $r^{2}_{-3}\in H$, and $V$ is an object in $^{H}_{H}\mathcal{YD}$}, we have $(r^{2}_{-3}\rhd v_{0})_{0}\in V$. The expression $\lambda_{V\otimes W}(r\otimes(v\otimes w))$ is a finite sum. Thus, by the definition of the rational module, there exists a natural number $n_{1}$ such that for $n\geq n_{1}$,
$\lambda_{V}(R(n)\otimes(r^{2}_{-3}\rhd v_{0})_{0})=0.$
Similarly, there is a natural number $n_{2}$ such that
$\lambda_{W}(R(n)\otimes w_{0})=0,$
for all $n\geq n_{2}$.
Now we assume $m\geq 2\operatorname{max}\{n_{1},n_{2}\}$, since $R$ is $\mathbb{N}_{0}$-graded, we have $\Delta_{R}(r)\in\bigoplus_{p+q=m}R(p)\otimes R(q).$ Therefore
$\lambda_{V\otimes W}(R(m)\otimes(v\otimes w))=0.$
This implies $V\otimes W\in{{}_{R}\mathcal{C}_{\operatorname{rat}}}$. Therefore, ${{}_{R}\mathcal{C}}_{\operatorname{rat}}$ is a monoidal subcategory of $^{R}_{R}\mathcal{YD}(\mathcal{C})$}. The category ${{}_{R}\mathcal{C}}_{\operatorname{rat}}$ is closed under arbitrary direct sums, subobjects and quotient objects.
(2) The forgetful functor ${{}^{R}_{R}\mathcal{YD}(\mathcal{C})}\rightarrow{{}_{R}\mathcal{C}}$ is monoidal. Thus, for $V,W\in{{}^{R}_{R}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}},$ we have $V\otimes W\in{{}_{R}\mathcal{C}}_{\operatorname{rat}}$. By definition it implies $V\otimes W\in{{}^{R}_{R}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$.
3.2 Properties of pairings
We now turn to the study of pairings in the category $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$, which will later facilitate the construction of dualities between comodules and rational modules.
Let $A$ and $B$ be objects in $\mathcal{C}$, and $\omega:A\otimes B\rightarrow\mathbbm{k}$ is a morphism in $\mathcal{C}$, named pairing. For subsets $X\subseteq A$ and $Y\subseteq B$, we define
$X^{\perp}$ $:=\{b\in B\mid\omega(x,b)=0,\ \text{for all}\ x\in X\},$
$Y^{\perp}$ $:=\{a\in A\mid\omega(a,y)=0,\ \text{for all}\ y\in Y\}.$
We say $\omega$ is non-degenerate, if $A^{\perp}=0$ and $B^{\perp}=0$.
Lemma 3.4.
If $E$ is a subobject of $A$, then $E^{\perp}$ is a subobject of $B$. Similarly, if $F$ is a subobject of $B$, then $F^{\perp}$ is a subobject of $A$.
Proof.
Let $E$ be a subobject of $A$, we need to show that $E^{\perp}$ is an object in $\mathcal{C}$. The proof of $E^{\perp}$ being a $H$-comodule is the same as in the Hopf algebra case, see [Heckenberger and Schneider, 2013a, Lemma 2.3].
To show $H\rhd E^{\perp}\subseteq E^{\perp}$. One need to verify that for any $b\in E^{\perp}$, $h\in H$ and $e\in E$,
$\omega(e,h\rhd b)=0.$
By hexagon axiom in $\mathcal{C}$, we have the following commutative diagram:
$A\otimes(H\otimes B)$$A\otimes(B\otimes H)$$(A\otimes H)\otimes B$$(H\otimes A)\otimes B$$H\otimes(A\otimes B)$$(A\otimes B)\otimes H$$\operatorname{id}\otimes c_{H,B}$$a^{-1}_{A,H,B}$$c^{-1}_{H,A}\otimes\operatorname{id}$$a_{H,A,B}$$c_{H,A\otimes B}$$a_{A,B,H}^{-1}$
Since $E$ is an object in $\mathcal{C}$, for simplicity, we write $c_{A,H}^{-1}(e\otimes h)=\sum_{1\leq i\leq n}h^{i}\otimes e^{i}$ for some natural number $n$. Here $h^{i}\in H$, $e^{i}\in E$. Now
$c_{H,A\otimes B}\circ a_{H,A,B}\circ(c_{A,H}^{-1}\otimes\operatorname{id})((e\otimes h)\otimes b)$ $=\sum_{1\leq i\leq n}c_{H,A\otimes B}\circ a_{H,A,B}((h^{i}\otimes e^{i})\otimes b)$ (3.1)
$=\sum_{1\leq i\leq n}\Phi(h^{i}_{1},e^{i}_{-1},b_{-1})c_{H,A\otimes B}(h^{i}_{2}\otimes(e^{i}_{0}\otimes b_{0}))$
$=\sum_{1\leq i\leq n}\Phi(h^{i}_{1},e^{i}_{-1},b_{-1})(h^{i}_{2}\rhd(e_{0}^{i}\otimes b_{0}))\otimes h^{i}_{3}.$
Here $b_{0}\in E^{\perp}$, because we have shown $E^{\perp}$ is a $H$-comodule. Also, $e_{0}^{i}\in E$ for all $i$. Note that $\omega$ is a morphism in $\mathcal{C}$, and the braiding $c$ is a natural isomorphism, we have
$(\omega\otimes\operatorname{id})\circ c_{H,A\otimes B}\circ a_{H,A,B}\circ(c_{A,H}^{-1}\otimes\operatorname{id})((e\otimes h)\otimes b)=\sum_{1\leq i\leq n}\Phi(h^{i}_{1},e^{i}_{-1},b_{-1})\omega(e_{0}^{i},b_{0})\otimes h^{i}_{2}=0.$ (3.2)
Note that $a_{A,H,B}^{-1}(e\otimes(h\otimes b))\in(A\otimes H)\otimes B$, thus by the hexagon axiom, we obtain:
$(\omega\otimes\varepsilon)\circ a_{A,B,H}^{-1}\circ(\operatorname{id}\otimes c_{H,B})\circ a_{A,H,B}(a_{A,H,B}^{-1}(e\otimes(h\otimes b)))=0.$
That is
$\Phi(e_{-1},(h_{1}\rhd b)_{-1},h_{2})(\omega\otimes\varepsilon)((e_{0}\otimes(h_{1}\rhd b)_{0})\otimes h_{3})=\Phi*(\omega\otimes\varepsilon)(e\otimes(h_{1}\rhd b)\otimes h_{2})=0.$ (3.3)
Note that $\Phi$ is convolution invertible, thus
$\omega(e,h\rhd b)=0$ (3.4)
for all $e\in E$, $b\in E^{\perp}$, $h\in H$. Consequently, $E^{\perp}$ is a subobject of $B$. The proof of another direction is similar, we omit it for simplicity. ∎
The following construction of the Yetter-Drinfeld submodule generated by a subcomodule is crucial for our subsequent analysis of rational modules.
Let $V\in{}_{H}^{H}\mathcal{YD}$ and let $V^{\prime}\subseteq V$ be an $H$-subcomodule. We define the Yetter-Drinfeld submodule generated by $V^{\prime}$, denoted by $\langle V^{\prime}\rangle$, as the smallest Yetter-Drinfeld submodule of $V$ containing $V^{\prime}$.
We provide an explicit filtration to construct $\langle V^{\prime}\rangle$.
  • Set $X_{0}:=V^{\prime}$.
  • Assume $X_{n}$ is defined. Let $Y_{n}:=\text{span}\{h\rhd x\mid h\in H,x\in X_{n}\}$.
  • Define $X_{n+1}$ as the smallest $H$-subcomodule of $V$ containing $Y_{n}$:
    $X_{n+1}:=\sum_{y\in Y_{n}}\text{span}\{y_{0}\mid\delta_{V}(y)=y_{-1}\otimes y_{0}\}.$
Lemma 3.5.
With the notation above, $\langle V^{\prime}\rangle=\bigcup_{n\geq 0}X_{n}$.
Proof.
It is direct to verify that $\bigcup_{n\geq 0}X_{n}$ is closed under the $H$-action and $H$-coaction by construction. Since any Yetter-Drinfeld submodule containing $V^{\prime}$ must contain $X_{0}$ and be closed under the action and coaction, it must contain each $X_{n}$. Thus, the union is minimal. ∎
Let $V,W\in C$, and $\omega$ a pairing of $A,B$ in $C$. We denote by $\operatorname{Hom}_{C,\text{rat}}(A\otimes V,W)$ the set of all $g$ in $\operatorname{Hom}_{C}(A\otimes V,W)$ such that for all $v\in V$ there is a finite-dimensional subobject $F\subseteq B$ in $C$ with $g(F^{\perp}\otimes v)=0$.
Proposition 3.6.
Let $A,B\in C$, $\omega$ a non-degenerate pairing of $A,B$ in $C$, and $W\in C$. Assume that for every $b\in B$ there is a finite-dimensional subobject $F\subseteq B$ in $C$ containing $b$. Then for all $V\in C$, the map
$G_{V}:\operatorname{Hom}_{C}(V,B\otimes W)\to\operatorname{Hom}_{C,\operatorname{rat}}(A\otimes V,W),$
$f\mapsto(A\otimes V\xrightarrow{\operatorname{id}_{A}\otimes f}A\otimes(B\otimes W)\xrightarrow{a^{-1}_{A,B,W}}(A\otimes B)\otimes W\xrightarrow{\omega\otimes\operatorname{id}}W)$
is bijective.
Proof.
Step1:
We first show the map $G_{V}$ is well-defined and injective. Let $f\in\operatorname{Hom}_{C}(V,B\otimes W)$, and $g=G_{V}(f)$. For each $v\in V$, there is a finite-dimensional subobject $F\subseteq B$ with $f(v)=\sum_{i\in I}b^{i}\otimes w^{i}\in F\otimes W$. Now for any $a\in F^{\perp}$, by definition of $G_{V}$ and Lemma 3.4, we have
$g(a\otimes v)=\Phi(a_{-1},b^{i}_{-1},w^{i}_{-1})\omega(a_{0},b^{i}_{0})\otimes w^{i}_{0}=0.$
This shows $G_{v}(f)\in\operatorname{Hom}_{\mathcal{C},\operatorname{rat}}(A\otimes V,W)$. If $g=0$, since $\omega$ is non-degenerate, we deduce that $f=0$. Therefore $G_{V}$ is injective.
Step2:
Assume that $B$ is finite-dimensional. Then $B$ has a left dual $B^{*}=\operatorname{Hom}(B,\mathbbm{k})$. Since $\omega$ is non-degenerate, there is an isomorphism $A\cong B^{*}$ in $\mathcal{C}$ via the following isomorphism:
$\operatorname{Hom}_{\mathcal{C}}(A\otimes B,\mathbbm{k})\cong\operatorname{Hom}_{\mathcal{C}}(A,B^{*}).$
Therefore $A$ is a left dual of $B$, which is finite-dimensional. We have such an isomorphism
$\operatorname{Hom}_{\mathcal{C}}(V,B\otimes W)\cong\operatorname{Hom}_{\mathcal{C}}(A\otimes V,W).$
Since $\operatorname{Hom}_{\mathcal{C},\operatorname{rat}}(A\otimes V,W)\subseteq\operatorname{Hom}_{\mathcal{C}}(A\otimes V,W),$ and $G_{V}$ is injective, we deduce that $G_{V}$ is bijective when $B$ is finite-dimensional.
Step3:
We now extend the bijectivity to the general case when $B$ may be infinite-dimensional. Let $V^{\prime}\subseteq V$ be a finite-dimensional $H$-comodule, we denote $\langle V^{\prime}\rangle$ as the Yetter-Drinfeld submodule of $V$ generated by $V^{\prime}$. Assume that $\langle V^{\prime}\rangle=V$, we are going to prove that $G_{V}$ is surjective in this case. Let $g\in\operatorname{Hom}_{\mathcal{C},\operatorname{rat}}(A\otimes V,W)$. Since $V^{\prime}$ is finite-dimensional, the rationality of $g$ implies the existence of a finite-dimensional subobject $F\subseteq B$ in $\mathcal{C}$ such that $g(F^{\perp}\otimes V^{\prime})=0$. By Lemma 3.4, $F^{\perp}$ is an object in $\mathcal{C}$. Since $g$ is a morphism in $\mathcal{C}$, by similar method in proof of Lemma 3.4, for all $h\in H$, we have
$g(F^{\perp}\otimes(h\rhd V^{\prime}))=0.$
By construction of $\langle V^{\prime}\rangle$, if we denote $V^{\prime}$ as $X_{0}$, this shows $g(F^{\perp}\otimes Y_{0})=0$. This property implies that for all $h\in H$, $v\in V^{\prime}$, we have $f(h\rhd v)\in F\otimes W$.
Recall that $X_{1}=\sum_{y\in Y_{0}}\operatorname{span}\{y_{0}\mid\delta_{V}(y)=y_{-1}\otimes y_{0}\},$ We are going to show $g(F^{\perp}\otimes X_{1})=0$. This is equivalent to verify $f(X_{1})\in F\otimes W$. Since $f$ is a morphism in $\mathcal{C}$, we have
$(\operatorname{id}\otimes f)\delta(h\rhd v)$ $=\delta(f(h\rhd v)).$
Recall $F$ is an object in $^{H}_{H}\mathcal{YD}$}, we know that $\delta(f(h\rhd v))\in\delta(F\otimes W)\subseteq H\otimes(F\otimes W)$. This implies $f((h\rhd v)_{0})\in F\otimes W.$ Hence $g(F^{\perp}\otimes X_{1})=0$. By iterated statements, we can prove that
$g(F^{\perp}\otimes\langle V\rangle)=0.$
Note that we have such a non-degenerate pairing
$A/{F^{\perp}}\otimes F\rightarrow\mathbbm{k},\ \overline{a}\otimes b\mapsto\omega(a,b).$
Since $F$ is finite-dimensional, by the situation considered before, we have a bijective map
$\operatorname{Hom}_{\mathcal{C}}(V,F\otimes W)\cong\operatorname{Hom}_{\mathcal{C},\operatorname{rat}}(A/F^{\perp}\otimes V,W).$
Let $\overline{g}\in\operatorname{Hom}_{\mathcal{C},\operatorname{rat}}(A/F^{\perp}\otimes V,W)$ be the morphism induced by $g$ and $f\in\operatorname{Hom}(V,F\otimes W)$ be the preimage of $\overline{g}$. Then the preimage of $g$ is
$i\circ f:V\rightarrow B\otimes W,$
where $i:F\otimes W\rightarrow B\otimes W$ is an inclusion. Thus $G_{V}:\operatorname{Hom}_{\mathcal{C}}(V,B\otimes W)\cong\operatorname{Hom}_{\mathcal{C},\operatorname{rat}}(A\otimes V,W)$ is bijective in this case.
Step4:
It remains to prove the bijectivity for an arbitrary object $V\in\mathcal{C}$. Let
$\mathcal{V}=\{\langle V^{\prime}\rangle\mid V^{\prime}\subseteq V\ \operatorname{finite-dimensional}\ H-\operatorname{subcomodule}\}.$
By the finiteness theorem for comodules, we have $V=\bigcup\limits_{V^{\prime}\ \operatorname{finite-dimensional}}V^{\prime}$ as comodules. Note that for all $U_{1},U_{2}\in\mathcal{V}$, there is an object $U\in\mathcal{V}$ with $U_{1}\cup U_{2}\subseteq U$, since $U$ is closed under direct sums. Thus
$V=\bigcup_{U\in\mathcal{V}}U$
as an object in $\mathcal{C}$. Given $g\in\operatorname{Hom}_{\mathcal{C},\operatorname{rat}}(A\otimes V,W),$ let $g_{U}$ be the restriction of $g$ to $A\otimes U$, then there is a morphism $f_{U}:U\rightarrow B\otimes W$ with $G_{V}(f_{U})=g_{U}$. For $U_{1},U_{2}\in\mathcal{V}$, let $U_{1}\subseteq U_{2}$, we have $f_{U_{2}}\mid_{U_{1}}=f_{U_{1}}$, since $G_{V}$ is injective. Hence the maps $f_{U}$ deine a linear map $f:V\rightarrow B\otimes W$ by $f(v)=f_{U}(v)$, where $U$ is an element in $\mathcal{V}$ containing $v$. This construction yields $G_{V}(f)=g$. Thus $G_{V}$ is bijective for arbitrary $V$. ∎
3.3 Bijection between comodules and rational modules
We begin by introducing the notion of Hopf pairings. For our purpose, we assume that $A$ and $B$ are locally finite $\mathbb{N}_{0}$-graded Hopf algebras in $\mathcal{C}$ in this section. Since this does not affect our study of Nichols algebras, we assume that $A$ and $B$ have bijective antipodes in $\mathcal{C}$.
Definition 3.7.
A morphism $\omega:A\otimes B\to\mathbbm{k}$ in $\mathcal{C}$ is called a Hopf pairing, if the following identities hold:
$\omega\circ(\mu_{A}\otimes\operatorname{id}_{B})$ $=\omega\circ(\operatorname{id}_{A}\otimes\omega\otimes\operatorname{id}_{B})\circ(\operatorname{id}_{A}\otimes a^{-1}_{A,B,B})\circ a_{A,A,B\otimes B}\circ(\operatorname{id}_{A\otimes A}\otimes\Delta_{B}),$ (3.5)
$\omega\circ(\eta_{A}\otimes\operatorname{id}_{B})$ $=\varepsilon_{B},$ (3.6)
$\omega\circ(\operatorname{id}_{A}\otimes\mu_{B})$ $=\omega\circ(\operatorname{id}_{A}\otimes\omega\otimes\operatorname{id}_{B})\circ(\operatorname{id}_{A}\otimes a^{-1}_{A,B,B})\circ a_{A,A,B\otimes B}\circ(\Delta_{A}\otimes\operatorname{id}_{B\otimes B}),$ (3.7)
$\omega\circ(\operatorname{id}_{A}\otimes\eta_{B})$ $=\varepsilon_{A}.$ (3.8)
Moreover, we call a Hopf pairing $\omega:A\otimes B\to\mathbbm{k}$ is a dual pair of locally finite $\mathbb{N}_{0}$-graded Hopf algebras in $\mathcal{C}$, if $\omega$ is non-degenerate and
$\omega(A(n),B(m))=0,\ \operatorname{if}\ n\neq m.$
There are some useful properties of Hopf pairings, we list here for further use.
Remark 3.8.
[Heckenberger and Schneider, [2020] ©2020, Proposition 3.3.8] Under above assumptions, we have the following properties.
  • (1)
    Suppose $\omega:A\otimes B\rightarrow\textbf{1}$ is a dual pair in $\mathcal{C}$, then
    $\omega\circ(\operatorname{id}\otimes\mathcal{S}_{B})=\omega\circ(\mathcal{S}_{A}\otimes\operatorname{id}).$ (3.9)
  • (2)
    If $\omega:A\otimes B\rightarrow\textbf{1}$ is a dual pair, then also
    $\omega^{+}:=\omega\circ c_{B,A}\circ(\mathcal{S}_{B}\otimes\mathcal{S}_{A}):B\otimes A\rightarrow\mathbbm{k},$ (3.10)
    $\omega^{+\operatorname{cop}}:=\omega^{+}\circ(\operatorname{id}_{B}\otimes\mathcal{S}^{-1}_{A}):B^{\operatorname{cop}}\otimes A^{\operatorname{cop}}\rightarrow\mathbbm{k}.$ (3.11)
The following lemma holds in general boards of monoidal category, here we cite this lemma for simplicity.
Lemma 3.9.
[Heckenberger and Schneider, [2020] ©2020, Proposition 3.3.9] Let $\mathcal{C}$ be an arbitrary braided monoidal category. Suppose $A$, $B$ are Hopf algebras in $\mathcal{C}$. Assume that $p:A\otimes B\rightarrow\textbf{1}$ is a Hopf pairing, then we have two strict monoidal functors:
$D_{1}:$ ${{}^{B}\mathcal{C}}\longrightarrow{{}_{A^{\operatorname{cop}}}\overline{\mathcal{C}}},\ \ (V,\delta)\mapsto(V,\overline{\lambda}),$ (3.12)
with $\overline{\lambda}=(A\otimes V\xrightarrow{\operatorname{id}_{A}\otimes\delta}A\otimes(B\otimes V)\xrightarrow{a^{-1}_{A,B,V}}(A\otimes B)\otimes V\xrightarrow{p\otimes\operatorname{id}_{V}}V)$
and
$D_{2}:$ ${{}^{A^{\operatorname{cop}}}\overline{\mathcal{C}}}\longrightarrow{{}_{B}\mathcal{C}},\ \ (V,\overline{\delta})\mapsto(V,\lambda),$ (3.13)
with $\lambda=(B\otimes V\xrightarrow{\operatorname{id}_{B}\otimes\overline{\delta}}B\otimes(A\otimes V)\xrightarrow{a^{-1}_{B,A,V}}(B\otimes A)\otimes V\xrightarrow{p^{+\operatorname{cop}}\otimes\operatorname{id}_{V}}V).$
The main result of this subsection is to give a bijection between $B$-comodules in $\mathcal{C}$ and rational $A^{\operatorname{cop}}$-modules in $\overline{\mathcal{C}}$.
Proposition 3.10.
Let $(A,B,\omega)$ be a dual pair of locally finite $\mathbb{N}_{0}$-graded Hopf algberas in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$. The functor
$D_{1}:\$${{}^{B}\mathcal{C}}\longrightarrow{{}_{A^{\operatorname{cop}}}\overline{\mathcal{C}}_{\operatorname{rat}}},\ \ (V,\delta)\mapsto(V,\lambda),$
with $\lambda=(A\otimes V\xrightarrow{\operatorname{id}_{A}\otimes\delta}A\otimes(B\otimes V)\xrightarrow{a^{-1}_{A,B,V}}(A\otimes B)\otimes V\xrightarrow{\omega\otimes\operatorname{id}_{V}}V)$
is a strict monoidal isomorphism of categories.
Proof.
Let $V$ be an object in $^{H}_{H}\mathcal{YD}$}. By Proposition 3.6, we have a bijection:
$G_{V}:\operatorname{Hom}_{C}(V,B\otimes V)\to\operatorname{Hom}_{C,\text{rat}}(A\otimes V,V),$
$f\mapsto\lambda=(A\otimes V\xrightarrow{\operatorname{id}_{A}\otimes f}A\otimes(B\otimes V)\xrightarrow{a^{-1}_{A,B,V}}(A\otimes B)\otimes V\xrightarrow{\omega\otimes\operatorname{id}}V).$
Note that $\operatorname{Hom}_{C,\text{rat}}(A\otimes V,V)$ is the set of all $\lambda$ in $\operatorname{Hom}_{C}(A\otimes V,V)$ such that for all $v\in V$, there is a finite-dimensional object $F\subseteq B$, such that $\lambda(F^{\perp}\otimes v)=0.$ Since $B$ is locally finite, there is a natural number $n_{0}$ such that $F\subseteq\bigoplus_{i=0}^{n_{0}}B(i)$. Therefore $\lambda(A(n)\otimes v)=0$ for all $n>n_{0}$. Thus if $(V,\lambda)\in{{}_{A}\mathcal{C}}$, then $(V,\lambda)\in{{}_{A}\mathcal{C}_{\operatorname{rat}}}$ automatically.
Now we are going to prove $(V,\lambda)\in{{}_{A}\mathcal{C}}_{\operatorname{rat}}$ if and only if $(V,\delta)\in{{}^{B}\mathcal{C}}$. The condition $(V,\delta)\in{{}^{B}\mathcal{C}}$ implies $(V,\lambda)\in{{}_{A}\mathcal{C}}_{\operatorname{rat}}$ follows from Lemma 3.9 and the above statement. On the other hand, $(V,\lambda)\in{{}_{A}\mathcal{C}}$ is equivalent to
$\lambda\circ(\operatorname{id}_{A}\otimes\lambda)=\lambda\circ(\mu_{A}\otimes\operatorname{id}_{V})\circ a_{A,A,V}^{-1}.$
According to the definition of $\lambda$,
$\lambda\circ(\operatorname{id}_{A}\otimes\lambda)=(\omega\otimes\operatorname{id}_{V})\circ a^{-1}_{A,B,V}\circ(\operatorname{id}_{A}\otimes\delta)\circ(\operatorname{id}_{A}\otimes(\omega\otimes\operatorname{id}_{V}))\circ(\operatorname{id}_{A}\otimes a^{-1}_{A,B,V})\circ(\operatorname{id}_{A}\otimes(\operatorname{id}_{A}\otimes\delta)).$
Since $\delta$, $\omega$ are morphisms in $\mathcal{C}$, and the associative isomorphism is natural, we have
$\lambda\circ(\operatorname{id}_{A}\otimes\lambda)$ $=(\omega\otimes\operatorname{id}_{V})\circ a^{-1}_{A,B,V}\circ(\operatorname{id}_{A}\otimes(\omega\otimes\operatorname{id}_{B\otimes V}))\circ(\operatorname{id}_{A}\otimes a^{-1}_{A,B,B\otimes V})\circ(\operatorname{id}_{A}\otimes(\operatorname{id}_{A}\otimes(\operatorname{id}_{B}\otimes\delta)\circ\delta)).$ (3.14)
$=(\omega\otimes\operatorname{id}_{V})\circ((\operatorname{id}_{A}\otimes(\omega\otimes\operatorname{id}_{B}))\otimes\operatorname{id}_{V})\circ(a^{-1}_{A,A\otimes B,B}\otimes\operatorname{id}_{V})\circ(\operatorname{id}_{A}\otimes a^{-1}_{A\otimes B,B,V})$
$\circ(\operatorname{id}_{A}\circ a^{-1}_{A,B,B\otimes V})\circ(\operatorname{id}_{A}\otimes(\operatorname{id}_{A}\otimes(\operatorname{id}_{B}\otimes\delta)\circ\delta)).$
Now we write down explicit formula of $\lambda\circ(\mu_{A}\otimes\operatorname{id}_{V})\circ a_{A,A,V}^{-1}$.
$\lambda\circ(\mu_{A}\otimes\operatorname{id}_{V})\circ a_{A,A,V}^{-1}=(\omega\otimes\operatorname{id}_{V})\circ a^{-1}_{A,B,V}\circ(\operatorname{id}\otimes\delta)\circ(\mu_{A}\otimes\operatorname{id})\circ a^{-1}_{A,A,V}.$
Since $\mu_{A}$ is a morphism in $\mathcal{C}$, we have
$\lambda\circ(\mu_{A}\otimes\operatorname{id}_{V})\circ a_{A,A,V}^{-1}=(\omega\otimes\operatorname{id}_{V})\circ((\mu_{A}\otimes\operatorname{id}_{B})\otimes\operatorname{id}_{V})\circ a^{-1}_{A\otimes A,B,V}\otimes(\operatorname{id}_{A\otimes A}\otimes\delta)\circ a^{-1}_{A,A,V}.$
By equation (3.5), and by naturality of associative isomorphism,
$\lambda\circ(\mu_{A}\otimes\operatorname{id}_{V})\circ a_{A,A,V}^{-1}$ (3.15)
$=(\omega\otimes\operatorname{id}_{V})\circ((\operatorname{id}_{A}\otimes(\omega\otimes\operatorname{id}_{B}))\otimes\operatorname{id}_{V})\circ((\operatorname{id}_{A}\otimes a^{-1}_{A,B,B})\circ a_{A,A,B\otimes B}\circ(\operatorname{id}_{A\otimes A}\otimes\Delta_{B})\otimes\operatorname{id}_{V})$
$\circ a^{-1}_{A\otimes A,B,V}\circ(\operatorname{id}_{A\otimes A}\otimes\delta)\circ a^{-1}_{A,A,V}$
$=(\omega\otimes\operatorname{id}_{V})\circ((\operatorname{id}_{A}\otimes(\omega\otimes\operatorname{id}_{B}))\otimes\operatorname{id}_{V})\circ((\operatorname{id}_{A}\otimes a^{-1}_{A,B,B})\circ a_{A,A,B\otimes B}\otimes\operatorname{id}_{V})$
$\circ a^{-1}_{A\otimes A,B\otimes B,V}\circ a^{-1}_{A,A,(B\otimes B)\otimes V}\circ(\operatorname{id}_{A}\otimes(\operatorname{id}_{A}\otimes a_{B,B,V}^{-1}))\circ(\operatorname{id}_{A}\otimes(\operatorname{id}_{A}\otimes(a_{B,B,V}\circ((\Delta_{B}\otimes\operatorname{id}_{V})\circ\delta)))).$
For all $a,a^{\prime}\in A$, $v\in V$, we have $\lambda\circ(\operatorname{id}_{A}\otimes\lambda)(a\otimes(a^{\prime}\otimes v))=\lambda\circ(\mu_{A}\otimes\operatorname{id}_{V})\circ a_{A,A,V}^{-1}(a\otimes(a^{\prime}\otimes v)).$ By non-degeneracy of $\omega$ and Mac-Lane’s coherence theorem, this is equivalent to
$(\operatorname{id}\otimes\delta)\circ\delta(v)=a_{B,B,V}\circ(\Delta_{B}\otimes\operatorname{id}_{V})\circ\delta(v).$
Furthermore, we have $v=(\varepsilon_{B}\otimes\operatorname{id}_{V})\circ\delta(v)$ by equation (3.6). Therefore we deduce that $(V,\lambda)\in{{}_{A}\mathcal{C}_{\operatorname{rat}}}$ if and only if $(V,\delta)\in{{}^{B}\mathcal{C}}$.
Now let $(V,\delta),(V^{\prime},\delta^{\prime})\in{{}^{B}\mathcal{C}}$, and $(V,\lambda)$, $(V^{\prime},\lambda^{\prime})$ be the corresponding modules in $_{A}\mathcal{C}_{\operatorname{rat}}$}, where $\lambda=G_{V}(\delta)$, $\lambda^{\prime}=G_{V}^{\prime}(\delta^{\prime})$. It is direct that a map $f\in\operatorname{Hom}_{\mathcal{C}}(V,V^{\prime})$ is a $B$-comodule map if and only if $f$ is a $A$-module map in $\mathcal{C}$.
Note that $(V,\lambda)\in{{}_{A}\mathcal{C}_{\operatorname{rat}}}$ if and only if $(V,\lambda)\in{{}_{A^{\operatorname{cop}}}\overline{\mathcal{C}}_{\operatorname{rat}}}$. Since the category $_{A^{\operatorname{cop}}}\overline{\mathcal{C}}_{\operatorname{rat}}$} is a monoidal subcategory of $_{A^{\operatorname{cop}}}\overline{\mathcal{C}}$} by Lemma 3.4 and $D_{1}$ is a strict monoidal functor by Lemma 3.9, we deduce that $D_{1}$ is a strict monoidal equivalence. Since $G_{V}$ is bijective, it is easy to see that the inverse functor is given by
$D_{1}^{\prime}:{{}_{A^{\operatorname{cop}}}\overline{\mathcal{C}}_{\operatorname{rat}}}\rightarrow{{}^{B}\mathcal{C}},\ \ (V,\lambda^{\prime})\mapsto(V,\delta^{\prime}),$
$\text{where}\ \delta^{\prime}\ \text{is determined by}\ G_{V}^{-1}(\lambda^{\prime}).$ Thus $D_{1}$ is a strict monoidal isomorphism. ∎
3.4 Monoidal equivalence between Yetter-Drinfeld module categories related by a dual pair
Let $(A,B,\omega)$ be a dual pair of locally finite $\mathbb{N}_{0}$-graded Hopf algberas with bijective antipodes in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$. The main result in this subsection is to show there is a braided monoidal equivalence $(\Omega,\beta):{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}\rightarrow{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$.
By Remark 3.8, we have a dual pair of locally finite $\mathbb{N}_{0}$-graded Hopf algebras in $\overline{\mathcal{C}}$:
$\omega^{+\operatorname{cop}}$ $:B^{\operatorname{cop}}\otimes A^{\operatorname{cop}}\longrightarrow\mathbbm{k},$
$\omega^{+\operatorname{cop}}$ $=\omega\circ c_{B,A}\circ(\mathcal{S}_{B}\otimes\mathcal{S}_{A})\circ(\operatorname{id}_{B}\otimes\mathcal{S}^{-1}_{A})$
$=\omega\circ c_{B,A}\circ(\mathcal{S}_{B}\otimes\operatorname{id}_{A})$
$=\omega\circ c_{B,A}\circ(\operatorname{id}_{B}\otimes\mathcal{S}_{A}).$
Recall that by Proposition 3.10, we now have two monoidal isomorphisms:
$D_{1}:{{}^{B}\mathcal{C}}\rightarrow{{}_{A^{\operatorname{cop}}}\overline{\mathcal{C}}_{\operatorname{rat}}},\ D_{2}:{{}^{A^{\operatorname{cop}}}\overline{\mathcal{C}}}\rightarrow{{}_{B}\mathcal{C}_{\operatorname{rat}}},$ (3.16)
$D_{1}(V,\delta)=(V,\overline{\lambda}),\ \overline{\lambda}=(A\otimes V\xrightarrow{\operatorname{id}_{A}\otimes\delta}A\otimes(B\otimes V)\xrightarrow{a^{-1}_{A,B,V}}(A\otimes B)\otimes V\xrightarrow{\omega\otimes\operatorname{id}_{V}}V),$ (3.17)
$D_{2}(V,\overline{\delta})=(V,{\lambda}),\ {\lambda}=(B\otimes V\xrightarrow{\operatorname{id}_{B}\otimes\overline{\delta}}B\otimes(A\otimes V)\xrightarrow{a^{-1}_{B,A,V}}(A\otimes B)\otimes V\xrightarrow{\omega^{+\operatorname{cop}}\otimes\operatorname{id}_{V}}V).$ (3.18)
Our first objective is to prove the following braided monoidal equivalence:
Theorem 3.11.
The functor
$\Gamma:\overline{{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}\longrightarrow{{}^{A^{\operatorname{cop}}}_{A^{\operatorname{cop}}}\mathcal{YD}(\overline{\mathcal{C}})}_{\operatorname{rat}},\ (V,\lambda,\delta)\mapsto(V,\overline{\lambda},\overline{\delta}),$ (3.19)
where $\overline{\lambda},\overline{\delta}$ are defined by (3.17), (3.18), and the morphism $f$ are mapped onto $f$, is an equivalence of braided monoidal categories.
The following lemma, which holds in the general setting of a braided monoidal category $\mathcal{C}$ , is used to establish this theorem. Without loss of generality, we assume $\mathcal{C}$ is strict. Let $A$, $B$ be Hopf algebras in $\mathcal{C}$, and $p:A\otimes B\rightarrow\textbf{1}$ is a Hopf pairing, now we have two braided monoidal categories $\overline{{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}$ and ${{}^{A^{\operatorname{cop}}}_{A^{\operatorname{cop}}}\mathcal{YD}(\overline{\mathcal{C}})}_{\operatorname{rat}}$, with braiding $\overline{c}^{\mathcal{YD}(\mathcal{C})}$ and $c^{\mathcal{YD}(\overline{\mathcal{C}})}$ respectively.
Lemma 3.12.
With above assumptions, let $(X,\overline{\delta}_{X})\in{{}^{A^{\operatorname{cop}}}\overline{\mathcal{C}}}$, $(V,\delta)\in{{}^{B}\mathcal{C}}$, and define
$(X,\lambda_{X})=D_{2}(X,\overline{\delta}_{X}),\ (V,\overline{\lambda})=D_{1}(V,\delta).$
where $D_{1},D_{2}$ is given by Lemma 3.9, then we have
$\overline{c}^{\mathcal{YD}(\mathcal{C})}_{(X,\lambda_{X}),(V,\delta)}=c^{\mathcal{YD}(\overline{\mathcal{C}})}_{(X,\overline{\delta}_{X}),(V,\overline{\lambda})}.$ (3.20)
Proof.
By definition of $c^{\mathcal{YD}}$, properties of Hopf pairing and naturality of braiding, we have
$\overline{c}^{\mathcal{YD}(\mathcal{C})}_{(X,\lambda_{X}),(V,\delta)}$ $=\overline{c}_{X,V}\circ((p^{+\operatorname{cop}}\otimes\operatorname{id}_{X})\circ(\operatorname{id}_{B}\otimes\overline{\delta}_{X})\otimes\operatorname{id}_{V})\circ(\mathcal{S}_{B}^{-1}\otimes\operatorname{id}_{X}\otimes\operatorname{id}_{V})\circ(\overline{c}_{X,B}\otimes\operatorname{id}_{V})\circ(\operatorname{id}_{X}\otimes\delta)$
$=\overline{c}_{X,V}\circ(p\circ c_{B,A}\otimes\operatorname{id}_{X}\otimes\operatorname{id}_{V})\circ(\operatorname{id}_{B}\otimes\overline{\delta}_{X}\otimes\operatorname{id}_{V})\circ(\overline{c}_{X,B}\otimes\operatorname{id}_{V})\circ(\operatorname{id}_{X}\otimes\delta)$
$=(p\otimes\overline{c}_{X,V})\circ(c_{B,A}\otimes\operatorname{id}_{X})\circ(\overline{c}_{A\otimes X,B}\otimes\operatorname{id}_{V})\circ(\operatorname{id}_{A}\otimes\operatorname{id}_{X}\otimes\delta)\circ(\overline{\delta}_{X}\otimes\operatorname{id}_{V})$
$=(p\otimes\overline{c}_{X,V})\circ(\operatorname{id}_{A}\otimes\overline{c}_{X,B}\otimes\operatorname{id}_{V})\circ(\operatorname{id}_{A}\otimes\operatorname{id}_{X}\otimes\delta)\circ(\overline{\delta}_{X}\otimes\operatorname{id}_{V})$
$=(p\otimes\overline{c}_{X,B\otimes V})\circ(\operatorname{id}_{A}\otimes\operatorname{id}_{X}\otimes\delta)\circ(\overline{\delta}_{X}\otimes\operatorname{id}_{V})$
$=((p\otimes\operatorname{id}_{V})\circ(\operatorname{id}_{A}\otimes\delta)\otimes\operatorname{id}_{X})\circ(\operatorname{id}_{A}\otimes\overline{c}_{X,V})\circ(\overline{\delta}_{X}\otimes\operatorname{id}_{V})=c^{\mathcal{YD}(\overline{\mathcal{C}})}_{(X,\overline{\delta}_{X}),(V,\overline{\lambda})}.$
Here the second equality uses (3.9) and (3.11), the third and the sixth equalities use naturality of braiding, the fourth and fifth equalities use braiding axiom. ∎
Now we return to the setting of coquasi-Hopf algebras. We want to restrict the Yetter-Drinfeld module criterion to rational modules. Let $B$ be a locally finite $\mathbb{N}_{0}$-graded Hopf algbera in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$ with bijective antipode. Let $(V,\lambda)\in{{}_{B}\mathcal{C}_{\operatorname{rat}}}$, and $(V,\delta)\in{{}^{B}\mathcal{C}}$.
Lemma 3.13.
Let $X\in{{}_{B}\mathcal{C}}$, and assume that there is an index set $I$, a family $X_{i},i\in I$, of objects in $_{B}\mathcal{C}$, and morphisms $\pi_{i}:X\rightarrow X_{i}$ in $_{B}\mathcal{C}$ for all $i\in I$ with $\bigcap_{i\in I}\operatorname{ker}(\pi_{i})=0$. Then if $c^{\mathcal{YD}}_{V,X_{i}}$ is a morphism in $_{B}\mathcal{C}$ for all $i\in I$, then $c^{\mathcal{YD}}_{V,X}$ is a morphism in $_{B}\mathcal{C}$.
Proof.
The proof of this lemma is parallel to [Heckenberger and Schneider, [2020] ©2020, Lemma 12.2.5]. We omit here for simplicity. ∎
Lemma 3.14.
The following are equivalent.
(1) For all $(X,\lambda_{X})\in{{}_{B}\mathcal{C}_{\operatorname{rat}}}$, $c^{\mathcal{YD}}_{V,X}$ is a morphism in $_{B}\mathcal{C}_{\operatorname{rat}}$.
(2) The object $(V,\lambda,\delta)$ belongs to ${{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$.
Proof.
Suppose for all $(X,\lambda_{X})\in{{}_{B}\mathcal{C}_{\operatorname{rat}}}$, the map $c^{\mathcal{YD}}_{V,X}$ is a morphism in $_{B}\mathcal{C}_{\operatorname{rat}}$. Then by Remark 2.12, we only need to prove that
$c^{\mathcal{YD}}_{V,B}:V\otimes B\rightarrow B\otimes V$
is a morphism in $_{B}\mathcal{C}$. Now for all $n\geq 0$, let $X_{n}=B/{\bigoplus_{i\geq n}B(i)}$. The quotient map $\pi_{n}:B\rightarrow X_{n}$ is a morphism in $_{B}\mathcal{C}$. We have
$\bigcap_{n\geq 0}\operatorname{ker}(\pi_{n})=0,\ \ \mu_{B}(B(m)\otimes X_{n})=0$
for all $m\geq n$. Since $X_{n}$ is a rational $B$-module in $\mathcal{C}$, $c^{\mathcal{YD}}_{V,X_{n}}$ is a morphism in $_{B}\mathcal{C}$. By Lemma 3.13, $c^{\mathcal{YD}}_{V,B}$ is a morphism in $_{B}\mathcal{C}$. Hence $(V,\lambda,\delta)\in{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$.
Suppose $(V,\lambda,\delta)\in{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$, we deduce that $c^{\mathcal{YD}}_{V,X}$ is a morphism in $_{B}\mathcal{C}$ for any $(X,\lambda_{X})\in{{}_{B}\mathcal{C}_{\operatorname{rat}}}$ by Remark 2.12. Furthermore, $_{B}\mathcal{C}_{\operatorname{rat}}$} is a monoidal full subcategory of $_{B}\mathcal{C}$, and $V\otimes X,X\otimes V\in{{}_{B}\mathcal{C}_{\operatorname{rat}}}$. Thus $c^{\mathcal{YD}}_{V,X}$ is a morphism in $_{B}\mathcal{C}_{\operatorname{rat}}$. ∎
Proposition 3.15.
The following are equivalent.
(1) The object $(V,\lambda,\delta)$ belongs to ${{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$.
(2) The object $(V,\overline{\lambda},\overline{\delta})$ belongs to ${{}^{A^{\operatorname{cop}}}_{A^{\operatorname{cop}}}\mathcal{YD}(\overline{\mathcal{C}})}_{\operatorname{rat}}$.
(3) For all $(X,\lambda_{X})\in{{}_{B}\mathcal{C}_{\operatorname{rat}}}$, the following map is a morphism in $_{B}\mathcal{C}_{\operatorname{rat}}$
$\overline{c}^{\mathcal{YD}(\mathcal{C})}_{(X,\lambda_{X}),(V,\delta)}:(X,\lambda_{X})\otimes(V,\lambda)\rightarrow(V,\lambda)\otimes(X,\lambda_{X}).$
(4) For all $(X,\overline{\delta}_{X})\in{{}^{A^{\operatorname{cop}}}\overline{\mathcal{C}}}$, the following map is a morphism in $^{A^{\operatorname{cop}}}\overline{\mathcal{C}}$.
$c^{\mathcal{YD}(\overline{\mathcal{C}})}_{(X,\overline{\delta}_{X}),(V,\overline{\lambda})}:(X,\overline{\delta}_{X})\otimes(V,\overline{\delta})\rightarrow(V,\overline{\delta})\otimes(X,\overline{\delta}_{X}).$
Proof.
(1) $\Leftrightarrow$ (3): The condition (1), by Lemma 3.6, is equivalent to the condition that for all $(X,\lambda_{X})\in{{}_{B}\mathcal{C}_{\operatorname{rat}}}$, $c^{\mathcal{YD}}_{(V,\delta),(X,\lambda_{X})}$ is a morphism in $_{B}\mathcal{C}_{\operatorname{rat}}$. Since we have assumed $B$ has bijective antipode, $c^{\mathcal{YD}}_{(V,\delta),(X,\lambda_{X})}$ is an isomorphism with inverse $\overline{c}^{\mathcal{YD}}_{(X,\lambda_{X}),(V,\delta)}.$ Hence (1) is equivalent to (3).
(2) $\Leftrightarrow$ (4): This follows from Remark 2.12.
(3) $\Leftrightarrow$ (4): By Lemma 3.12,
$\overline{c}^{\mathcal{YD}(\mathcal{C})}_{(X,\lambda_{X}),(V,\delta)}=c^{\mathcal{YD}(\overline{\mathcal{C}})}_{(X,\overline{\delta}_{X}),(V,\overline{\lambda})}$
as a morphism in $\mathcal{C}$. Then the equivalence of (3) and (4) follows from $D_{2}:{{}^{A^{\operatorname{cop}}}\overline{\mathcal{C}}}\rightarrow{{}_{B}\mathcal{C}_{\operatorname{rat}}}$ is a monoidal isomorphism. ∎
Proof of Theorem 3.11.
Let $V\in\mathcal{C}$, and
$S_{1}=\{(\lambda,\delta)\mid(V,\lambda)\in{{}_{B}\mathcal{C}_{\operatorname{rat}}},(V,\delta)\in{{}^{B}\mathcal{C}}\}$
$S_{2}=\{(\overline{\lambda},\overline{\delta})\mid(V,\overline{\lambda})\in{{}_{A^{\operatorname{cop}}}\overline{\mathcal{C}}_{\operatorname{rat}}},(V,\overline{\delta})\in{{}^{A^{\operatorname{cop}}}\overline{\mathcal{C}}}\}$
We define a map $\phi:S_{1}\rightarrow S_{2}$, by $(\lambda,\delta)\mapsto(\overline{\lambda},\overline{\delta})$, where $\overline{\lambda},\overline{\delta}$ are given by
$D_{1}(V,\delta)=(V,\overline{\lambda}),\ \ \ D_{2}(V,\overline{\delta})=(V,\lambda).$
Thus $\phi$ is bijective since $D_{1}$, $D_{2}$ are monoidal isomorphisms.
Now that given $(V,\lambda,\delta)\in{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$, the object $(V,\overline{\lambda},\overline{\delta})\in{{}^{A^{\operatorname{cop}}}_{A^{\operatorname{cop}}}\mathcal{YD}(\overline{\mathcal{C}})}_{\operatorname{rat}}$ by Proposition 3.15. Since $\phi$ is bijective, we deduce that $\Gamma$ is a monoidal equivalence. It is strict since $D_{1}$ and $D_{2}$ is strict. To show $\Gamma$ is braided, let $X=(X,\lambda_{X},\delta_{X})\in{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$, and $\Gamma(X)=(X,\overline{\lambda}_{X},\overline{\delta}_{X})\in{{}^{A^{\operatorname{cop}}}_{A^{\operatorname{cop}}}\mathcal{YD}(\overline{\mathcal{C}})}_{\operatorname{rat}}$. We know that $\overline{c}^{\mathcal{YD}(\mathcal{C})}_{(X,\lambda_{X}),(V,\delta)}$ is the braiding in $\overline{{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}$, and $c^{\mathcal{YD}(\overline{\mathcal{C}})}_{(X,\overline{\delta}_{X}),(V,\overline{\lambda})}$ is the braiding of $\Gamma(X),\Gamma(V)$ in ${{}^{A^{\operatorname{cop}}}_{A^{\operatorname{cop}}}\mathcal{YD}(\overline{\mathcal{C}})}_{\operatorname{rat}}$. By Lemma 3.12, the functor $\Gamma$ is braided. Hence $\Gamma$ is a strict braided monoidal equivalence.
Our next goal is to build a monoidal equivalence between ${{}^{A^{\operatorname{cop}}}_{A^{\operatorname{cop}}}\mathcal{YD}(\overline{\mathcal{C}})}_{\operatorname{rat}}$ and $\overline{{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}}_{\operatorname{rat}}$. It is standard that
$_{A}\mathcal{C}\rightarrow{\mathcal{C}_{A}},\ (V,\lambda)\mapsto(V,\lambda_{+}),\ \text{where}\ \lambda_{+}=\lambda\circ c_{V,A}\circ(\operatorname{id}_{V}\otimes\mathcal{S}_{A}),$
$\mathcal{C}_{A}\rightarrow{{}_{A}\mathcal{C}},\ (V,\lambda)\mapsto(V,\lambda_{-}),\ \text{where}\ \lambda_{-}=\lambda\circ\overline{c}_{A,V}\circ(\mathcal{S}_{A}^{-1}\otimes\operatorname{id}_{V}),$
are inverse isomorphisms of monoidal categories. Furthermore, suppose $(V,\lambda)\in{{}_{A}\mathcal{C}_{\operatorname{rat}}}$, for $v\in V$, $c_{V,A}(v\otimes a)$ is a finite sum of elements in $A\otimes V$. By $\mathcal{S}_{A}$ is a $\mathbb{N}_{0}$-graded map, we have $(V,\lambda_{+})\in{\mathcal{C}_{A}}_{,\operatorname{rat}}$. Similarly, suppose $(V,\lambda)\in{\mathcal{C}_{A}}_{,\operatorname{rat}}$, we can deduce that $(V,\lambda_{-})\in{{}_{A}\mathcal{C}_{\operatorname{rat}}}$.
Now we restrict the monoidal isomorphisms in [Heckenberger and Schneider, [2020] ©2020, Theorem 3.4.15] to the rational Yetter-Drinfeld modules, we have a braided monoidal equivalence:
$(F_{rl},\alpha):{{\mathcal{YD}(\mathcal{C})^{A}_{A}}}_{,\operatorname{rat}}\rightarrow{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}},\$$(V,\lambda,\delta)\mapsto(V,\lambda_{-},(\mathcal{S}_{A}\otimes\operatorname{id}_{V})\circ c_{V,A}\circ\delta),$ (3.21)
$\alpha_{X,Y}=c_{Y,X}^{{\mathcal{YD}(\mathcal{C})^{A}_{A}}}\circ\overline{c}_{X,Y},$ (3.22)
where morphisms $f$ are mapped onto $f$. From [Heckenberger and Schneider, [2020] ©2020, Theorem 3.4.16], we have a braided strict monoidal equivalence:
$F_{lr}:{{}^{A^{\operatorname{cop}}}_{A^{\operatorname{cop}}}\mathcal{YD}(\overline{\mathcal{C}})}_{\operatorname{rat}}\rightarrow\overline{{\mathcal{YD}(\mathcal{C})^{A}_{A}}_{,\operatorname{rat}}},\ (V,\lambda,\delta)\mapsto(V,\lambda_{+},c_{A,V}\circ\delta),$ (3.23)
where morphisms $f$ are mapped onto $f$.
Our main result in this section establishes the braided monoidal equivalence between the categories of rational modules over dual pairs. The construction of the functor $\Omega$ relies on the interplay between the equivalences established in the previous subsections, as illustrated in the following diagram:
$\overline{{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}$${{}^{A^{\operatorname{cop}}}_{A^{\operatorname{cop}}}\mathcal{YD}(\overline{\mathcal{C}})}_{\operatorname{rat}}$$\overline{{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}$$\overline{{\mathcal{YD}(\mathcal{C})^{A}_{A}}_{,\operatorname{rat}}}$$\scriptstyle{\Gamma}\scriptstyle{\cong}$$\scriptstyle{\Omega}$$\scriptstyle{F_{lr}}$$\scriptstyle{(F_{rl},\alpha)}$
Theorem 3.16.
The following functor is a braided monoidal equivalence:
$(\Omega,\beta):{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}\rightarrow{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}},\ \text{where}\ (V,\lambda_{B},\delta_{B})\mapsto(V,\lambda_{A},\delta_{A}),\text{with}$
$\lambda_{A}=(A\otimes V\xrightarrow{\operatorname{id}\otimes\delta_{B}}A\otimes(B\otimes V)\xrightarrow{a^{-1}_{A,B,V}}(A\otimes B)\otimes V\xrightarrow{\omega\otimes\operatorname{id}}V),$
$\delta_{A}=(V\xrightarrow{\delta^{\prime}}A\otimes V\xrightarrow{\mathcal{S}_{A}^{2}\otimes\operatorname{id}_{V}}A\otimes V\xrightarrow{c^{2}_{A,V}}A\otimes V),\ \text{ where }\ \delta^{\prime}\ \text{is defined by}$
$\lambda_{B}=(B\otimes V\xrightarrow{\operatorname{id}_{V}\otimes\delta^{\prime}}B\otimes(A\otimes V)\xrightarrow{a^{-1}_{B,A,V}}(B\otimes A)\otimes V\xrightarrow{\omega^{+}\otimes\operatorname{id}_{V}}V),$
where morphisms $f$ are mapped onto $f$. The monoidal structure is given by
$(\Omega,\beta):{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}\rightarrow{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}},\ \beta_{X,Y}=c_{Y,X}^{{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}}\circ\overline{c}_{X,Y}.$
Proof.
s depicted in the diagram above, we define the functor $\Omega$ as the composition:
$\Omega:=F_{rl}\circ F_{lr}\circ\Gamma.$
We now compute the explicit structure of this compositionLet $(V,\lambda_{B},\delta_{B})\in{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$. Then
$F_{rl}F_{lr}\Gamma(V,\lambda_{B},\delta_{B})=$ $F_{rl}F_{lr}(V,\overline{\lambda_{A}},\overline{\delta_{A}})$
$=F_{rl}(V,\overline{\lambda_{A}}_{+},c_{A,V}\circ\overline{\delta_{A}})$
$=(V,(\overline{\lambda_{A}}_{+})_{-},(\mathcal{S}_{A}\otimes\operatorname{id}_{V})\circ c_{V,A}\circ c_{A,V}\circ\overline{\delta_{A}}).$
Note that $(\overline{\lambda_{A}}_{+})_{-}=\overline{\lambda_{A}}$, while
$\overline{\lambda_{A}}=(A\otimes V\xrightarrow{\operatorname{id}_{A}\otimes\delta_{B}}A\otimes(B\otimes V)\xrightarrow{a^{-1}_{A,B,V}}(A\otimes B)\otimes V\xrightarrow{\omega\otimes\operatorname{id}_{V}}V).$
This coincides with our definition of $\lambda_{A}$.
Now we need to show $(\mathcal{S}_{A}\otimes\operatorname{id}_{V})\circ c_{V,A}\circ c_{A,V}\circ\overline{\delta_{A}}$ equals $\delta_{A}$. Since $\mathcal{S}_{A}$ and $c_{A,V}$ are isomorphisms, we only need to prove:
$\delta^{\prime}=(\mathcal{S}_{A}^{-2}\otimes\operatorname{id}_{V})\circ\overline{c}^{2}_{A,V}\circ(\mathcal{S}_{A}\otimes\operatorname{id}_{V})\circ c^{2}_{A,V}\circ\overline{\delta_{A}}=(\mathcal{S}_{A}^{-1}\otimes\operatorname{id}_{V})\circ\overline{\delta_{A}}.$
It suffices to show
$(\omega^{+}\otimes\operatorname{id}_{V})\circ a^{-1}_{B,A,V}\circ(\operatorname{id}_{B}\otimes(\mathcal{S}_{A}^{-1}\otimes\operatorname{id}_{V})\circ\overline{\delta_{A}})=\lambda_{B}.$
Since $a$ is a natural isomorphism, and $\omega^{+\operatorname{cop}}=\omega^{+}\circ(\operatorname{id}_{B}\otimes\mathcal{S}_{A}^{-1})$, we have
$(\omega^{+}\otimes\operatorname{id}_{V})\circ a^{-1}_{B,A,V}\circ(\operatorname{id}_{B}\otimes(\mathcal{S}_{A}^{-1}\otimes\operatorname{id}_{V})\circ\overline{\delta_{A}})$ $=(\omega^{+}\otimes\operatorname{id}_{V})\circ((\operatorname{id}_{B}\otimes\mathcal{S}_{A}^{-1})\otimes\operatorname{id}_{V})\circ a^{-1}_{B,A,V}\circ(\operatorname{id}_{B}\otimes\overline{\delta_{A}})$
$=(\omega^{+\operatorname{cop}}\otimes\operatorname{id}_{V})\circ a^{-1}_{B,A,V}\circ(\operatorname{id}_{B}\otimes\overline{\delta_{A}})$
$=\lambda_{B}.$
Where the last equality follows from the relation between $\overline{\delta_{A}}$ and $\lambda_{B}$ via the functor $D_{2}$. Thus $\Omega$ induces a monoidal equivalence $\Omega:\overline{{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}\rightarrow\overline{{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}$.
For the monoidal structure of $\Omega$, note that only the functor $(F_{rl},\alpha)$ is non-strict. Since $F_{lr}$ and $\Gamma$ are braided, we have
$\overline{\beta}_{X,Y}=\overline{c}^{\overline{{\mathcal{YD}(\mathcal{C})^{A}_{A}}}}_{F_{lr}\Gamma(Y),F_{lr}\Gamma(X)}\circ\overline{c}_{F_{lr}\Gamma(X),F_{lr}\Gamma(Y)}=\overline{c}^{\overline{{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}}}_{Y,X}\circ\overline{c}_{X,Y}.$
Thus $(\Omega,\overline{\beta}):\overline{{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}\rightarrow\overline{{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}$ is a braided monoidal equivalence. Now we take reverse braiding, and it is direct that
$(\Omega,\beta):{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}\rightarrow{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}},\ \beta_{X,Y}=c_{Y,X}^{{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}}\circ\overline{c}_{X,Y}.$
is a braided monoidal equivalence.
4 Applications to Nichols algebras over coquasi-Hopf algebras
We now specialize to the setting of Nichols algebras.
4.1 Hopf pairings for Nichols algebras
Let $\mathcal{C}$ be an abelian braided monoidal category. We first recall the definition of Nichols algebras $\mathcal{B}(V)$ in $\mathcal{C}$, where $V\in\mathcal{C}$ is an arbitrary object. Then we are going to show that there is a non-degenerate Hopf pairing between $\mathcal{B}(V^{*})$ and $\mathcal{B}(V)$ when $V$ has a left dual $V^{*}$.
We denote by $T(V)$ the tensor algebra in $\mathcal{C}$ generated freely by $V$. Here we denote that
$V^{\otimes n}:=\left(\cdots\left(\left(V\otimes V\right)\otimes V\right)\cdots\otimes V\right).$
Then $T(V)$ is isomorphic to $\bigoplus_{n\geq 0}V^{\otimes n}$ as an object. The tensor algebra $T(V)$ is naturally a graded Hopf algebra in $\mathcal{C}$.
Let $i_{n}:V^{\otimes n}\rightarrow T(V)$ be the canonical injection. We denote the total Woronowicz symmetriser by
$\operatorname{Wor}(c)=\bigoplus_{n\geq 1}i_{n}[n]!_{c},$
where $n]!_{c}\in\operatorname{End}(V^{\otimes n})$ is the braided symmertriser. For explicit definition of $n]!_{c}$, one may refer to [Bazlov and Berenstein, 2013, Section 5.6].
Definition 4.1.
Let $V\in\mathcal{C}$. The Nichols algebra of $V$ is defined to be the quotient Hopf algebra
$\mathcal{B}(V):=T(V)/\operatorname{ker}\operatorname{Wor}(c).$
To facilitate further analysis, we assume $\mathcal{C}$ is a $\mathbbm{k}$-linear braided monoidal abelian category. In this setting, the following equivalent definition of the Nichols algebra is more convenient.
Definition 4.2.
[Meir, 2022, Definition 2.4] For a given object $V\in\mathcal{C}$, the Nichols algebra $\mathcal{B}(V)$ is the unique Hopf algebra in $\mathcal{C}$ that satisfies the following conditions:
$(1)$ The Hopf algebra $\mathcal{B}(V)$ is graded by the non-negative integers.
$(2)$ The zeroth component of the grading satisfies $\mathcal{B}(V)_{0}=\mathbbm{k}$.
$(3)$ The first component of the grading satisfies $\mathcal{B}(V)_{1}=V$ , and $\mathcal{B}(V)$ is generated by $V$ as an algebra in $\mathcal{C}$.
$(4)$ The subobject of primitive elements of $\mathcal{B}(V)$ is $V$.
We are going to investigate the Hopf pairing between $\mathcal{B}(V^{*})$ and $\mathcal{B}(V)$ when $V$ has a left dual $V^{*}$. There is a unique Hopf pairing $\omega:T(V^{*})\otimes T(V)\to\textbf{1}$. It is shown in [Majid, 1993] that $\omega$ is given by
$\omega\mid_{V^{*}\otimes V}=\operatorname{ev}_{V}.$
and
$\omega_{n}:=\omega|_{V^{*\otimes n}\otimes V^{\otimes n}}=\operatorname{ev}_{V^{\otimes n}}\circ(\operatorname{id}_{V^{*\otimes n}}\otimes[n]!_{c})=\operatorname{ev}_{V^{\otimes n}}\circ([n]!_{c^{*}}\otimes\operatorname{id}_{V^{\otimes n}})$
The Hopf pairing between $V^{\otimes n}$ and $V^{*\otimes m}$ is 0 unless $n=m$. Here $c^{*}$ denotes the the braided symmertriser of $V^{*}$.
The Hopf pairing between $T(V)$ and $T(V^{*})$ inducing a Hopf pairing between $\mathcal{B}(V^{*})$ and $\mathcal{B}(V)$. Note that the Hopf pairing $\omega$ vanishes on the kernels of the symmetrisers, and for all positive integer $n$, we have
$\omega_{n}({\operatorname{ker}[n]!_{c}^{*},\ V^{\otimes n}})=0,\ \omega_{n}({V^{*\otimes n},\ \operatorname{ker}[n]!_{c}})=0.$
Now we return to the case of coquasi-Hopf algebra $H$ with bijective antipode. Let $V$ be a finite-dimensional object in $^{H}_{H}\mathcal{YD}$}. Consequently, the Hopf pairing on the tensor algebras descends to a well-defined Hopf pairing
$\omega:\mathcal{B}(V^{*})\otimes\mathcal{B}(V)\rightarrow\mathbbm{k}.$
In particular,
$\omega(V^{*\otimes m},V^{\otimes n})=0,\ \text{for all}\ m\neq n,$
and both $\mathcal{B}(V)$ and $\mathcal{B}(V^{*})$ are locally finite $\mathbb{N}_{0}$-graded Hopf algebras in $^{H}_{H}\mathcal{YD}$}.
We consider $T(V)^{\perp}$, by expression of $\omega$ it equals $\operatorname{ker}\operatorname{Wor}(c^{*})=\bigoplus_{n\geq 1}i_{n}\operatorname{ker}([n]!_{c}^{*})$. Similarly, $T(V^{*})^{\perp}=\operatorname{ker}\operatorname{Wor}(c)$. Since $\mathcal{B}(V)=T(V)/\operatorname{ker}\operatorname{Wor}(c)$ and $\mathcal{B}(V^{*})=T(V^{*})/\operatorname{ker}\operatorname{Wor}(c^{*})$, we deduce that $\omega:\mathcal{B}(V^{*})\otimes\mathcal{B}(V^{*})\rightarrow\mathbbm{k}$ is a dual pair of locally finite $\mathbb{N}_{0}$-graded Hopf algebras in $^{H}_{H}\mathcal{YD}$}.
Corollary 4.3.
For any finite-dimensional $V\in{{}^{H}_{H}\mathcal{YD}}$, there exists a braided monoidal equivalence:
$(\Omega_{V},\beta_{V}):{{}^{\mathcal{B}(V)}_{\mathcal{B}(V)}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}\longrightarrow{{}^{\mathcal{B}(V^{*})}_{\mathcal{B}(V^{*})}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}.$ (4.1)
Proof.
This follows immediately from the existence of a dual pair $\omega:\mathcal{B}(V^{*})\otimes\mathcal{B}(V)\rightarrow\mathbbm{k}$ and Theorem 3.16
Let $A$, $B$ be $\mathbb{N}_{0}$-graded locally finite Hopf algebra in $\mathcal{C}$ with bijective antipodes and $\omega:A\otimes B\rightarrow\mathbbm{k}$ a dual pair. Recall the braided monoidal equivalence in Theorem 3.16:
$\Omega:{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}\longrightarrow{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}.$
Corollary 4.4.
The Nichols algebra $\mathcal{B}(V)$ of any $V\in{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}$ is rational if $V$ is.
Proof.
If $V$ is rational, then for any $n\in\mathbb{N}_{\geq 1}$, $V^{\otimes n}$ is rational in $^{B}_{B}\mathcal{YD}(\mathcal{C})$}. Since ${{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$ is closed under direct sum and quotient by Lemma 3.3 both objects $T(V)=\bigoplus_{n\geq 0}V^{\otimes n}$ and $\mathcal{B}(V)$ are rational. ∎
Now for $Q\in{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$, $\mathcal{B}(Q)$ is an $\mathbb{N}_{0}$-graded Hopf algebra in ${{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$ as well. The following result is straightforward.
Corollary 4.5.
Let $Q\in{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$, then
$\Omega(\mathcal{B}(Q))\cong\mathcal{B}(\Omega(Q))$ (4.2)
as $\mathbb{N}_{0}$-graded Hopf algebras in ${{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$.
Proof.
It follows from [Li and Liu, 2025b, Lemma 2.16] immediately. ∎
4.2 The functor $\Omega$ under $\mathbb{Z}$-gradings
[…]
Furthermore, suppose $H$ is a $\mathbb{N}_{0}$-graded coquasi-Hopf algebra. A $\mathbb{Z}$-graded Yetter-Drinfeld module $V$ over $H$ is by definition an object $V$ in ${{}^{H}_{H}\mathcal{YD}}(\mathbb{Z}\text{-Gr}\mathcal{M}_{k})$. In other words, $V$ is a $\mathbb{Z}$-graded vector space such that the map $H\otimes V\rightarrow V$ and comodule structure maps $V\rightarrow H\otimes V$ are $\mathbb{Z}$-graded.
Lemma 4.6.
[Li and Liu, 2025b, Lemma 3.9] Suppose $H$ is a $\mathbb{Z}$-graded coquasi-Hopf algebra.
$(1)$ Let $K$ be a Nichols algebra in ${{}^{H}_{H}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$, and $K(1)=\oplus_{\gamma\in\mathbb{Z}}K(1)_{\gamma}$ a $\mathbb{Z}$-graded object in ${{}^{H}_{H}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$. Then $K$ admits a unique $\mathbb{Z}$-grading that extends the grading on $K(1)$. Moreover, $K(n)$ is $\mathbb{Z}$-graded in ${{}^{H}_{H}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$ for all $n\geq 0$.
$(2)$ Let $K$ be a $\mathbb{Z}$-graded braided Hopf algebra in ${{}^{H}_{H}\mathcal{YD}}(\mathbb{Z}\text{-Gr}\mathcal{M}_{k})$. Then the bosonization K#HK\#H is a $\mathbb{Z}$-graded Hopf algebra with $\deg K(\gamma)\#H(\lambda)=\gamma+\lambda$ for all $\gamma,\lambda\in\mathbb{Z}$.
$(3)$ Let $H_{0}\subseteq H$ be a coquasi-Hopf subalgebra of degree 0, and $\pi:H\rightarrow H_{0}$ a coquasi-Hopf algebra map with $\pi|_{H_{0}}=\mathrm{id}$. Define $R=H^{\operatorname{co}H_{0}}$. Then $R$ is a $\mathbb{Z}$-graded braided Hopf algebra in ${{}^{H_{0}}_{H_{0}}\mathcal{YD}}(\mathbb{Z}\text{-Gr}\mathcal{M}_{k})$ with $R(\gamma)=R\cap H(\gamma)$ for all $\gamma\in\mathbb{Z}$.
From now on, $H$ represents a coquasi-Hopf algebra with bijective antipode, let $A,B\in{{}^{H}_{H}\mathcal{YD}}$ be locally finite $\mathbb{N}_{0}$-graded Hopf algebras with dual pairing $\omega$. Moreover,
$\omega(A(m),B(n))=0,\ \ \text{for all}\ m\neq n.$
To extend this construction to the $\mathbb{Z}$-graded setting, we extend the $\mathbb{N}_{0}$-grading to $\mathbb{Z}$-grading by
A(n)=0,B(n)=0for alln<0.A(n)=0,\ B(n)=0\ \ $\text${for all}\ n<0.
We endow $A$ with a new $\mathbb{Z}$-grading as follows: $\operatorname{deg}(A(n))=-n$ for all $n\in\mathbb{Z}$, which ensures that $\omega:A\otimes B\rightarrow\mathbb{k}$ is $\mathbb{Z}$-graded. Here, the grading of $\mathbbm{k}$ is given by $\mathbbm{k}(n)=0$, if $n\neq 0$, and $\mathbbm{k}(0)=\mathbbm{k}$.
Now we return to the purpose of this subsection. We first introduce the following notations for further use.
Definition 4.7.
Let $H$ be a coquasi-Hopf algebra with bijective antipode, and $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$. Suppose $R$ is an $\mathbb{N}_{0}$-graded Hopf algebra in $^{H}_{H}\mathcal{YD}$}. Let $(X,\lambda_{X})$ be a $R$-module in $\mathcal{C}$ and $(Y,\delta_{Y})$ be a $R$-comodule in $\mathcal{C}$. For $n\geq 0$, we define:
$\mathcal{F}_{n}X$ $=\{x\in X\mid\lambda_{X}(R(n)\otimes x)=0\text{ for all }i>n\},\$
$\mathcal{F}^{n}Y$ $=\{y\in Y\mid\delta_{Y}(y)\in\bigoplus_{i=0}^{n}R(i)\otimes Y\}.$
With above notations, the functor $\Omega$ relates the operator $\mathcal{F}_{n}$ to $\mathcal{F}^{n}$.
Lemma 4.8.
Let $N$ be a finite-dimensional object in $^{H}_{H}\mathcal{YD}$}. Recall we have the following equivalence of monoidal categories:
$\Omega:{{}^{\mathcal{B}(N)}_{\mathcal{B}(N)}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}\cong{{}^{\mathcal{B}(N^{*})}_{\mathcal{B}(N^{*})}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}.$
Let $V\in{{}^{\mathcal{B}(N)}_{\mathcal{B}(N)}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$, then we have
(1) $\mathcal{F}_{n}\Omega(V)=\mathcal{F}^{n}V$ for all $n\geq 0$.
(2) $\mathcal{F}^{n}\Omega(V)=\mathcal{F}_{n}(V)$ for all $n\geq 0$.
(3) Suppose $V$ is a $\mathbb{Z}$-graded object in $^{\mathcal{B}(N)}_{\mathcal{B}(N)}\mathcal{YD}(\mathcal{C})$} , then $\Omega(V)$ is a $\mathbb{Z}$-graded object in $^{\mathcal{B}(N^{*})}_{\mathcal{B}(N^{*})}\mathcal{YD}(\mathcal{C})$}, where the grading of $\Omega(V)$ is given by $\Omega(V)(n)=V(-n)$ for all $n\in\mathbb{Z}$.
Proof.
Recall that the braided monoidal equivalence of abelian category
$\Omega:{{}^{\mathcal{B}(N)}_{\mathcal{B}(N)}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}\longrightarrow{{}^{\mathcal{B}(N^{*})}_{\mathcal{B}(N^{*})}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}},$
sends
$(V,\lambda_{\mathcal{B}(N)},\delta_{\mathcal{B}(N)})\mapsto(V,\lambda_{\mathcal{B}(N^{*})},\delta_{\mathcal{B}(N^{*})}).$
which is induced by the dual pair $\omega:\mathcal{B}(N^{*})\otimes\mathcal{B}(N)\rightarrow\mathbbm{k}$.
(1) Recall
$\lambda_{\mathcal{B}(N^{*})}=(\mathcal{B}(N^{*})\otimes V\xrightarrow{\operatorname{id}\otimes\delta_{\mathcal{B}(N)}}\mathcal{B}(N^{*})\otimes(\mathcal{B}(N)\otimes V)\xrightarrow{a^{-1}_{\mathcal{B}(N^{*}),\mathcal{B}(N),V}}(\mathcal{B}(N^{*})\otimes\mathcal{B}(N))\otimes V\xrightarrow{\omega\otimes\operatorname{id}_{V}}\mathcal{B}(N^{*})),$
Now for fixed $n\geq 0$, the kernel of the induced map:
$\mathcal{B}(N^{*})\otimes V\longrightarrow\operatorname{Hom}(\mathcal{B}(N)(n),V),\ \ a\otimes v\mapsto(b\mapsto\Phi(a_{-1},b_{-1},v_{-1})^{-1}\langle a_{0},b_{0}\rangle v_{0})$ (4.3)
is $\bigoplus_{m\neq n}\mathcal{B}(N^{*})(m)\otimes V$. This means, if $v\in\mathcal{F}^{n}V$, then for each $i>n$, $\lambda_{\mathcal{B}(N^{*})}(\mathcal{B}(N^{*})(i)\otimes v)=0$ by (4.3). On the other hand, if $v\in\mathcal{F}_{n}(V,\lambda_{\mathcal{B}(N^{*})})$, this will imply $\delta(v)\in\bigoplus_{i=0}^{n}\mathcal{B}(N)(i)\otimes V$. which completes the proof of (1): $\mathcal{F}_{n}\Omega(V)=\mathcal{F}_{n}(V,\lambda_{\mathcal{B}(N^{*})})=\mathcal{F}^{n}V$.
(2) For the second statement, recall $\delta_{\mathcal{B}(N^{*})}$ is given by
$\delta_{\mathcal{B}(N^{*})}=(V\xrightarrow{\delta^{\prime}}\mathcal{B}(N^{*})\otimes V\xrightarrow{\mathcal{S}_{\mathcal{B}}(N^{*})^{2}\otimes\operatorname{id}_{V}}\mathcal{B}(N^{*})\otimes V\xrightarrow{c^{2}_{\mathcal{B}(N^{*}),V}}\mathcal{B}(N^{*})\otimes V),\ \text{ where }\ \delta^{\prime}\ \text{is defined by}$
$\lambda_{\mathcal{B}(N)}=(\mathcal{B}(N)\otimes V\xrightarrow{\operatorname{id}_{V}\otimes\delta^{\prime}}\mathcal{B}(N)\otimes(\mathcal{B}(N^{*})\otimes V)\xrightarrow{a^{-1}_{\mathcal{B}(N),\mathcal{B}(N^{*}),V}}(\mathcal{B}(N)\otimes\mathcal{B}(N^{*}))\otimes V\xrightarrow{\omega^{+}\otimes\operatorname{id}_{V}}V).$
By [Heckenberger and Schneider, [2020] ©2020, Corollary 3.3.6, Lemma 12.2.11], $(V,\delta^{\prime})$ is $\mathcal{B}(N^{*})$-comodule in $^{H}_{H}\mathcal{YD}$}, and we have:
$\mathcal{F}^{n}(V,\delta^{\prime})=\mathcal{F}^{n}\Omega(V).$
A similar argument to that in (1) shows that $\mathcal{F}_{n}(V)=\mathcal{F}^{n}(V,\delta^{\prime})$. This combines together shows $\mathcal{F}_{n}(V)=\mathcal{F}^{n}(V,\delta^{\prime})=\mathcal{F}^{n}\Omega(V)$.
(3) We regard $\mathcal{B}(N)$ and $\mathcal{B}(N^{*})$ as $\mathbb{Z}$-graded Hopf algebras in $^{H}_{H}\mathcal{YD}$}. Let $V=\bigoplus_{n\in\mathbb{Z}}V(n)$ be its gradation, and set $\Omega(V)(n)=V(-n)$ for all $n\in\mathbb{Z}$. Then for $m\geq 0$, by definition of $\rho_{\mathcal{B}(N^{*})}$, we have
$\lambda_{\mathcal{B}(N^{*})}(\mathcal{B}(N^{*})(m)\otimes\Omega(V)(n))=\lambda_{\mathcal{B}(N^{*})}(\mathcal{B}(N^{*})(m)\otimes V(-n))\subseteq V(-m-n)=\Omega(V(m+n)).$
On the other hand, by definition of $\delta_{\mathcal{B}(N^{*})}$
$\delta_{\mathcal{B}(N^{*})}(\Omega(V)(n))=\delta_{\mathcal{B}(N^{*})}(V(-n))\subseteq\bigoplus_{i\in\mathbb{Z}}\mathcal{B}(N^{*})(i)\otimes V(i-n)=\bigoplus_{i\in\mathbb{Z}}\mathcal{B}(N^{*})(i)\otimes\Omega(V(n-i)).$
The above two equations imply that $\Omega(V)$ is a well-defined $\mathbb{Z}$-graded object in $^{\mathcal{B}(N^{*})}_{\mathcal{B}(N^{*})}\mathcal{YD}(\mathcal{C})$}. ∎
5 Reflection of Nichols algebras over coquasi-Hopf algebra with bijective antipode
Let $H$ be a coquasi-Hopf algebra with bijective antipode. In this section, we study projection of Nichols algebras in $^{H}_{H}\mathcal{YD}$}, which will play an important role when considering reflection of simple Yetter-Drinfeld modules. Some proofs in this section are parallel to those in the pointed coquasi-Hopf algebra setting Li and Liu (2025b), we will omit them for simplicity.
5.1 The structure of the space of coinvariants $K$
We first consider the projection of Nichols algebras for further study of reflection of Nichols algebras.
For any $M,N\in{{}^{H}_{H}\mathcal{YD}}$, there is a canonical surjection of braided Hopf algebra in $^{H}_{H}\mathcal{YD}$}
$\pi_{\mathcal{B}(N)}:\mathcal{B}(M\oplus N)\rightarrow\mathcal{B}(N),\ \ \ \pi_{\mathcal{B}(N)}\mid_{N}=\operatorname{id},\ \ \ \pi_{\mathcal{B}(N)}\mid_{M}=0.$
This surjection naturally induces a canonical projection of coquasi-Hopf algebras:
$\pi=\pi_{\mathcal{B}(N)}\#\operatorname{id}_{H}:\mathcal{B}(M\oplus N)\#H\rightarrow\mathcal{B}(N)\#H.$
To simplify our notation, we introduce the following conventions:
$\mathcal{A}(M\oplus N):=\mathcal{B}(M\oplus N)\#H,\ \mathcal{A}(N):=\mathcal{B}(N)\#H.$
This projection allows us to construct the associated space of coinvariants. There is a natural injection $\iota:\mathcal{A}(N)\rightarrow\mathcal{A}(M\oplus N)$ such that $\pi\circ\iota=\operatorname{id}_{\mathcal{A}(N)}$.
Denote $\pi_{H}:\mathcal{A}(M\oplus N)\rightarrow H$, and $\pi^{\prime}_{H}:\mathcal{A}(N)\rightarrow H$ be the natural projection. These maps satisfy a compatibility condition, namely, there exists a commutative diagram relating these projections:
[…]
Let
$K=\mathcal{A}(M\oplus N)^{\operatorname{co}\mathcal{A}(N)}$
be the space of right $\mathcal{A}(N)$-coinvariant elements with respect to the projection $\pi$. By virtue of Lemma 2.9, $K$ inherits the structure of a braided Hopf algebra in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} with $\mathcal{A}(N)$-coaction:
$\delta:K\rightarrow\mathcal{A}(N)\otimes K,\ \ \ x_{-1}\otimes x_{0}=:\delta(x):=\pi(x_{1})\otimes x_{2},$
and a linear map:
$\operatorname{ad}:\mathcal{A}(N)\otimes K\rightarrow K,\ \ \ \ a\otimes x\mapsto\operatorname{ad}(a)(x)=\Phi(a_{1}x_{-1},\mathbb{S}(a_{3})_{1},a_{4})(a_{2}x_{0})\mathbb{S}(a_{3})_{2}.$ (5.1)
Here, $\mathbb{S}$ denotes preantipode of $\mathcal{A}(N)$. We have the following two observations.
Lemma 5.1.
[Li and Liu, 2025b, Lemma 4.1, Lemma 4.2]
(1) The space $K$ coincides with $\mathcal{B}(M\oplus N)^{\operatorname{co}\mathcal{B}(N)}$, which is precisely the space of right $\mathcal{B}(N)$-coinvariant elements with respect to $\pi_{\mathcal{B}(N)}$.
(2) The space of primitive elements $P(K)=\{x\in K\mid\Delta_{K}(x)=x\otimes 1+1\otimes x\}$ is a subobject of $K$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}.
One may expect that $K$ is a Nichols algebra generated by $\operatorname{ad}(\mathcal{B}(N))(M)$. The following lemma establish that $\operatorname{ad}(\mathcal{B}(N))(M)$ is a subobject of $K$.
Lemma 5.2.
Let $M,N\in{{}^{H}_{H}\mathcal{YD}}$, $K=(\mathcal{A}(M\oplus N))^{\operatorname{co}\mathcal{A}(N)}$. The standard $\mathbb{N}_{0}$-grading of $\mathcal{B}(M\oplus N)$ induces an $\mathbb{N}_{0}$-grading on
$L:=\operatorname{ad}(\mathcal{B}(N))(M)=\bigoplus_{n\in\mathbb{N}}\operatorname{ad}(N)^{n}(M)$
with degree $\operatorname{deg}(\operatorname{ad}(N)^{n}(M))=n+1$. Then $L$ is a $\mathbb{N}_{0}$-graded object in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}. Moreover, $L\subseteq K$ is a subobject of $K$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}.
Proof.
Let $a\in N$ and $x\in\mathcal{B}(M\oplus N)$ be homogeneous elements. By equation (2.13), we have
$\operatorname{ad}(a)(x)=ax-\operatorname{ad}(a_{-1})(x)a_{0}.$
The element $\operatorname{ad}(a)(x)$ is of degree $\operatorname{deg}(X)+1$ in the Nichols algebra $\mathcal{B}(M\oplus N)$. Moreover, we have $\operatorname{ad}(N)^{n}(M)\cap\operatorname{ad}(N)^{m}(M)=0$ for $n\neq m$. This is implies the direct sum decomposition of $L$.
Since $M\subseteq K$ and $K\in{{}^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}}$, we conclude that $L=\operatorname{ad}(\mathcal{B}(N))(M)\subseteq K$. Note that
$\operatorname{ad}(\mathcal{B}(N))(M)\subseteq\operatorname{ad}(\mathcal{A}(N))(M).$ Conversely, since $M\in{{}^{H}_{H}\mathcal{YD}}$, we have
$\operatorname{ad}(\mathcal{A}(N))(M)\subseteq\operatorname{ad}(\mathcal{B}(N))(\operatorname{ad}(H)(M))\subseteq\operatorname{ad}(\mathcal{B}(N))(M).$
Hence $L=\operatorname{ad}(\mathcal{A}(N))(M)$.
It is clear that the map
$\operatorname{ad}:\mathcal{A}(N)\otimes L\rightarrow L$
is well-defined since $\operatorname{ad}$ satisfies the first axiom of Yetter-Drinfeld modules. Furthermore, the map $\operatorname{ad}:\mathcal{A}(N)\otimes L\rightarrow L$ is $\mathbb{N}_{0}$-graded. Next, we want to show $L$ is a $\mathcal{A}(N)$-comodule. The comodule structure is given by $\delta_{K}$. For $a\in N$ and $x\in L$. Since $a$ is primitive in $\mathcal{B}(M\oplus N)$, we have
$\Delta_{\mathcal{A}(M\oplus N)}(a)=a_{-1}\otimes a_{0}+a\otimes 1.$
A direct computation shows:
$\delta_{K}(\operatorname{ad}(a)(x))$ $=p_{R}((\operatorname{ad}(a_{3})(x_{0}))_{-1},a_{4})q_{R}(a_{1}x_{-2},\mathcal{S}(a_{6}))(a_{2}x_{-1})\mathcal{S}(a_{5})\otimes(\operatorname{ad}(a_{3})(x_{0}))_{0}$
$=p_{R}((\operatorname{ad}(a_{0})(x_{0}))_{-1},1)q_{R}(a_{-2}x_{-2},1)a_{-1}x_{-1}\otimes(\operatorname{ad}(a_{0})(x_{0}))_{-1}$
$+p_{R}((\operatorname{ad}(a_{-2})(x_{0}))_{-1},a_{-1})q_{R}(a_{-4}x_{-2},1)(a_{-3}x_{-1})\mathcal{S}(a_{0})\otimes(\operatorname{ad}(a_{-2})(x_{0}))_{0}$
$=a_{-1}x_{-1}\otimes\operatorname{ad}(a_{0})(x_{0})+p_{R}((\operatorname{ad}(a_{-2})(x_{0}))_{-1},a_{-1})(a_{-3}x_{-1})\mathcal{S}(a_{0})\otimes(\operatorname{ad}(a_{-2})(x_{0}))_{0}.$
Since $x\in L=\bigoplus_{n\geq 0}\operatorname{ad}(N)^{n}(M)$. We proceed by induction on $n$. For $n=0$, note that $\operatorname{ad}(a_{0})(x_{0})\in L$, hence the first term lies in $\mathcal{A}(N)\otimes L$. For the second term, note that $\operatorname{ad}(a_{-2})(x_{0})\in M$, since $a_{-2}\in H$ has degree zero. Therefore $(\operatorname{ad}(a_{-2})(x_{0}))_{0}\in M$. Consequently, $\delta_{K}(\operatorname{ad}(a)(x))\in\mathcal{A}(N)\otimes L$.
Now we assume that for some fixed integer $n_{0}$, we have $\delta_{K}(\operatorname{ad}(a)(x))\in\mathcal{A}(N)\otimes L$ for all $x\in\operatorname{ad}(N)^{n^{\prime}}(M)$, where $n^{\prime}\leq n_{0}$. Then for $a\in N$ and $x\in\operatorname{ad}(N)^{n_{0}+1}(M)$. Since by definition, $\operatorname{ad}(N)^{n_{0}+1}(M)=\operatorname{ad}(N)(\operatorname{ad}(N)^{n_{0}}(M))$. We have $\delta_{K}(x)\in\mathcal{A}(N)\otimes L$ by inductive hypothesis. Thus $\operatorname{ad}(a_{0})(x_{0})\in\operatorname{ad}(N)(L)\subseteq L$. On the other hand, $\operatorname{ad}(a_{-2})(x_{0})\in L$ as well. Since $\operatorname{ad}(a_{-2})(x_{0})\in\bigoplus_{i=0}^{n_{0}+1}\operatorname{ad}(N)^{i}(M)$, we have $(\operatorname{ad}(a_{-2})(x_{0}))_{0}\in L$ by inductive hypothesis. This completes the induction, and we shows $\delta_{K}:L\rightarrow\mathcal{A}(N)\otimes L$ is well-defined. The third axiom holds automatically since $L\subseteq K$. Thus $L$ is a $\mathbb{N}_{0}$-graded object in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} and is a subobject of $K$.
The following lemma is used to prove the semisimplicity of the object $L$ in the case that $M$ is semisimple. It guarantees $L$ to be a $\mathcal{B}(N)$-comodule in $^{H}_{H}\mathcal{YD}$} via restricting the following composition to $L$,
$\delta:\mathcal{B}(M\oplus N)\xrightarrow{\Delta_{\mathcal{B}(M\oplus N)}}\mathcal{B}(M\oplus N)\otimes\mathcal{B}(M\oplus N)\xrightarrow{\pi_{\mathcal{B}(N)}\otimes\operatorname{id}}\mathcal{B}(N)\otimes\mathcal{B}(M\oplus N).$
Lemma 5.3.
For all $x\in L$, we have $\Delta_{\mathcal{B}(M\oplus N)}(x)-x\otimes 1\in\mathcal{B}(N)\otimes L$.
Proof.
Recall $L=\bigoplus_{n\in\mathbb{N}_{0}}\operatorname{ad}(N)^{n}(M)$. We proceed by induction on $n$. Now let $a\in N$ and $x\in M$ be homogeneous elements. A direct computation yields:
$\Delta_{\mathcal{B}(M\oplus N)}(\operatorname{ad}(a)(x))=\Delta_{\mathcal{B}(M\oplus N)}(ax-(\operatorname{ad}(a_{-1})(x)a_{0}))$
$=ax\otimes 1+1\otimes ax+a\otimes x+\operatorname{ad}(a_{-1})(x)\otimes a_{0}-\operatorname{ad}(a_{-1})(x)a_{0}\otimes 1-1\otimes\operatorname{ad}(a_{-1})(x)a_{0}-\operatorname{ad}(a_{-1})(x)\otimes a_{0}$
$-\operatorname{ad}((\operatorname{ad}(a_{-1})(x))_{-1})(a_{0})\otimes(\operatorname{ad}(a_{-1})(x))_{0}$
$=\operatorname{ad}(a)(x)\otimes 1+1\otimes\operatorname{ad}(a)(x)+a\otimes x-\operatorname{ad}((\operatorname{ad}(a_{-1})(x))_{-1})(a_{0})\otimes(\operatorname{ad}(a_{-1})(x))_{0}.$
Note that $x=\operatorname{ad}(1)(x)\in L$. Since $\operatorname{ad}(a_{-1})(x)\in L$, we have $(\operatorname{ad}(a_{-1})(x))_{0}\in L$ by Lemma 5.2. Therefore $\Delta_{\mathcal{B}(M\oplus N)}(\operatorname{ad}(a)(x))-\operatorname{ad}(a)(x)\otimes 1\in\mathcal{B}(N)\otimes L$.
For a fixed $m\in\mathbb{N}_{0}$. We assume that for all $y\in\operatorname{ad}(N)^{m^{\prime}}(M)$, $1\leq m^{\prime}\leq m$, we have
$\Delta_{\mathcal{B}(M\oplus N)}(y)-y\otimes 1\in\mathcal{B}(N)\otimes L.$
Now let $a\in N$ and $x\in\operatorname{ad}(N)^{m}(M)$ be homogeneous, we have
$\Delta_{\mathcal{B}(M\oplus N)}(\operatorname{ad}(a)(x))$ $=\Delta_{\mathcal{B}(M\oplus N)}(ax-(\operatorname{ad}(a_{-1})(x))a_{0})$
$=(a\otimes 1+1\otimes a)(x^{1}\otimes x^{2})$
$-\frac{\Phi(a_{-5},x^{1}_{-1},x^{2}_{-2})\Phi((\operatorname{ad}(a_{-4})(x^{1}_{0}))_{-1},(\operatorname{ad}(a_{-2})(x^{2}_{0}))_{-1},a_{-1})}{\Phi((\operatorname{ad}(a_{-4})(x^{1}_{0}))_{-2},a_{-3},x^{2}_{-1})}\times$
$((\operatorname{ad}(a_{-4})(x^{1}_{0}))_{0}\otimes((\operatorname{ad}(a_{-2})(x^{2}_{0}))_{0})(1\otimes a_{0}+a_{0}\otimes 1).$
Since the term $x\otimes 1$ lies in $\Delta_{\mathcal{B}(M\oplus N)}(x)$, let us consider this expression separately. The term $\operatorname{ad}(a)(x)\otimes 1$ will occur in $\Delta_{\mathcal{B}(M\oplus N)}(\operatorname{ad}(a)(x))$. Now the left terms are
$\sum_{x^{1}\otimes x^{2}\neq x\otimes 1}\Phi(a_{-1},x^{1}_{-1},x^{2}_{-1})a_{0}x^{1}_{0}\otimes x^{2}_{0}+\frac{\Phi(a_{-3},x^{1}_{-1},x^{2}_{-2})}{\Phi((\operatorname{ad}(a_{-2})(x^{1}_{0}))_{-1},a_{-1},x^{2}_{-1})}(\operatorname{ad}(a_{-2})(x^{1}_{0}))_{0}\otimes a_{0}x^{2}_{0}$
$-\frac{\Phi(a_{-6},x^{1}_{-1},x^{2}_{-2})\Phi((\operatorname{ad}(a_{-5})(x^{1}_{0}))_{-2},(\operatorname{ad}(a_{-3})(x^{2}_{0}))_{-2},a_{-2})}{\Phi((\operatorname{ad}(a_{-5})(x^{1}_{0}))_{-3},a_{-4},x^{2}_{-1})\Phi((\operatorname{ad}(a_{-5})(x^{1}_{0}))_{-1},(\operatorname{ad}(a_{-3})(x^{2}_{0}))_{-1},a_{-1})}(\operatorname{ad}(a_{-5})(x^{1}_{0}))_{0}\otimes(\operatorname{ad}(a_{-3})(x^{2}_{0}))_{0}a_{0}$
$-\frac{\Phi(a_{-5},x^{1}_{-1},x^{2}_{-2})\Phi((\operatorname{ad}(a_{-4})(x^{1}_{0}))_{-2},(\operatorname{ad}(a_{-2})(x^{2}_{0}))_{-3},a_{-1})}{\Phi((\operatorname{ad}(a_{-4})(x^{1}_{0}))_{-3},a_{-3},x^{2}_{-1})}\times$
$\Phi((\operatorname{ad}(a_{-4})(x^{1}_{0}))_{-1},(\operatorname{ad}((\operatorname{ad}(a_{-2})(x^{2}_{0}))_{-2})(a_{0}))_{-1},(\operatorname{ad}(a_{-2})(x^{2}_{0}))_{-1})\times$
$(\operatorname{ad}(a_{-4})(x^{1}_{0}))_{0}(\operatorname{ad}((\operatorname{ad}(a_{-2})(x^{2}_{0}))_{-2})(a_{0}))_{0}\otimes(\operatorname{ad}(a_{-2})(x^{2}_{0}))_{0}.$
By assumptions,
$\Delta_{\mathcal{B}(M\oplus N)}(x)-x\otimes 1\in\mathcal{B}(N)\otimes L.$
Therefore the first term lies in $\mathcal{B}(N)\otimes{L}$. For the third term, note that $\Phi$ is convolution invertible, thus the third term becomes
$\sum_{x^{1}\otimes x^{2}\neq x\otimes 1}\frac{\Phi(a_{-3},x^{1}_{-1},x^{2}_{-2})}{\Phi((\operatorname{ad}(a_{-2})(x^{1}_{0}))_{-1},a_{-1},x^{2}_{-1})}(\operatorname{ad}(a_{-2})(x^{1}_{0}))_{0}\otimes\operatorname{ad}(a_{-1})(x_{0}^{2})a_{0}.$
Now we combine the second and third term together, yields
$\sum_{x^{1}\otimes x^{2}\neq x\otimes 1}\frac{\Phi(a_{-3},x^{1}_{-1},x^{2}_{-2})}{\Phi((\operatorname{ad}(a_{-2})(x^{1}_{0}))_{-1},a_{-1},x^{2}_{-1})}(\operatorname{ad}(a_{-2})(x^{1}_{0}))_{0}\otimes\operatorname{ad}(a_{0})(x^{2}_{0}).$
By assumption, $\sum_{x^{1}\otimes x^{2}\neq x\otimes 1}x^{1}\otimes x^{2}\in\mathcal{B}(N)\otimes L$, therefore this term lies in $\mathcal{B}(N)\otimes L$. The last term belongs to $\mathcal{B}(N)\otimes{L}$ by similar reason, which completes the induction and proves the lemma.
The proof of next lemma proceeds exactly as in the cosemisimple case, so we omit here.
Lemma 5.4.
[Li and Liu, 2025b, Lemma 4.5]
$(1)$ Assume $M=\bigoplus_{i\in I}M_{i}$ is a direct sum in $^{H}_{H}\mathcal{YD}$}. Let $L_{i}=\operatorname{ad}\mathcal{B}(N)(M_{i})$ for all $i\in I$. Then we have such a decomposition in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}
$L=\bigoplus_{i\in I}L_{i}.$
$(2)$ If $M$ is irreducible in $^{H}_{H}\mathcal{YD}$}, then $L=\operatorname{ad}\left(\mathcal{B}(N)\right)(M)$ is irreducible in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}.
We now proceed to prove that $K$ possesses the structure of a Nichols algebra in an appropriate Yetter-Drinfeld module category.
Theorem 5.5.
There is an isomorphism
$K\cong\mathcal{B}(L)$ (5.2)
of Hopf algebras in the category $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}. In particular, $P(K)=L$.
Proof.
The proof is parallel to proof of group case [Li and Liu, 2025b, Theorem 4.9]. However, some processes have become even more difficult, we give a proof here for completeness.
We endow a non-standard grading on $\mathcal{B}(M\oplus N)$ via setting $\operatorname{deg}(M)=1$ and $\operatorname{deg}(N)=0$. Under this grading, $\mathcal{B}(M\oplus N)$ remains a $\mathbb{N}_{0}$-graded Hopf algebra in $^{H}_{H}\mathcal{YD}$}. Furthermore, $\mathcal{B}(M\oplus N)\#H$ becomes a $\mathbb{N}_{0}$-graded coquasi-Hopf algebra with $\operatorname{deg}(H)=0$. Now we are going to show $K$ is a Nichols algebra by verifying the axiom in Definition 4.2.
(1) It follows direct that $K$ inherits an $\mathbb{N}_{0}$-graded Hopf algebra in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} with $K(n)=K\cap(\mathcal{B}(M\oplus N)\#H)(n)$.
(2) We have $K(0)=K\cap(\mathcal{B}(M\oplus N)\#H)(0)=K\cap(\mathcal{B}(M\oplus N)(0)=\mathbbm{k}$.
(3) Note that $K(1)=K\cap(\mathcal{B}(M\oplus N)\#H)(1)=K\cap(\mathcal{B}(M\oplus N)(1))=\operatorname{ad}(\mathcal{B}(N))(M)=L$. We are going to show $K$ is generated its degree one component $L$. Let $K^{\prime}$ be the subalgebra of $K$ generated by $L$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}, which is an object in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} with structures induced by tensor product of $L$. Now we let $W$ be the image of $K^{\prime}\#\mathcal{B}(N)$ under the isomorphism $K\#\mathcal{B}(N)\cong\mathcal{B}(M\oplus N)$. It suffices to show that $W=\mathcal{B}(M\oplus N)$. Note that $M\oplus N\subseteq W$ since $M,N\subseteq W$. Thus, we need only verify that $W$ forms a subalgebra of $\mathcal{B}(M\oplus N)$.
The key observation is the stability of $K^{\prime}$ under the map of $\mathcal{B}(N)$. Indeed, $L$ itself is stable under the map of $\mathcal{B}(N)$. Now consider arbitrary homogeneous elements $y,z\in L$ and $a\in\mathcal{A}(N)$, by (2.5),
$\operatorname{ad}(a)(y\otimes z)$
$=\frac{\Phi(a_{1},y_{-1},z_{-2})\Phi((\operatorname{ad}(a_{2})(y_{0}))_{-1},(\operatorname{ad}(a_{4})(z_{0}))_{-1},a_{5}))}{\Phi(({\operatorname{ad}(a_{2})(y_{0})})_{-2},a_{3},z_{-1})}(\operatorname{ad}(a_{2})(y_{0}))_{0}\otimes(\operatorname{ad}(a_{4})(z_{0}))_{0}.$
Since $L$ is an object in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}, it follows that $\operatorname{ad}(a)(y\otimes z)\in L\otimes L$. Thus $K^{\prime}$ remains stable under the map induced by $\mathcal{B}(N)$ by induction.
As both $K^{\prime}$ and $\mathcal{B}(N)$ are subalgebras of $\mathcal{B}(M\oplus N)$, we only need to examine the case when $a\in N$ and $x\in K^{\prime}$ . In this situation:
$\operatorname{ad}(a)(x)=ax-\operatorname{ad}(a_{-1})(x)a_{0}.$
Since $\operatorname{ad}(a)(x)\in K^{\prime}$ and $\operatorname{ad}(a_{-1})(x)a_{0}\in K^{\prime}\#\mathcal{B}(N)$, we conclude that $ax\in K^{\prime}\#\mathcal{B}(N)$. Hence its image lies in $W$. Thus $W$ is a subalgebra of $\mathcal{B}(M\oplus N)$, which establishes that $K$ is generated by $L$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}.
(4) To show $P(K)=K(1)$, that is, there is no primitive element in $K(n)$, $n\geq 2$. Suppose, for contradiction, that there exists a nonzero subspace $U\subseteq P(K(n))$ for some positive integer $n$. By Lemma 5.1, $U$ is an object in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}, since $\operatorname{deg}(N)=\operatorname{deg}(H)=0$. Now, consider the coalgebra filtration on $\mathcal{B}(N)\#H$ given by $(\mathcal{B}(N)\#H)_{0}=H$, $(\mathcal{B}(N)\#H)_{1}=H+N$. Applying [Li and Liu, 2025b, Lemma 4.11], we find a nonzero element $u\in U$ such that $\delta_{K}(u)=u_{-1}\otimes u_{0}\in H\otimes U$. Note that
$\delta_{K}(u)=(\pi\otimes\operatorname{id})\Delta_{\mathcal{A}(M\oplus N)}(u).$
Therefore
$\Delta_{\mathcal{A}(M\oplus N)}(u)=u\otimes 1+u_{-1}\otimes u_{0}.$
Applying the map $\mathrm{id}\#\varepsilon\otimes\mathrm{id}$, we obtain
$\Delta_{\mathcal{B}(M\oplus N)}(u)=(\operatorname{id}\#\varepsilon\otimes\operatorname{id})\Delta(u)=1\otimes u+u\otimes 1,$
which implies that $u$ is a primitive element in $\mathcal{B}(M\oplus N)$. However, since $K(n)$ is generated by $L=\operatorname{ad}(\mathcal{B}(N))(M)$, the element $u$ must have degree at least $n$ in the standard grading of $\mathcal{B}(M\oplus N)$. This leads to a contradiction, as primitive elements in a Nichols algebra lie in degree one.
We now turn to a converse of the preceding result under an additional restriction.
Theorem 5.6.
Let $N\in{{}^{H}_{H}\mathcal{YD}}$ and $(W,\rhd,\delta)$ be a semisimple object in the category ${{}^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$ with linear map $\rhd:\mathcal{A}(N)\otimes W\rightarrow W$. Here $\mathcal{A}(N)$ is equipped with the standard $\mathbb{N}_{0}$-grading. Let $\mathcal{B}(W)$ be the Nichols algebra of $W$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} and define $M=\{w\in W\mid\delta(w)\in H\otimes W\}$. Then there is a unique isomorphism
$\mathcal{B}(W)\#\mathcal{B}(N)\cong\mathcal{B}(M\oplus N)$ (5.3)
of braided Hopf algebras in $^{H}_{H}\mathcal{YD}$} which is the identity on $M\oplus N$.
Proof.
Let $W=\bigoplus_{i\in I}W_{i}$ be the decomposition of $W$ into irreducible objects in the category
${{}^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$. By [Li and Liu, 2025b, Lemma 4.12], $M$ is an object in $^{H}_{H}\mathcal{YD}$} with decomposition into irreducible objects $M=\bigoplus_{i\in I}M_{i}:=\bigoplus_{i\in I}M\cap W_{i}$. Moreover, we have $W_{i}=\mathcal{B}(N)\rhd M_{i}=\bigoplus_{n\in\mathbb{N}_{0}}N^{\otimes n}\rhd M_{i}$ for each $i\in I$.
Let $\operatorname{deg}(M_{i})=m_{i}$ for each $i\in I$. We endow $W_{i}$ a new grading by $\widetilde{W_{i}}=W_{i}$ with
$\widetilde{W_{i}}(n)=W(n+m_{i}-1)=N^{n-1}\rhd M_{i},\ \text{for all}\ n\in\mathbb{N}_{0}.$
[…]
$\mathcal{B}(\widetilde{W})(0)=\mathbbm{k},\ \ \mathcal{B}(\widetilde{W})(1)=M.$
Moreover, $\mathcal{B}(W)\#(\mathcal{B}(N)\#H)$ is a $\mathbb{N}_{0}$-graded coquasi-Hopf algebra, and
$R:=\mathcal{B}(W)\#\mathcal{B}(N)=(\mathcal{B}(W)\#(\mathcal{B}(N)\#H))^{\operatorname{co}H}$
is a $\mathbb{N}_{0}$-graded Hopf algebra in $^{H}_{H}\mathcal{YD}$}. Its degree 0’s part is $\mathbbm{k}$, and its degree $1$’s part is $M\oplus N$.
We are going to show $R$ is a pre-Nichols algebra in $^{H}_{H}\mathcal{YD}$}, it remains to show $R$ is generated by $M\oplus N$ as an algebra in $^{H}_{H}\mathcal{YD}$}. It is direct to see that $R$ is generated by $K(1)=\mathcal{B}(N)\rhd M$ and $N$. It therefore suffices to prove that $\mathcal{B}(N)\rhd M=\bigoplus_{n\geq 0}N^{\otimes n}\rhd M$ is contained in the subalgebra generated by $\mathcal{B}(N)$ and $M$. To see this, we proceed by induction on $n$. For the base case, take elements $x\in M$ and $y\in N$. Inside the coquasi-Hopf algebra $\mathcal{B}(W)\#(\mathcal{B}(N)\#H)$, we have
$yx=(y_{1}\rhd x)y_{2}=(y_{-1}\rhd x)y_{0}+(y\rhd x).$
Since $M\in{{}^{H}_{H}\mathcal{YD}}$ and $y_{-1}\in H$, the term $y\rhd x$ is contained in the subalgebra generated by $\mathcal{B}(N)$ and $M$. Since $\mathcal{B}(N)\rhd M$ is an object in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}, $\delta(y\rhd x)\in\mathcal{A}(N)\otimes(N\rhd M)$. This implies terms like $(y\rhd x)_{0}$ are contained in the subalgebra generated by $\mathcal{B}(N)$ and $M$. Let us fixed $k\geq 0$, for all $z\in\mathcal{B}(N)(k^{\prime})$ and $x\in M$, where $k^{\prime}\geq k$. We assume the element $z\rhd x$ as well as $(z\rhd x)_{0}$ lie in the subalgebra generated by $\mathcal{B}(N)$ and $M$, then for homogeneous $y\in N$,
$(yz)\rhd x$ $=\frac{\Phi(y_{-1},(z_{2}\rhd x_{0})_{-1},z_{3})}{\Phi(y_{-2},z_{1},x_{-1})}y_{0}\rhd(z_{2}\rhd x_{0})_{0}$
$=\frac{\Phi(y_{-2},(z_{2}\rhd x_{0})_{-1},z_{3}))}{\Phi(y_{-3},z_{1},x_{-1})}(y_{-1}\rhd(z_{2}\rhd x_{0})_{0})y_{0}-\frac{\Phi(y_{-1},(z_{2}\rhd x_{0})_{-1},z_{3})}{\Phi(y_{-2},z_{1},x_{-1})}y_{0}(z_{2}\rhd x_{0})_{0}.$
By the induction hypothesis, $(yz)\rhd x$ lies in the subalgebra generated by $\mathcal{B}(N)$ and $M$. Therefore, $R$ is generated by $M\oplus N$, and hence a pre-Nichols algebra of $M\oplus N$.
By the universal property of $\mathcal{B}(M\oplus N)$, there exists a surjective morphism of $\mathbb{N}_{0}$-graded Hopf algebras in $^{H}_{H}\mathcal{YD}$}:
$\rho:R\rightarrow\mathcal{B}(M\oplus N),\rho\mid_{M\oplus N}=\operatorname{id}.$
This induces a surjective coquasi-Hopf algebra map:
$\rho\#\operatorname{id}:R\#H\rightarrow\mathcal{B}(M\oplus N)\#H.$
Let $K=(\mathcal{B}(M\oplus N)\#H)^{\operatorname{co}\mathcal{B}(N)\#H}$. Then we obtain two bijective coquasi-Hopf algebra maps:
$R\#H\rightarrow\mathcal{B}(W)\#(\mathcal{B}(N)\#H),\ \ K\#(\mathcal{B}(N)\#H)\rightarrow\mathcal{B}(M\oplus N)\#H.$
Then the map $\rho\#\operatorname{id}$ thus induces a surjective map of coquasi-Hopf algebras
$\rho^{\prime}:\mathcal{B}(W)\#(\mathcal{B}(N)\#H)\rightarrow K\#(\mathcal{B}(N)\#H),\ \ \rho^{\prime}\mid_{(M\oplus N)}=\operatorname{id}.$
Moreover, the following diagram commutes:
$\mathcal{B}(W)\#(\mathcal{B}(N)\#H)$$K\#(\mathcal{B}(N)\#H)$$\mathcal{B}(N)\#H$$\mathcal{B}(N)\#H$$\rho^{\prime}$$\varepsilon_{\mathcal{B}(W)}\#\operatorname{id}_{\mathcal{B}(N)\#H}$$\varepsilon_{K}\#\operatorname{id}_{\mathcal{B}(N)\#H}$$\operatorname{id}$
since ${\rho\mid}_{M\oplus N}=\operatorname{id}$. As $\mathcal{B}(N)\#H$ acts on $K$ via adjoint map. Theorem 5.5 implies that $K\cong\mathcal{B}(\operatorname{ad}(\mathcal{B}(N))(M))$. Therefore $\rho^{\prime}$ induces a surjective map
$\phi:\mathcal{B}(W)\rightarrow\mathcal{B}(\operatorname{ad}(\mathcal{B}(N))(M))$
in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} between the right coinvariant subspaces of $\varepsilon_{\mathcal{B}(W)}\#\operatorname{id}_{\mathcal{B}(N)\#H}$ and $\varepsilon_{K}\#\operatorname{id}_{\mathcal{B}(N)\#H}$, satisfying $\phi\mid_{M}=\operatorname{id}$. Furthermore, there is a surjective map in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}:
$\phi_{1}:\mathcal{B}(N)\rhd M\rightarrow\operatorname{ad}(\mathcal{B}(N))(M),\ \ \phi_{1}\mid_{M}=\operatorname{id}.$
Since $M=\bigoplus_{i\in I}M_{i}$ is a decomposition into irreducible objects in $^{H}_{H}\mathcal{YD}$}, each $\operatorname{ad}(\mathcal{B}(N))(M_{i})$ is irreducible in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}. On the other hand, each $\mathcal{B}(N)\rhd M_{i}$ is irreducible as well. Thus $\phi_{1}$ is an isomorphism in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}, and it follows that $\phi$ is an isomorphism of Hopf algebra in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}. Consequently, $\rho\#\operatorname{id}_{H}=\phi\#\operatorname{id}_{\mathcal{A}(N)}$ is an isomorphism of coquasi-Hopf algebras. Therefore,
$\rho=(\rho\#\operatorname{id}_{H})^{\operatorname{co}H}:\mathcal{B}(W)\#\mathcal{B}(N)\rightarrow\mathcal{B}(M\oplus N)$
is an isomorphism of Hopf algebras in $^{H}_{H}\mathcal{YD}$}.
5.2 The semi-Cartan graph of a Nichols algebra
In this subsection, we develop reflections of Nichols algebras over arbitrary coquasi-Hopf algebra with bijective antipode. The reflection allows us to systematically relate different Nichols algebra realizations through a well-defined transformation procedure.
Fix a positive integer $\theta$ and denote the index set $\mathbb{I}=\{1,2,...,\theta\}$.
Definition 5.7.
Let $\mathcal{F}_{\theta}$ denote the class of all $\theta$-tuples $M=(M_{1},\ldots,M_{\theta})$, where $M_{1},\ldots,M_{\theta}\in{{}^{H}_{H}\mathcal{YD}}$ are finite-dimensional Yetter-Drinfeld modules. If $M\in\mathcal{F}_{\theta}$, we define
$\mathcal{B}(M):=\mathcal{B}(M_{1}\oplus\cdots\oplus M_{\theta}).$
Two tuples $M,M^{\prime}\in\mathcal{F}_{\theta}$ are called isomorphic, denoted $M\cong M^{\prime}$, if $M_{j}\cong M_{j}^{\prime}$ in $^{H}_{H}\mathcal{YD}$} for all j. The isomorphism class of $M\in\mathcal{F}_{\theta}$ is denote by $[M]$.
For $1\leq i\leq\theta$ and $M\in\mathcal{F}_{\theta}$, we say the tuple $M$ admits the $i$-th reflection $R_{i}(M)$ if for all $j\neq i$ there is a natural number $m_{ij}^{M}\geq 0$ such that $(\mathrm{ad}\,M_{i})^{m_{ij}^{M}}(M_{j})$ is a non-zero finite-dimensional subspace of $\mathcal{B}(M)$, and $(\mathrm{ad}\,M_{i})^{m_{ij}^{M}+1}(M_{j})=0$. Assume $M$ admits the $i$-th reflection. Then we set $R_{i}(M)=(V_{1},\ldots,V_{\theta})$, where
$V_{j}=\begin{cases}M_{i}^{*},&\text{if }j=i,\\ (\mathrm{ad}\,M_{i})^{m_{ij}^{M}}(M_{j}),&\text{if }j\neq i.\end{cases}$
Such a question arises naturally: how do the irreducibility properties behave under reflections? The following result provides a reassuring answer.
Lemma 5.8.
Suppose $M\in\mathcal{F}_{\theta}$ admits the $i$-th reflection for some $i\in\mathbb{I}$, and $M_{j}$ is irreducible for each $j\in\mathbb{I}$. Then each $R_{i}(M)_{j}$ is irreducible in $^{H}_{H}\mathcal{YD}$} for $1\leq j\leq\theta$.
Proof.
By definition, $R_{i}(M)$ is defined, thus ${R_{i}(M)}_{i}\cong M_{i}^{*}$ is irreducible since $M_{i}$ is. For $j\neq i$, we observe that $\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})=\bigoplus_{n=0}^{m_{ij}}\operatorname{ad}(M_{i})^{n}(M_{j})$. Consider the embedding $H\rightarrow\mathcal{A}(M_{i})$ and the projection $\mathcal{A}(M_{i})\rightarrow H$. The object $\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})$ belongs to $^{H}_{H}\mathcal{YD}$}. Now suppose $0\neq P\subset\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})$ is a Yetter-Drinfeld submodule over $^{H}_{H}\mathcal{YD}$}. Let $\langle P\rangle$ be the $\mathcal{A}(M_{i})$-subcomodule of $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ generated by $P$, defined explicitly as
$\langle P\rangle:=\{\langle f,x_{-1}\rangle x_{0}\mid x\in P,f\in\operatorname{Hom}(\mathcal{A}(M_{i}),\mathbbm{k})\}.$
To establish this, we prove $\langle P\rangle$ is a Yetter-Drinfeld submodule of $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$}. To see this, for all $a\in\mathcal{A}(M_{i})$ and $x\in P$, by Lemma 2.6,
$\operatorname{ad}(a)$ $(\langle f,x_{-1}\rangle x_{0})=\langle f,x_{-1}\rangle\operatorname{ad}(a)(x_{0})$
$=\langle f,p_{L}(\mathcal{S}(a_{1}),(\operatorname{ad}(a_{4})(x_{0}))_{-1}a_{6})q_{L}(a_{3},x_{-1})\mathcal{S}(a_{2})((\operatorname{ad}(a_{4})(x_{0}))_{-2}a_{5})\rangle(\operatorname{ad}(a_{4})(x_{0}))_{0}.$
[…]
$\operatorname{ad}(a)(\langle f,x_{-1}\rangle x_{0})=t*s\left((\operatorname{ad}(a_{4})(x_{0}))_{-1}\right)(\operatorname{ad}(a_{4})(x_{0}))_{0}.$
Hence $\operatorname{ad}(a)(\langle f,x_{-1}\rangle x_{0})\in\langle P\rangle.$ This ensures the first axiom of a Yetter-Drinfeld module, and the third axiom holds automatically. Since $\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})$ is generated by $\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})$ as a Yetter-Drinfeld module and is irreducible in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$} by Lemma 5.4 (2). Then $\langle P\rangle=\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})$ since $P\neq 0$. This forces $P=\operatorname{ad}(M_{i})^{m_{ij}}(M_{j}),$ establishing the irreducibility.
The following lemma is standard, we list here for completeness.
Lemma 5.9.
[Li and Liu, 2025b, Lemma 5.4, Lemma 5.5]
(1) Suppose $M\in\mathcal{F}_{\theta}$ and $M$ admits $i$-th reflection for each $i\in\mathbb{I}$. We define $a_{ii}^{M}=2$ for all $1\leq i\leq\theta$ and define $a_{ij}^{M}=-m_{ij}^{M}$. Then $(a_{ij}^{M})_{i,j\in\mathbb{I}}$ is a generalized Cartan matrix.
(2) Suppose $M\cong N$ in $\mathcal{F}_{\theta}$, if $M$ admits the $i$-th reflection for some $i\in\mathbb{I}$, so is $N$. Furthermore, $R_{i}(M)\cong R_{i}(N)$ and $a_{ij}^{M}=a_{ij}^{N}$ for each $j\in\mathbb{I}$.
Recall by Remark 2.12, let $A$ be a Hopf algebra in $^{H}_{H}\mathcal{YD}$} with bijective antipode, we have such a braided monoidal isomorphism
$F:{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}\cong{{}^{A\#H}_{A\#H}\mathcal{YD}}.$
Now restricting to rational Yetter-Drinfeld modules, we denote the image of $F$ by ${{}^{A\#H}_{A\#H}\mathcal{YD}}_{\operatorname{rat}}$, which is a monoidal full subcategory of $^{A\#H}_{A\#H}\mathcal{YD}$.
From now on, we always fix a tuple $M=(M_{1},\ldots,M_{\theta})\in\mathcal{F}_{\theta}$, with each component $M_{i}$ is irreducible for $i\in\mathbb{I}$. By Corollary 4.3, we have such a braided tensor equivalence for each $i\in\mathbb{I}$.
$\Omega_{i},:{{}^{\mathcal{B}(M_{i})}_{\mathcal{B}(M_{i})}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}\longrightarrow{{}^{\mathcal{B}(M_{i}^{*})}_{\mathcal{B}(M_{i}^{*})}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}.$ (5.4)
By abusing notation, we again denote the following monoidal equivalence by $\Omega_{i}$:
${{}^{A(M_{i})}_{A(M_{i})}\mathcal{YD}}_{\operatorname{rat}}\cong{{}^{\mathcal{B}(M_{i})}_{\mathcal{B}(M_{i})}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}\rightarrow{{}^{\mathcal{B}(M_{i}^{*})}_{\mathcal{B}(M_{i}^{*})}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}\cong{{}^{A(M_{i}^{*})}_{A(M_{i}^{*})}\mathcal{YD}}_{\operatorname{rat}}.$
Lemma 5.10.
The following are equivalent:
(1) $M$ admits $i$-th reflection for some $i\in I$.
(2) $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}$ belongs to ${{}^{A(M_{i})}_{A(M_{i})}\mathcal{YD}}_{\operatorname{rat}}$.
Proof.
(1) $\Rightarrow$ (2): Let $W=\bigoplus_{j\neq i}M_{j}$, and define $Q=\operatorname{ad}\mathcal{B}(M_{i})(W)$. By Lemma 5.4, we have the decomposition $Q=\bigoplus_{j\neq i}Q_{j}$, where each $Q_{j}=\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ is irreducible in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$}. Applying Theorem 5.5 yields an isomorphism of Hopf algebras:
$\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\cong\mathcal{B}(Q).$
Now by assumption, there is an integer $m$ such that $\operatorname{ad}(M_{i})^{m}(W)=0$. This implies that $\operatorname{ad}(M_{i})^{m}(Q)=0$. Thus $Q$ is rational, thus $\mathcal{B}(Q)$ as well as $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}$ are rational by Lemma 3.3(3).
(2) $\Rightarrow$ (1): The condition (2) implies that $Q$ is rational in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$}. Furthermore, $M$ admits the $i$-th reflection by definition. ∎
Lemma 5.11.
With above assumptions on $M$, and suppose $M$ admits the $i$-th reflection with $R_{i}(M)=(V_{1},\ldots,M_{i}^{*},\ldots,V_{\theta})$, then the following equalities hold.
$\operatorname{ad}(\mathcal{B}(M_{i})(M_{j}))=\bigoplus_{n=0}^{m_{ij}}\operatorname{ad}(M_{i})^{n}(M_{j})$ $,\ \ \ \ V_{j}=\mathcal{F}_{0}(\operatorname{ad}\mathcal{B}(M_{i})(M_{j})),$
$\Omega_{i}(\operatorname{ad}(\mathcal{B}(M_{i})(M_{j})))=\bigoplus_{n=0}^{m_{ij}}\operatorname{ad}(M_{i}^{*})^{n}(V_{j})$ $,\ \ \ \ M_{j}\cong\operatorname{ad}(M_{i}^{*})^{m_{ij}}(V_{j}).$
Proof.
This proof is parallel to [Li and Liu, 2025b, Lemma 5.6]. ∎
The next theorem gives a natural explanation of reflections of tuples of Yetter-Drinfeld modules. Although the proof strategy parallels the group case [Li and Liu, 2025b, Theorem 5.7], we include the details here for the sake of completeness and the reader’s convenience.
Theorem 5.12.
With above assumptions on $M$, and suppose $M$ admits the $i$-th reflection, then there is an isomorphism of Hopf algebras in $^{H}_{H}\mathcal{YD}$}:
$\Theta_{i}:\mathcal{B}(R_{i}(M))\cong\Omega_{i}\left(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\right)\#\mathcal{B}(M_{i}^{*}).$ (5.5)
Proof.
Let $N=\bigoplus_{j\neq i}M_{j}$, and define $Q=\operatorname{ad}\mathcal{B}(M_{i})(N)$. By Lemma 5.4(1), we have the decomposition $Q=\bigoplus_{j\neq i}Q_{j}$, where each $Q_{j}=\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ is irreducible in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$}. Applying Theorem 5.5 yields an isomorphism of Hopf algebras:
$\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\cong\mathcal{B}(Q).$
Now observe that the functor $\Omega_{i}$ sends Nichols algebras to Nichols algebras by Corollary 4.5. Therefore,
$\Omega_{i}\left(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\right)\cong\Omega_{i}(\mathcal{B}(Q))\cong\mathcal{B}(\Omega_{i}(Q)).$
Hence, in the category $^{H}_{H}\mathcal{YD}$},
$\Omega_{i}\left(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\right)\#\mathcal{B}(M_{i}^{*})\cong\mathcal{B}(\Omega_{i}(Q))\#\mathcal{B}(M_{i}^{*}).$
Since $Q$ is a direct sum of irreducible objects, it is semisimple. The equivalence $\Omega_{i}$ preserves semisimplicity, so $\Omega_{i}(Q)$ is also semisimple in ${{}^{A(M_{i}^{*})}_{A(M_{i}^{*})}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$ by Lemma 4.8 (3). Then we apply Proposition 5.6 to obtain
$\mathcal{B}\left(\Omega_{i}\left(Q\right)\right)\#\mathcal{B}(M_{i}^{*})\cong\mathcal{B}\left(\mathcal{F}^{0}(\Omega_{i}(Q))\oplus M_{i}^{*}\right).$
According to Lemma 5.11 and Lemma 4.8(2), $\mathcal{F}^{0}(\Omega_{i}(Q))=\mathcal{F}_{0}(Q)=\bigoplus_{j\neq i}V_{j}$. Thus
$\mathcal{B}(\Omega_{i}(Q))\#\mathcal{B}(M_{i}^{*})\cong\mathcal{B}\left(\bigoplus_{j\neq i}V_{j}\oplus M_{i}^{*}\right)\cong\mathcal{B}(R_{i}(M)).$
Combining the above isomorphisms, we conclude that in $^{H}_{H}\mathcal{YD}$}, there is an isomorphism of Hopf algebras:
$\Theta_{i}:\mathcal{B}(R_{i}(M))\cong\Omega_{i}\left(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\right)\#\mathcal{B}(M_{i}^{*}).$
Having defined individual reflections, we now extend this notion to sequences of reflections, which will be important for our study of repeated reflections of tuples.
Definition 5.13.
Let $M\in\mathcal{F}_{\theta}$, with each $M_{i}$ is irreducible for $i\in I$. For $l\in\mathbb{N}_{0}$ and $i_{1},i_{2},...,i_{l}\in\mathbb{I}$.
$(1)$ We say $M$ admits the reflection sequence $(i_{1},i_{2},...,i_{l})$ if $l=0$ or $M$ amdits the $i_{1}$-th reflection and $R_{i_{1}}(M)$ satisfies the reflection sequence $(i_{2},i_{3},...,i_{l})$.
$(2)$ We say $M$ admits all reflection sequences if $M$ admits all reflection sequence $(i_{1},i_{2},...,i_{l})$ for all $l\in\mathbb{N}_{0}$ and $i_{1},i_{2},...,i_{l}\in\mathbb{I}$.
For $M\in\mathcal{F}_{\theta}$ admitting all reflections, we denote
$\mathcal{F}_{\theta}(M)=\{R_{i_{1}}(...(R_{i_{l}}(M))...)\mid l\in\mathbb{N}_{0},\ i_{1},..,i_{l}\in\mathbb{I}\}.$
Definition 5.14.
Let $\mathbb{I}$ be a non-empty finite set, $\mathcal{X}$ a non-empty set, $r:\mathbb{I}\times\mathcal{X}\rightarrow\mathcal{X}$, $A:\mathbb{I}\times\mathbb{I}\times\mathcal{X}\rightarrow\mathbb{Z}$ maps. For all $i,j\in\mathbb{I}$ and ${X}\in\mathcal{X}$ we write $r_{i}(X)=r(i,X)$, $a_{ij}^{X}=A(i,j,X)$ and $A^{X}=(a_{ij}^{X})_{i,j\in\mathbb{I}}\in\mathbb{Z}^{\mathbb{I}\times\mathbb{I}}$. The quadruple $\mathcal{G}=\mathcal{G}(\mathbb{I},\mathcal{X},(r_{i})_{i\in I},(A^{X})_{X\in\mathcal{X}})$ is called a semi-Cartan graph if for all $X\in\mathcal{X}$, the matrix $A^{X}$ is a generalized Cartan matrix, and the following axioms hold.
$(CG1)$ For all $i\in I$, the map $r_{i}$ satisfies $r_{i}^{2}=\mathrm{id}_{\mathcal{X}}$.
$(CG2)$ For all $i\in I$, $X\in\mathcal{X}$, $A^{X}$ and $A^{r_{i}(X)}$ have the same $i$-th row.
As a consequence of Theorem 5.12, we obtain the following corollary.
Corollary 5.15.
Let $M=(M_{1},\ldots,M_{\theta})\in\mathcal{F}_{\theta}$ with each component $M_{i}$ is irreducible.
(1) For each $i\in\mathbb{I}$, suppose $M$ admits $i$-th reflection, then $R_{i}(M)$ admits $i$-th reflection. Furthermore, we have:
$R_{i}^{2}(M)\cong M,$ (5.6)
and
$a_{ij}^{M}=a_{ij}^{R_{i}(M)},$ (5.7)
for all $1\leq j\leq\theta$.
(2) Suppose $M$ admits all reflections. We define the set
$\mathcal{X}=\{[P]\mid P\in\mathcal{F}_{\theta}(M)\},$
and the map
$r:\mathbb{I}\times\mathcal{X}\rightarrow\mathcal{X},\ i\times[X]\mapsto[R_{i}(X)].$
Furthermore, we assume each $\mathcal{B}(P_{i})$ is finite-dimensional for $P\in\mathcal{X}$ and $i\in\mathbb{I}$, then
$\mathcal{G}(M)=(\mathbb{I},\mathcal{X},r,(A^{X})_{X\in\mathcal{X}}),$
where $A^{[X]}=(a_{ij}^{X})_{i,j\in\mathbb{I}}$ for all $X]\in\mathcal{X}$, is a semi-Cartan graph.
In [Li and Liu, 2025b], we proved that reflection preserves the finite-dimensionality of Nichols algebra, explicitly, we proved
$\operatorname{dim}\mathcal{B}(M)=\operatorname{dim}\mathcal{B}(R_{i}(M)).$
The next proposition generalizes to the case Gelfand-Kirillov dimension.
Proposition 5.16.
Let $M=(M_{1},\ldots,M_{\theta})\in\mathcal{F}_{\theta}$ with each component $M_{i}$ is irreducible. Suppose $M$ admits $i$-th reflection, then
$\operatorname{GKdim}\mathcal{B}(M)=\operatorname{GKdim}\mathcal{B}(R_{i}(M)).$
Proof.
Since $M_{i}$ is finite-dimensional, the Nichols algebra $\mathcal{B}(M_{i})$ is finitely generated in degree 1. Consequently, the Gelfand-Kirillov dimension is well-defined. Now that $M$ admits $i$-th reflection, we have
$\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)=\bigoplus_{j\neq i}(\operatorname{ad}\mathcal{B}(M_{i}))(M_{j})$
is finite-dimensional. We have $\mathcal{B}(M)\simeq\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\#\mathcal{B}(M_{i})$ by Theorem 5.12. Further, $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}$ is generated by $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)$ by Theorem 5.5. Then we have $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)+M_{i}$ generates $\mathcal{B}(M)$ and $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)+M_{i}^{*}$ generates $\Omega_{i}(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})})\#\mathcal{B}(M_{i}^{*})$, and
$(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)+M_{i})^{n}$ $=\bigoplus_{k=0}^{n}\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)^{n-k}\#M_{i}^{k}\quad\text{in }\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\#\mathcal{B}(M_{i}),$
$(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)+M_{i}^{*})^{n}$ $=\bigoplus_{k=0}^{n}\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)^{n-k}\#(M_{i}^{*})^{k}\quad\text{in }\Omega_{i}(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})})\#\mathcal{B}(M_{i}^{*})$
for all $n$, where $M_{i}^{k}$ means $k$-fold product of $M_{i}$ in $\mathcal{B}(M_{i})$. Recall that both $\mathcal{B}(M_{i})$ and $\mathcal{B}(M_{i}^{*})$ are finitely generated $\mathbb{N}_{0}$-graded algebras, and there is a non-degenerate dual pairing between them which is compatible with the grading. Thus $\dim M_{i}^{l}=\dim(M_{i}^{*})^{l}<\infty$ for all $l\in\mathbb{N}_{0}$. Therefore the definition of $\operatorname{GKdim}$ implies that
$\operatorname{GKdim}\mathcal{B}(M)$ $=\limsup_{n\to\infty}\frac{\log\dim(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)+M_{i})^{n}}{\log n}$
$=\limsup_{n\to\infty}\frac{\log\dim(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)+M_{i}^{*})^{n}}{\log n}=\operatorname{GKdim}\Omega_{i}(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})})\#\mathcal{B}(M_{i}^{*}).$
Hence the theorem holds since $\mathcal{B}(R_{i}(M))\simeq\Omega_{i}(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})})\#\mathcal{B}(M_{i}^{*})$ as Hopf algebras in $^{H}_{H}\mathcal{YD}$}. ∎
6 An example of affine Nichols algebra
In this section, we consider a rank three Nichols algebra $\mathcal{B}(M)$ of non-diagonal type in [Li and Liu, 2025b], which plays a central role in classification of finite-dimensional Nichols algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$}, where $G$ is a finite abelian group and $\Phi$ is a $3$-cocycle on $G$. It was proved to be infinite-dimensional via different ways in [Huang et al., 2026,Li and Liu, 2025a,Li and Liu, 2025b]. Moreover, it gives rise to a semi-Cartan graph $\mathcal{G}(M)$. In this section, we prove $\mathcal{G}(M)$ is actually a Cartan graph and find a root system over $\mathcal{G}(M)$. Moreover, we show that $\mathcal{B}(M)$ is affine in the sense of Cuntz et al. (2019).
6.1 The Cartan graph of $\mathcal{B}(M)$
We begin by recalling the specific Nichols algebra $\mathcal{B}(M)$ under consideration. Let $G=\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}=\langle h_{1}\rangle\times\langle h_{2}\rangle\times\langle h_{3}\rangle$, and
$\Phi\left(h_{1}^{i_{1}}h_{2}^{i_{2}}h_{3}^{i_{3}},h_{1}^{j_{1}}h_{2}^{j_{2}}h_{3}^{j_{3}},h_{1}^{k_{1}}h_{2}^{k_{2}}h_{3}^{k_{3}}\right)=(-1)^{k_{1}j_{2}i_{3}},$
be a $3$-cocycle on $G$, where $0\leq i_{l},j_{l},k_{l}\leq 1$, $1\leq l\leq 3$.
By [Huang et al., 2026, Lemma 3.5], we may give a complete list of irreducible Yetter-Drinfeld modules over $^{G}_{G}\mathcal{YD}^{\Phi}$}. In particular, restricting to those irreducible Yetter–Drinfeld modules that generate finite-dimensional Nichols algebras, the isomorphism classes are given by the set
$S=\{[^{h}M]\mid\ ^{h}M\ \text{is irreducible},\ h\in\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}-\{1,h_{1}h_{2}h_{3}\}\}.$ (6.1)
There are exactly six isomorphism classes. Here $^{h}M$ represents that the simple Yetter–Drinfeld module has comodule structure
$^{h}M:=\{w\in M\mid\delta_{M}(w)=h\otimes w\}.$
We now choose explicit representatives of $S$. For $1\leq i\leq 6$, let $M_{i}\in S$ be pairwise non-isomorphic simple modules such that $\operatorname{dim}(M_{i})=2$, and $\operatorname{deg}(M_{1})=h_{1}$, $\operatorname{deg}(M_{2})=h_{2}$, $\operatorname{deg}(M_{3})=h_{3}$, $\operatorname{deg}(M_{4})=h_{1}h_{2}$, $\operatorname{deg}(M_{5})=h_{1}h_{3}$, $\operatorname{deg}(M_{6})=h_{2}h_{3}$. One may refer to Li and Liu (2025b) for explicit expression of Yetter-Drinfeld structure, we omit here for simplicity.
The Nichols algebra $\mathcal{B}(M)$ which we concern is generated by $M_{1}\oplus M_{2}\oplus M_{3}$. Furthermore, we proved that:
Theorem 6.1.
[Li and Liu, 2025b, Theorem 6.11] Let $M=(M_{1},M_{2},M_{3})$ be the $3$-tuple, then $M$ admits all reflections and $\mathcal{G}(M)$ is a standard semi-Cartan graph. In particular, the Cartan matrix is
$\begin{pmatrix}2&-1&-1\\ -1&2&-1\\ -1&-1&2\end{pmatrix}$.
Given a semi-Cartan graph, its Weyl groupoid is defined as follows.
Definition 6.2.
We denote by $\mathcal{D}(\mathcal{X},\operatorname{End}(\mathbb{Z}^{\mathbb{I}}))$ to be the category with objects $\operatorname{Ob}\mathcal{D}(\mathcal{X},\operatorname{End}(\mathbb{Z}^{\mathbb{I}}))=\mathcal{X}$, and morphisms
$\operatorname{Hom}(X,Y)=\{(Y,f,X)\mid f\in\operatorname{End}(\mathbb{Z}^{\mathbb{I}})\},$
where the composition of morphisms is defined by
$(Z,g,Y)\circ(Y,f,X)=(Z,gf,X),\ \text{for all}\ X,Y,Z\in\mathcal{X},\ f,g\in\operatorname{End}(\mathbb{Z}^{\mathbb{I}}).$
Let $\alpha_{i}$, $1\leq i\leq\theta$ be the standard basis of $\mathbb{Z}^{\mathbb{I}}$, and
$s_{i}^{X}\in\operatorname{Aut}(\mathbb{Z}^{\mathbb{I}}),\ s_{i}^{X}(\alpha_{j})=\alpha_{j}-a_{ij}^{X}\alpha_{i},\text{for all}\ j.$
We call the smallest subcategory of $\mathcal{D}(\mathcal{X},\operatorname{End}(\mathbb{Z}^{\mathbb{I}}))$ which contains all morphisms $(r_{i}(X),s_{i}^{X},X)$ with $i\in\mathbb{I}$, $X\in\mathcal{X}$ the Weyl groupoid of $\mathcal{G}$, denoted by $\mathcal{W}(\mathcal{G})$.
We denote the Weyl groupoid structure on $\mathcal{G}(M)$ by $\mathcal{W}(\mathcal{G}(M))$. Furthermore, the set of real roots of $\mathcal{G}(M)$ at $X$
$\Delta^{X,\operatorname{re}}=\{\omega(\alpha_{i})\in\mathbb{Z}^{\mathbb{I}}\mid\omega\in\operatorname{Hom}(\mathcal{W}(\mathcal{G}(M)),X),i\in\mathbb{I}\}$
can be formulated explicitly. Since $\mathcal{G}(M)$ is standard, it is obvious that $\Delta^{X,\operatorname{re}}=\Delta^{Y,\operatorname{re}}$ for all $X,Y\in F_{3}(M)$ and
$\operatorname{Hom}(\mathcal{W}(\mathcal{G}(M)),X)\rightarrow W(A_{2}^{(1)}),\ \ (Y,s,X)\mapsto s$
is bijective, where $W(A_{2}^{(1)})$ is the Weyl group of the affine Lie algebra $A_{2}^{(1)}$.
For an affine Lie algebra $A_{2}^{(1)}$ (and more generally for any symmetrizable Kac–Moody algebra), the set of real roots is closed under the action of the Weyl group, and every real root can be obtained from a simple root by applying a sequence of fundamental reflections, see [Kac, 1990, Proposition 5.1]. If we denote $\delta=\alpha_{1}+\alpha_{2}+\alpha_{3}$. Then
$\Delta^{M,\operatorname{re}}=\{\alpha+k\delta\mid\alpha\in\Phi_{A_{2}},\ k\in\mathbb{Z}\}.$ (6.2)
Here $\Phi_{A_{2}}=\pm\{\alpha_{2},\alpha_{3},\alpha_{2}+\alpha_{3}\}$, which corresponds to the root of Lie algebra $A_{2}$. This agreement is expected because the semi-Cartan graph is standard, and thus $\Delta^{M,\operatorname{re}}$ coincides with the set of real roots for the affine Lie algebra $A_{2}^{(1)}$. Unfortunately, $\Delta^{M,\operatorname{re}}$ is a infinite set, by definition $\mathcal{G}(M)$ is not a finite semi-Cartan graph.
Definition 6.3.
(1) A semi-Cartan graph $\mathcal{G}$ is called connected if the groupoid $\mathcal{W}(\mathcal{G})$ is connected, that is, if for any two objects $X,Y$ of $\mathcal{G}$ there is a morphism from $X$ to $Y$ in $\mathcal{W}(\mathcal{G})$.
(2) A semi-Cartan graph $\mathcal{G}$ is called simply connected if for any two points $X,Y$ of $\mathcal{G}$ there is at most one morphism from $X$ to $Y$ in $\mathcal{W}(\mathcal{G})$.
(3) For any $X\in\mathcal{X}$ and $i,j\in I$ let
$m_{ij}^{X}=|\Delta^{X\,\text{re}}\cap(\mathbb{N}_{0}\alpha_{i}+\mathbb{N}_{0}\alpha_{j})|.$
We say that $\mathcal{G}$ is a Cartan graph if the following hold.
  1. (CG3)
    For all $X\in\mathcal{X}$, the set $\Delta^{X\,\text{re}}$ consists of positive and negative roots.
  2. (CG4)
    Let $X\in\mathcal{X}$, and $i,j\in I$. If $m_{ij}^{X}<\infty$, then $(r_{i}r_{j})^{m_{ij}^{X}}(X)=X$.
Proposition 6.4.
The quadruple $\mathcal{G}(M)$ is a connected and simply connected Cartan graph.
Proof.
Since $M$ admits all reflections, $\mathcal{G}(M)$ is a connected semi-Cartan graph. Moreover, the fact that $\operatorname{Hom}([M],[M])=\operatorname{id}_{[M]}$ implies that $\mathcal{G}(M)$ is simply connected.
The condition (CG3) is obvious, since $\Delta^{X,\operatorname{re}}=\Delta^{Y,\operatorname{re}}$ for all $X,Y\in F_{3}(M)$ and $\Delta^{M,\operatorname{re}}$ consists of positive and negative roots.
For (CG4), by definition,
$m_{12}^{M}=\mid\Delta^{M,\operatorname{re}}\cap(\mathbb{N}_{0}\alpha_{1}+\mathbb{N}_{0}\alpha_{2})\mid=3.$
Now by direct computation of reflection, see [Li and Liu, 2025b, Lemma 6.12],
$(r_{1}r_{2})^{3}(M)$ $\cong r_{1}r_{2}r_{1}r_{2}r_{1}(M_{4},M_{2},M_{6})\cong r_{1}r_{2}r_{1}r_{2}(M_{4},M_{1},M_{5})\cong r_{1}r_{2}r_{1}(M_{2},M_{1},M_{3})$
$\cong r_{1}r_{2}(M_{2},M_{4},M_{6})\cong r_{1}(M_{1},M_{4},M_{5})\cong(M_{1},M_{2},M_{3}).$
We deduce that $(r_{1}r_{2})^{3}([M])=[M]$. Similarly, for all $i,j\in\{1,2,3\}$ and $X\in\mathcal{X}$, one can prove $m_{ij}^{X}=3$, and
$(r_{i}r_{j})^{3}([X])=[X].$
Hence $\mathcal{G}(M)$ is a Cartan graph. ∎
Definition 6.5.
Let $\mathcal{G}=\mathcal{G}(\mathbb{I},\mathcal{X},r,A)$ be a semi-Cartan graph. For all $X\in\mathcal{X}$, let $(R^{X})$ be a subset of $\mathbb{Z}^{I}$ with the following properties.
  1. (1)
    $0\notin R^{X}$ and $\alpha_{i}\in R^{X}$ for all $X\in\mathcal{X}$ and $i\in I$.
  2. (2)
    $R^{X}\subseteq\mathbb{N}_{0}^{I}\cup-\mathbb{N}_{0}^{I}$ for all $X\in\mathcal{X}$.
  3. (3)
    For any $X\in\mathcal{X}$ and $i\in I$, $s_{i}^{X}(R^{X})=R^{r_{i}(X)}$.
  4. (4)
    If $i,j\in\mathbb{I}$ and $X\in\mathcal{X}$ such that $i\neq j$ and $m_{ij}^{X}$ in Definition 6.3 is finite, then $(r_{i}r_{j})^{m_{ij}^{X}}(X)=X.$
Then we say that the pair $(\mathcal{G},(R^{X})_{X\in\mathcal{X}})$ is a root system over $\mathcal{G}$. A root system over $\mathcal{G}$ is said to be reduced if for all $X\in\mathcal{X}$ and $\alpha\in R^{X}$ the roots $\alpha$ and $\alpha$ are the only rational multiples of $\alpha$ in $R^{X}$. A root system is said to be reduced if for all $X\in\mathcal{X}$, $R^{X}$ is finite.
If $\mathcal{G}$ is a Cartan graph, the pair $(\mathcal{G},(\Delta^{X,\operatorname{re}})_{X\in\mathcal{X}})$ is a reduced root system over $\mathcal{G}$ by [Heckenberger and Schneider, [2020] ©2020, Example 10.4.4]. The following corollary follows immediately.
Corollary 6.6.
Let $M$ be the $3$-tuple as above, then $(\mathcal{G}(M),(\Delta^{X,\operatorname{re}})_{X\in\mathcal{X}})$ is a reduced root system over $\mathcal{G}(M)$. Furthermore, there is no finite root system over $\mathcal{G}$.
Proof.
The second claim follows from the fact that $\mathcal{G}$ is not a finite Cartan graph and [Heckenberger and Schneider, [2020] ©2020, Theorem 10.4.7]. ∎
6.2 The Tits cone induced by $(\mathcal{G}(M),(\Delta^{X,\operatorname{re}})_{X\in\mathcal{X}})$ is a half plane
In [Cuntz, 2018,Cuntz et al., 2019], the authors showed that most of Nichols algebras over Hopf algebras can define a Tits cone, via its root system. In particular, they call a Nichols algebra affine if the corresponding Tits cone is a half plane. The following is their main observation:
Theorem 6.7.
[Cuntz et al., 2019, Theorem 1.1] There exists a one-to-one correspondence between connected, simply connected Cartan graphs permitting a root system and crystallographic Tits arrangements with reduced root system.
Under this correspondence, equivalent Cartan graphs correspond to combinatorially equivalent Tits arrangements and vice versa, giving rise to a one-to-one correspondence between the respective equivalence classes.
A crystallographic Tits arrangements consists of a pair $(\mathcal{A},T)$, where $\mathcal{A}$ is a set of linear hyperplanes in $\mathbb{R}^{r}$, and $T$ is a Tits cone in $\mathbb{R}^{r}$, satisfying some additional conditions. For explicit definition of crystallographic Tits arrangements, one may refer to [Cuntz et al., 2019, Definition 3.1].
We give a brief recall of calculating Tits cone when given a connected simply connected Cartan graph $\mathcal{G}=\mathcal{G}(\mathbb{I},\mathcal{X},(r_{i})_{i\in I},(A^{X})_{X\in\mathcal{X}})$ with a root system $R=(\Delta^{X,\operatorname{re}})_{X\in\mathcal{X}}$. Fix an object $A\in\mathcal{X}$. Let $V=\mathbb{R}^{r}$ (where $r=|I|$) and fix a linear isomorphism $\psi:\mathbb{Z}^{I}\to V^{*}$ sending the standard basis to a basis $\{\phi_{1},\dots,\phi_{r}\}$ of $V^{*}$. Define the set of roots $\mathcal{R}=\psi(\Delta^{A,\operatorname{re}})\subset V^{*}$.
For any object $B\in\mathcal{X}$, since $\mathcal{G}$ is connected and simply connected, there exists a unique morphism $w_{B}\in\operatorname{Hom}(A,B)$ in the Weyl groupoid. Define the linear map $\psi_{B}=\psi\circ w_{B}^{-1}:\mathbb{Z}^{I}\to V^{*}$. Then the set $\mathcal{R}^{B}:=\psi_{B}(\{e_{i}\}_{i\in I})$ forms a basis of $V^{*}$, called the root basis at $B$. The corresponding chamber is defined as the open simplicial cone
$K^{B}:=\bigcap_{\beta\in\mathcal{R}^{B}}\beta^{+}=\{x\in V\mid\beta(x)>0\text{ for all }\beta\in\mathcal{R}^{B}\}.$
Let $\mathcal{K}=\{K^{B}\mid B\in\mathcal{X}\}$. The set of hyperplanes is
$\mathcal{A}:=\{\phi^{\perp}\mid\phi\in\mathcal{R}\}.$
Finally, the Tits cone is defined as the convex hull of all chambers:
$T:=\operatorname{conv}\left(\bigcup_{X\in\mathcal{X}}K^{X}\right).$
Theorem 6.7 guarantees that $(\mathcal{A},T)$ is a crystallographic Tits arrangement. The next proposition seems to be new, although those who are familiarly with affine Nichols algebras may regard it as a fact.
Proposition 6.8.
Let $\mathcal{G}=\mathcal{G}(\mathbb{I},\mathcal{X},r,A)$ be a connected, simply connected standard Cartan graph, with $(\Delta^{X,\operatorname{re}})_{X\in\mathcal{X}}$ a reduced root system over $\mathcal{G}$. If the Cartan matrix $A$ is affine, then the corresponding Tits cone $T$ in the sense of Theorem 6.7 is a half plane.
Proof.
We denote $r=|\mathbb{I}|$. Since $A$ is affine, it is automatically symmetrizable, then there exists a positive vector $v=(v_{1},v_{2},..,v_{r})\in\mathbb{Z}_{\geq 0}$ such that $v^{T}A=0$, that is
$\sum_{i=1}^{r}v_{i}c_{ij}=0\ \text{for all}\ j.$
For simplicity, we use $X$ to represent its isomorphism class $X]\in\mathcal{X}$. We fix a base object $M$. Let $V=\mathbb{R}^{r}$ and fix a linear isomorphism $\psi:\mathbb{Z}^{r}\to V^{*}$ sending the standard basis vectors $\{\alpha_{1},\alpha_{2},...,\alpha_{r}\}$ to a basis $\{\phi_{1},\phi_{2},...,\phi_{r}\}$ of $V^{*}$. Define the set of roots $\mathcal{R}=\psi(\Delta^{M,\operatorname{re}})\subset V^{*}$.
For any object $X\in\mathcal{X}$, there is a unique morphism $w_{X}\in\operatorname{Hom}(M,X)$ in the Weyl groupoid. Define $\psi_{X}=\psi\circ w_{X}^{-1}:\mathbb{Z}^{r}\to V^{*}$. The set of root basis at $X$ is
$B^{X}=\psi_{X}(\{\alpha_{1},\alpha_{2},...,\alpha_{r}\}),$
and the associated chamber is
$K^{X}=\bigcap_{\beta\in B^{X}}\beta^{+}=\{x\in V\mid\beta(x)>0\text{ for all }\beta\in B^{X}\}.$
We now prove that
$T=\operatorname{conv}\left(\bigcup_{X\in\mathcal{X}}K^{X}\right)$
coincides with the half-space $\delta^{+}$ where $\delta=v_{1}\phi_{1}+v_{2}\phi_{2}+...+v_{r}\phi_{r}$.
Now let $X,Y\in\mathcal{X}$, such that $r_{i}(X)=Y$ for some $i\in I$. By [Cuntz et al., 2019, Proposition 3.5], if $B^{X}=\{\beta_{1},\beta_{2},...\beta_{r}\}$ is indexed compatibly with $B^{Y}=\{\beta_{1}^{\prime},\beta_{2}^{\prime},...,\beta_{r}^{\prime}\}$, then
$\beta_{i}^{\prime}=-\beta_{i},\hskip 23.49976pt\beta_{j}^{\prime}=\beta_{j}-c_{ji}\beta_{i}\qquad(j\neq i).$
Consequently,
$\sum_{j=1}^{r}v_{j}\beta_{j}^{\prime}$ $=\sum_{j\neq i}v_{j}(\beta_{j}-c_{ji}\beta_{i})-v_{i}\beta_{i}$
$=\sum_{j\neq i}v_{j}\beta_{j}-\sum_{j\neq i}(v_{j}c_{ji}-v_{i})\beta_{i}$
$=\sum_{j\neq i}v_{j}\beta_{j}+v_{i}\beta_{i}=\sum_{j=1}^{r}v_{j}\beta_{j}.$
Note that in this case $K^{X}$ and $K^{Y}$ are adjacent chambers, hence the sum $\delta_{X}:=\sum_{\beta\in B^{X}}\beta$ is invariant under adjacent chambers. Since the Cartan graph is connected, this sum is independent of the chamber $X$; we denote this common value by $\delta$.
For any $X\in\mathcal{X}$ and any $x\in K^{X}$, we have $\beta(x)>0$ for all $\beta\in B^{X}$. Hence
$\delta(x)=\sum_{\beta\in B^{X}}\beta(x)>0,$
so $K^{X}\subset\delta^{+}$. Since $\delta^{+}$ is convex, the convex hull $T$ of the union of all chambers also satisfies $T\subseteq\delta^{+}$.
Consider the affine plane $H_{1}=\{x\in V\mid\delta(x)=1\}$. The intersection of each chamber $K^{X}$ with $H_{1}$ is non‑empty because for any $x\in K^{X}$ we have $\delta(x)>0$, so $x/\delta(x)\in K^{X}\cap H_{1}$. Moreover, each chamber $K^{X}$ is a r‑dimensional open simplicial cone. The intersection $K^{X}$ with $H_{1}$ is a non‑empty (r1)(r-1)-dimensional simplex. In the theory of affine reflection groups, $\Delta:=K^{X}\cap H_{1}$ is known as a fundamental alcove. A classical result [Humphreys, 1990, Chapter 4] states that the affine Weyl group $W=W(A)$ acts simply transitively on the set of alcoves and that the closures of the alcoves tile the space $H_{1}$, i.e.
$H_{1}=\bigcup_{w\in W}w(\overline{\Delta}).$
and the interiors of distinct alcoves are disjoint. Hence the collection $\{K^{X}\cap H_{1}\}_{X\in\mathcal{X}}$ forms a tessellation of $H_{1}$. It covers $H_{1}$ completely and the interiors of different members do not overlap.
The point $x/\delta(x)$ must lie in the closure of some simplex. Consequently, we can find finitely many vertices $p_{1},\dots,p_{n}$ of these triangles, where each vertex belongs to some chamber $K^{X_{i}}\cap H_{1}$ and non‑negative coefficients $t_{1},\dots,t_{n}$ with $\sum_{i=1}^{n}t_{i}=1$ such that
$x=\sum_{i=1}^{n}\delta(x)t_{i}p_{i}=\sum_{i=1}^{n}t_{i}(\delta(x)p_{i}).$
Each $\delta(x)p_{i}$ belongs to $K^{X_{i}}$ because $K^{X_{i}}$ is a cone. Hence $x$ is a convex combination of points from $\bigcup_{X}K^{X}$, so $x\in T$. This proves $\delta^{+}\subseteq T$. Now we obtain $T=\delta^{+}$. Thus by philosophy of Theorem 6.7, the Tits cone corresponding to $(\mathcal{G},(\Delta^{X,\operatorname{re}})_{X\in\mathcal{X}})$ is a half plane. ∎
Now return to the $3$-tuple $M$ in $^{G}_{G}\mathcal{YD}^{\Phi}$}, we have proved that $(\mathcal{G}(M),(\Delta^{X,\operatorname{re}})_{X\in\mathcal{X}})$ is a connected and simply connected standard Cartan graph with a reduced root system, thus it should produce a Tits cone in $\mathbb{R}^{3}$.
Corollary 6.9.
Under the equivalence of Theorem 6.7, the Tits cone corresponding to
$(\mathcal{G}(M),(\Delta^{X,\operatorname{re}})_{X\in\mathcal{X}})$ is a half plane. That is, the Nichols algebra $\mathcal{B}(M)$ is an affine Nichols algebra.
Remark 6.10.
(1) This means we realize an affine Nichols algebra over a coquasi-Hopf algebra, which broadens the realizing approaches of affine Nichols algebras.
(2) We hope to prove that a semi-Cartan graph induced by Nichols algebra over $^{H}_{H}\mathcal{YD}$} is actually a Cartan graph, which will endow them more beautiful properties.
Acknowledgment This work is supported by the National Key R&D Program of China
2024YFA1013802 and NSFC 12271243.

### v2

@@
-Reflection Theory of Nichols Algebras over Coquasi-Hopf Algebras with Bijective Antipode
+arXiv:2602.07946v2[math.QA] 27 Feb 2026
 Abstract
@@
 Nichols algebras play an important role in the theory of Hopf algebras and quantum groups, serving as key building blocks for the classification of pointed Hopf algebras and the structural analysis of braided tensor categories.
-Their combinatorial and representation-theoretic properties are intrinsically linked to the underlying Yetter–Drinfeld modules, and understanding their behavior under natural maps, such as reflections, has become a central theme in the research of Nichols algebra.
+Their combinatorial and representation-theoretic properties are intrinsically linked to the underlying Yetter–Drinfeld modules, and understanding their behavior under natural maps, such as reflections, has become a central theme in the research on Nichols algebra.
 Heckenberger introduced the concepts of semi-Cartan graphs and Weyl groupoids [Heckenberger, 2006,Heckenberger, 2009,Heckenberger, 2010], which provide powerful combinatorial tools for studying the structure and classification of Nichols algebras.
-A semi-Cartan graph is essentially a quadruple $\mathcal{G}=\mathcal{G}(\mathbb{I},\mathcal{X},r,A)$ consisting of a finite index set $\mathbb{I}$, a set of points $\mathcal{X}$, a set of reflection maps $r$, and a generalized Cartan matrix $A^{X}$ associated with each point, satisfying specific compatibility axioms (see Definition 5.14 for details) The most celebrated theorem states that a tuple of simple Yetter-Drinfeld modules over a Hopf algebra with bijective antipode gives rise to a semi-Cartan graph, provided that it admits all reflections [Andruskiewitsch et al., 2010,Heckenberger and Schneider, 2010b,Heckenberger and Schneider, 2013a].
-For Hopf algebras, the reflection theory of Nichols algebras, has been extensively developed and successfully applied to classify finite-dimensional Nichols algebras of diagonal type [Heckenberger, 2006,Heckenberger, 2010,Angiono, 2013,Angiono, 2015].
+A semi-Cartan graph is essentially a quadruple $\mathcal{G}=\mathcal{G}(\mathbb{I},\mathcal{X},r,A)$ consisting of a finite index set $\mathbb{I}$, a set of points $\mathcal{X}$, a set of reflection maps $r$, and a generalized Cartan matrix $A^{X}$ associated with each point $X\in\mathcal{X}$, satisfying specific compatibility axioms (see Definition 5.14 for details) The most celebrated theorem states that a tuple of simple Yetter-Drinfeld modules over a Hopf algebra with bijective antipode gives rise to a semi-Cartan graph, provided that it admits all reflections [Andruskiewitsch et al., 2010,Heckenberger and Schneider, 2010b,Heckenberger and Schneider, 2013a].
+For Hopf algebras, the reflection theory of Nichols algebras has been extensively developed and successfully applied to classify finite-dimensional Nichols algebras of diagonal type [Heckenberger, 2006,Heckenberger, 2010,Angiono, 2013,Angiono, 2015].
 Moreover, the classification of finite-dimensional pointed Hopf algebras over abelian groups has been completed with the aid of the classification of Nichols algebras of diagonal type [Andruskiewitsch and Schneider, 1998,Andruskiewitsch and Schneider, 2000,Andruskiewitsch and Schneider, 2010,Andruskiewitsch et al., 2014,Angiono and García Iglesias, 2019,Andruskiewitsch and Angiono, 2017].
@@
 We now consider two arbitrary Hopf algebras $A$, $B$ in $\mathcal{C}$ with bijective antipodes related by a non-degenerate Hopf pairing.
-Unfortunately, the monoidal categories $^{A}_{A}\mathcal{YD}(\mathcal{C})$} and $^{B}_{B}\mathcal{YD}(\mathcal{C})$} are not monoidal equivalence in general.
+Unfortunately, the monoidal categories $^{A}_{A}\mathcal{YD}(\mathcal{C})$} and $^{B}_{B}\mathcal{YD}(\mathcal{C})$} are not monoidal equivalent in general.
 More precisely, when $A$, $B$ are infinite-dimensional, a module of $A$ in $\mathcal{C}$ cannot give rise to a comodule of $B$ in $\mathcal{C}$.
 1.3 Braided monoidal equivalence induced by a dual pair and its application to reflection theory
-In the present work, we overcome several new difficulties that do not appear in pointed cosemisimple settings and remove the assumption on finite-dimensionality.
+In the present work, we overcome several new difficulties that do not appear in pointed cosemisimple settings and remove the assumption of finite-dimensionality.
 As a result, we extend the reflection theory to arbitrary coquasi‑Hopf algebras $H$ with bijective antipode.
@@
 This allows us to define the category ${{}^{R}_{R}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$, which is proven to be a monoidal subcategory of $^{R}_{R}\mathcal{YD}(\mathcal{C})$}.
-Now for two locally finite Hopf algebras $A$, $B$ in $\mathcal{C}$ with bijective antipodes related by a non-degenerate Hopf pairing compatible with the grading, called a dual pair.
-We prove the following result:
+For two locally finite Hopf algebras $A$, $B$ in $\mathcal{C}$ with bijective antipodes related by a non-degenerate Hopf pairing compatible with the grading, called a dual pair, we prove the following result:
 Theorem 1.1 (Theorem 3.16).
@@
 Then we set $R_{i}(M)=(V_{1},\ldots,V_{\theta})$, where
-$V_{j}=\begin{cases}M_{i}^{*},&\text{if }j=i,\\ (\mathrm{ad}\,M_{i})^{m_{ij}^{M}}(M_{j}),&\text{if }j\neq i.\end{cases}$
-It is still a tuple of finite-dimensional irreducible Yetter-Drinfeld modules. By above theorem, we have a braided monoidal equivalence:
+$V_{j}=\begin{cases}M_{i}^{*},&\text{if }j=i,\\ \mathrm{ad}(M_{i})^{m_{ij}^{M}}(M_{j}),&\text{if }j\neq i.\end{cases}$
+It is still a tuple of finite-dimensional irreducible Yetter-Drinfeld modules. By the above theorem, we have a braided monoidal equivalence:
 $\Omega_{i}:{{}^{\mathcal{B}(M_{i})}_{\mathcal{B}(M_{i})}\mathcal{YD}(\mathcal{\mathcal{C}})}_{\operatorname{rat}}\longrightarrow{{}^{\mathcal{B}(M_{i}^{*})}_{\mathcal{B}(M_{i}^{*})}\mathcal{YD}(\mathcal{\mathcal{C}})}_{\operatorname{rat}}.$
@@
 Theorem 1.2.
-(Theorem 5.12) With above assumptions on $M$. Suppose $M$ admits the $i$-th reflection for $1\leq i\leq\theta$. Then there is an isomorphism of Hopf algebras in $^{H}_{H}\mathcal{YD}$}:
+(Theorem 5.12) Under the above assumptions on $M$. Suppose $M$ admits the $i$-th reflection for $1\leq i\leq\theta$. Then there is an isomorphism of Hopf algebras in $^{H}_{H}\mathcal{YD}$}:
 $\Theta_{i}:\mathcal{B}(R_{i}(M))\cong\Omega_{i}\left(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\right)\#\mathcal{B}(M_{i}^{*}).$ (1.1)
@@
 is a semi-Cartan graph.
-Finally, we apply this machinery to a crucial test case: a rank-three Nichols algebra $\mathcal{B}(M)=\mathcal{B}(M_{1}\oplus M_{2}\oplus M_{3})$ of non-diagonal type, originally identified in [Huang et al., 2026, Proposition 4.1], which is proved to be infinite-dimensional via three different approachs [Huang et al., 2026,Li and Liu, 2025a,Li and Liu, 2025b].
-We compute the Weyl groupoid of $\mathcal{W}(\mathcal{G}(M))$, and show that the set of real roots of $\mathcal{G}(M)$ at each $X\in\mathcal{X}$ is just the set of real roots of Kac-Moody Lie algebra of $\mathcal{A}_{2}^{(1)}$.
-Furthermore, we prove that $\mathcal{G}(M)$ is a Cartan graph.
-Finally, utilizing the correspondence between Cartan graphs and Tits cones[Cuntz, 2018,Cuntz et al., 2019], we demonstrate that this algebra is affine, meaning its Tits cone is a half-plane.
-Our observations shows that affine Nichols algebras can be realized over coquasi-Hopf algebras.
+[…]
 1.4 Organization of this paper
@@
 Section 5 is devoted to the general reflection theory: we study the projection of a Nichols algebra and reflections of Nichols algebras.
-Finally, in Section 6 we work out a detailed example of an affine Nichols algebra and showing that it gives rise to a Cartan graph.
+Finally, in Section 6 we work out a detailed example of an affine Nichols algebra and show that it gives rise to a Cartan graph.
 2 Preliminaries
@@
 We now turn our attention to the Yetter-Drinfeld module category structure.
-The definition of Yetter-Drinfeld module category $^{H}_{H}\mathcal{YD}$} over an arbitrary coquasi Hopf-algebra $H$ was already given in [Ardizzoni and Pavarin, 2013].
+The definition of the Yetter-Drinfeld module category $^{H}_{H}\mathcal{YD}$} over an arbitrary coquasi-Hopf algebra $H$ was already given in [Ardizzoni and Pavarin, 2013].
 Since it plays a crucial role in our paper, we recall this definition.
@@
 A morphism $f:(V,\delta_{V},\rhd)\rightarrow(V^{\prime},\delta_{V^{\prime}},\rhd^{\prime})$ is a colinear map $f:(V,\delta_{V})\rightarrow(V^{\prime},\delta_{V^{\prime}})$ such that $f(h\rhd v)=h\rhd^{\prime}f(v)$ for all $h\in H$.
-The category $^{H}_{H}\mathcal{YD}$} is a $\mathbbm{k}$-linear braided monoidal abelian category over the field $\mathbbm{k}$.
+The category $^{H}_{H}\mathcal{YD}$} is a $\mathbbm{k}$-linear braided monoidal abelian category.
 The unit object of $^{H}_{H}\mathcal{YD}$} is $\mathbbm{k}$, which is regarded as an object in $^{H}_{H}\mathcal{YD}$} via trivial structures.
@@
 The proof for $^{H}_{H}\mathcal{YD}$} is completely dual.
-We need recall the definition of right center here for completeness.
+We need to recall the definition of right center here for completeness.
 Definition 2.4.
@@
 We construct mutually inverse braided monoidal functors between the two categories.
-Let $(V,c_{V,-})$ be an object in $\mathcal{Z}(^{H}\mathcal{M})$, we can define a left linear map on $V$ as follows:
-$\rhd:H\otimes V\longrightarrow V,\qquad h\rhd v=(\mathrm{id}_{V}\otimes\varepsilon)c_{H,V}(h\otimes v)$ (2.7)
+Let $(V,c_{V,-})$ be an object in $\mathcal{Z}(^{H}\mathcal{M})$; we can define a left linear map on $V$ as follows:
+$\rhd:H\otimes V\longrightarrow V,\qquad h\rhd v=(\mathrm{id}_{V}\otimes\varepsilon)c_{H,V}(h\otimes v).$ (2.7)
 Let $W\in{{}^{H}\mathcal{M}}$ and $w^{*}\in W^{*}$.
 Define the map $f_{w^{*}}:W\longrightarrow H,f_{w^{*}}(w)=w_{-1}w^{*}(w_{0})$.
-Then $f_{w^{*}}$ is $H$-comodule map.
+Then $f_{w^{*}}$ is an $H$-comodule map.
 By naturality of $c_{-,V}$, we have
-$c_{H,V}\circ(f_{w^{*}}\otimes\operatorname{id}_{V})$ $=(\operatorname{id}_{V}\otimes f_{w^{*}})\circ c_{W,V}$
+$c_{H,V}\circ(f_{w^{*}}\otimes\operatorname{id}_{V})(w\otimes v)$ $=(\operatorname{id}_{V}\otimes f_{w^{*}})\circ c_{W,V}(w\otimes v)$
 $=c_{H,V}(w_{-1}w^{*}(w_{0})\otimes v).$
+That is
+$c_{H,V}(w_{-1}w^{*}(w_{0})\otimes v)=(\operatorname{id}_{V}\otimes f_{w^{*}})\circ c_{W,V}(w\otimes v).$
 Applying $\operatorname{id}_{V}\otimes\varepsilon$ to both sides leads to
-$(\operatorname{id}_{V}\otimes w^{*})=w*(w_{0})(w_{-1}\rhd v).$
+$w^{*}(w_{0})(w_{-1}\rhd v)=(\operatorname{id}_{V}\otimes\varepsilon\circ f_{w^{*}})\circ c_{W,V}(w\otimes v)=(\operatorname{id}_{V}\otimes w^{*})\circ c_{W,V}(w\otimes v).$
 This implies
@@
 $(r\otimes h)(s\otimes k)$ $=\frac{\Phi(h_{2},s_{-1},k_{2})\Phi(r_{-1},(h_{3}\rhd s_{0})_{-1},h_{5}k_{4})}{\Phi(r_{-2},h_{1},s_{-2}k_{1})\Phi((h_{3}\rhd s_{0})_{-2},h_{4},k_{3})}r_{0}(h_{3}\rhd s_{0})_{0}\otimes h_{6}k_{5},$
-$u_{M}(k)$ $=k1_{R}\otimes 1_{M},$
+$u_{M}(k)$ $=k1_{R}\otimes 1_{H},$
 $\Delta_{M}(r\otimes h)$ $=\Phi^{-1}(r_{-1}^{1},r_{-2}^{2},h_{1})r_{0}^{1}\otimes r_{-1}^{2}h_{2}\otimes r_{0}^{2}\otimes h_{3},$
-$\varepsilon_{M}(r\otimes h)$ $=\varepsilon_{R}(r)\varepsilon_{M}(h),$
+$\varepsilon_{M}(r\otimes h)$ $=\varepsilon_{R}(r)\varepsilon_{H}(h),$
 $\Phi_{M}((r\otimes h),(s\otimes k),(t\otimes l))$ $=\varepsilon_{R}(r)\varepsilon_{R}(s)\varepsilon_{R}(t)\Phi(h,k,l),$
@@
 The antipode $\mathcal{S}$ is bijective if and only if $\mathcal{S}_{H}$ and $\mathcal{S}_{R}$ are bijective.
-Now suppose $N$ are $H$ are both coquasi-Hopf algebras. Furthermore, assume there exists morphisms of coquasi-Hopf algebras
+Now suppose $N$ and $H$ are both coquasi-Hopf algebras. Furthermore, assume there exist morphisms of coquasi-Hopf algebras
 $\pi:N\rightarrow H\ \text{ and }\sigma:H\rightarrow N$
@@
 Example 2.10.
-Given $R,R^{\prime}$ be two Hopf algebras in $^{H}_{H}\mathcal{YD}$} together with a surjective Hopf algebra map $\pi^{\prime}:R^{\prime}\rightarrow R$ and an injective map $\sigma^{\prime}:R\rightarrow R^{\prime}$ such that $\pi^{\prime}\circ\sigma^{\prime}=\operatorname{id}_{R}$.
+Let $R,R^{\prime}$ be two Hopf algebras in $^{H}_{H}\mathcal{YD}$} together with a surjective Hopf algebra map $\pi^{\prime}:R^{\prime}\rightarrow R$ and an injective map $\sigma^{\prime}:R\rightarrow R^{\prime}$ such that $\pi^{\prime}\circ\sigma^{\prime}=\operatorname{id}_{R}$.
 Let $M^{\prime}=R^{\prime}\#H$, and M=R#HM=R\#H.
 Obviously, these maps naturally extend to yield a surjective coquasi-Hopf algebra map $\pi:M^{\prime}\rightarrow M$ and an injective coquasi-Hopf algebra map $\sigma:M\rightarrow M^{\prime}$, such that $\pi\circ\sigma=\operatorname{id}_{M}$.
-Let $K:=M^{\prime\operatorname{co}M}$, then $L$ is an object in $^{M}_{M}\mathcal{YD}$}, with the following structures:
-$\operatorname{ad}:M\otimes K\rightarrow K,h\otimes k\mapsto\Phi_{H}(h_{1}k_{-1},\mathbb{S}_{H}(h_{3})_{1},h_{4})(h_{2}k_{0})\mathbb{S}_{H}(h_{3})_{2},$
-$\operatorname{\rho}:K\rightarrow H\otimes K,k\mapsto\pi(k_{1})\otimes k_{2}.$
+Let $K:=M^{\prime\operatorname{co}M}$, then $K$ is an object in $^{M}_{M}\mathcal{YD}$}, with the following structures:
+$\operatorname{ad}:M\otimes K\rightarrow K,h\otimes k\mapsto\Phi_{M}(h_{1}k_{-1},\mathbb{S}_{M}(h_{3})_{1},h_{4})(h_{2}k_{0})\mathbb{S}_{M}(h_{3})_{2},$
+$\operatorname{\rho}:K\rightarrow M\otimes K,k\mapsto\pi(k_{1})\otimes k_{2}.$
 To analyze the linear map $\operatorname{ad}$ in concrete calculations, the following computation is useful.
-Let $y\in R$ be a primitive element in $^{H}_{H}\mathcal{YD}$}, One can compute the preantipode $\mathbb{S}(y)$ in $M$.
-$\mathbb{S}(y)=\beta*\mathcal{S}*\alpha(y)=\beta(y_{1})\mathcal{S}_{M}(y_{2})\alpha(y_{3})=\beta(y_{-1})\mathcal{S}_{M}(y_{0})=-\beta(y_{-3})\mathcal{S}_{H}(y_{-2})\alpha(y_{-1})y_{0}.$
+Let $y\in R$ be a primitive element.
+One can compute the preantipode $\mathbb{S}_{M}(y)$ in $M$.
+$\mathbb{S}_{M}(y)=\beta*\mathcal{S}_{M}*\alpha(y)=\beta(y_{1})\mathcal{S}_{M}(y_{2})\alpha(y_{3})=\beta(y_{-1})\mathcal{S}_{M}(y_{0})=-\beta(y_{-3})\mathcal{S}_{H}(y_{-2})\alpha(y_{-1})y_{0}.$
 Then we can proceed to calculate $\operatorname{ad}(y)(x)$, where $x\in K$,
-$\operatorname{ad}(y)(x)$ $=yx+(y_{-1}x)\mathbb{S}(y_{0})$ (2.13)
-$=yx+(y_{-4}x)\beta(y_{-3})(\mathcal{S}_{H}(y_{-2})\alpha(y_{-1})y_{0})$
-$=yx-\Phi(y_{-7}x_{-1},\mathcal{S}(y_{-4}),y_{-1})\beta(y_{-6})((y_{-5}x_{0})\mathcal{S}(y_{-3}))\alpha(y_{-2})y_{0}.$
-On the other hand, note that for all $h\in H$, $\Delta_{H}(\mathbb{S}(h))=\beta(h_{1})\mathcal{S}(h_{3})\otimes\mathcal{S}(h_{2})\alpha(h_{4})$. Therefore, we have
-$yx-\operatorname{ad}(y_{-1})(x)y_{0}$ $=yx-\Phi(y_{-4}x_{-1},\mathbb{S}(y_{-2})_{1},y_{-1}))((y_{-3}x_{0})\mathbb{S}(y_{-2})_{2})y_{0}$
-$=yx-\Phi(y_{-7}x_{-1},\mathcal{S}(y_{-4}),y_{-1})\beta(y_{-6})((y_{-5}x_{0})\mathcal{S}(y_{-3}))\alpha(y_{-2})y_{0}.$
+$\operatorname{ad}(y)(x)$ $=yx+(y_{-1}x)\mathbb{S}_{M}(y_{0})$ (2.13)
+$=yx-(y_{-4}x)\beta(y_{-3})(\mathcal{S}_{H}(y_{-2})\alpha(y_{-1})y_{0})$
+$=yx-\Phi(y_{-7}x_{-1},\mathcal{S}_{H}(y_{-3}),y_{-1})\beta(y_{-6})((y_{-5}x_{0})\mathcal{S}_{H}(y_{-4}))\alpha(y_{-2})y_{0}.$
+On the other hand, note that for all $h\in H$, $\Delta_{H}(\mathbb{S}(h))=\beta(h_{1})\mathcal{S}_{H}(h_{3})\otimes\mathcal{S}_{H}(h_{2})\alpha(h_{4})$. Therefore, we have
+$yx-\operatorname{ad}(y_{-1})(x)y_{0}$ $=yx-\Phi(y_{-4}x_{-1},\mathbb{S}_{M}(y_{-2})_{1},y_{-1}))((y_{-3}x_{0})\mathbb{S}_{M}(y_{-2})_{2})y_{0}$
+$=yx-\Phi(y_{-7}x_{-1},\mathcal{S}_{H}(y_{-3}),y_{-1})\beta(y_{-6})((y_{-5}x_{0})\mathcal{S}_{H}(y_{-4}))\alpha(y_{-2})y_{0}.$
 This implies
 $\operatorname{ad}(y)(x)=yx-\operatorname{ad}(y_{-1})(x)y_{0}.$ (2.14)
-2.4 Yetter-Drinfeld modules in arbitrary braided monoidal category
+2.4 Yetter-Drinfeld modules in arbitrary braided monoidal categories
 In Section 3, we will investigate Yetter-Drinfeld modules over a Hopf algebra $A$ in the category $_{H}^{H}\mathcal{YD}$.
@@
 $\overline{c}^{\mathcal{YD}}_{X,Y}=\overline{c}_{X,Y}\circ(\lambda_{X}\otimes\operatorname{id}_{Y})\circ((\mathcal{S}^{-1}_{R}\otimes\operatorname{id}_{X})\otimes\operatorname{id}_{Y})\circ(\overline{c}_{X,R}\otimes\operatorname{id}_{Y})\circ a^{-1}_{X,R,Y}\circ(\operatorname{id}_{X}\otimes\delta_{Y}).$ (2.15)
-By [Heckenberger and Schneider, [2020] ©2020, Proposition 3.4.8], $c^{YD}_{Y,X}$ is an isomorphism in $\mathcal{C}$ with inverse $\overline{c}_{X,Y}^{\mathcal{YD}}$.
-The definition of the bosonization naturally extends to the case that $K$ is a Hopf algebra in the braided monoidal category $^{R}_{R}\mathcal{YD}(\mathcal{C})$}, where $\mathcal{C}$ is now an arbitrary braided monoidal category.
+By [Heckenberger and Schneider, [2020] ©2020, Proposition 3.4.8], $c^{\mathcal{YD}}_{Y,X}$ is an isomorphism in $\mathcal{C}$ with inverse $\overline{c}_{X,Y}^{\mathcal{YD}}$.
+The definition of the bosonization naturally extends to the case where $K$ is a Hopf algebra in the braided monoidal category $^{R}_{R}\mathcal{YD}(\mathcal{C})$}, where $\mathcal{C}$ is now an arbitrary braided monoidal category.
 We collect several key results from [Bespalov, 1997,Bespalov, 1995,Alonso Alvarez and Fernández Vilaboa, 2000] that will be important for the construction.
@@
 That is, we can talk about bosonization and taking coinvariants in $^{H}_{H}\mathcal{YD}$} without any restrictions.
-3 Rational Yetter-Drinfeld module over coquasi-Hopf algebras
+3 Rational Yetter-Drinfeld modules over coquasi-Hopf algebras
 In this section, $H$ always represents a coquasi-Hopf algebra with bijective antipode, and $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$ denotes the Yetter-Drinfeld module category of $H$ .
@@
 However, if $R$ is an algebra in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$, then the category $_{R}\mathcal{C}$ of left $R$-modules in $\mathcal{C}$ is well-defined.
-Motivated by [Heckenberger and Schneider, [2020] ©2020, Definition 12.2.3], we now introduce the concept of rationality under the setting of coquasi-Hopf algebras.
+Motivated by [Heckenberger and Schneider, [2020] ©2020, Definition 12.2.3], we now introduce the concept of rationality in the setting of coquasi-Hopf algebras.
 Definition 3.1.
@@
 (1) A left module $(X,\lambda_{X})$ over $R$ in $\mathcal{C}$ is called rational if for any $x\in X$, there exists a natural number $n_{0}$ such that $\lambda(R(n)\otimes x)=0$ for all $n\geq n_{0}$.
-We denote the category of left rational module over $R$ in $\mathcal{C}$ by ${{}_{R}\mathcal{C}}_{\operatorname{rat}}$.
-(2) A left Yetter-Drinfeld module $X$ over $^{R}_{R}\mathcal{YD}(\mathcal{C})$} is called rational if $X$ is rational as a module of $R$ in $\mathcal{C}$.
+We denote the category of left rational modules over $R$ in $\mathcal{C}$ by ${{}_{R}\mathcal{C}}_{\operatorname{rat}}$.
+(2) A left Yetter-Drinfeld module $X$ in $^{R}_{R}\mathcal{YD}(\mathcal{C})$} is called rational if $X$ is rational as a module of $R$ in $\mathcal{C}$.
 We denote the corresponding category by ${{}^{R}_{R}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$.
 Remark 3.2.
+We say that a $\mathbb{N}_{0}$-graded Hopf algebra $R$ is locally finite if $R(n)$ is finite-dimensional for all $n\geq 0$.
 It is worth noting that for a locally finite $\mathbb{N}_{0}$-graded Hopf algebra $R$, this graded definition of rationality coincides with the classical one.
-Intuitively, the condition $\lambda(R(n)\otimes x)=0$ for large $n$ implies that the action on any element $x$ factors through a finite-dimensional truncation of $R$, which recovers the standard notion of rationality in the non-graded setting.
-A natural question arises naturally, are rational modules closed under tensor products. The following lemma ensures this.
+Intuitively, the condition $\lambda(R(n)\otimes x)=0$ for large $n$ implies that the action on any element $x$ is nontrivial on a finite-dimensional truncation of $R$, which recovers the standard notion of rationality in the non-graded setting.
+A natural question is whether rational modules are closed under tensor products. The following lemma ensures this.
 Lemma 3.3.
@@
 Since $r^{2}_{-3}\in H$, and $V$ is an object in $^{H}_{H}\mathcal{YD}$}, we have $(r^{2}_{-3}\rhd v_{0})_{0}\in V$.
-The expression $\lambda_{V\otimes W}(r\otimes(v\otimes w))$ is a finite sum.
-Thus, by the definition of the rational module, there exists a natural number $n_{1}$ such that for $n\geq n_{1}$,
+The expression $\lambda_{V\otimes W}(r\otimes(v\otimes w))$ is a finite sum. Thus, by the definition of the rational module, there exists a natural number $n_{1}$ such that for $n\geq n_{1}$,
 $\lambda_{V}(R(n)\otimes(r^{2}_{-3}\rhd v_{0})_{0})=0.$
@@
 Therefore, ${{}_{R}\mathcal{C}}_{\operatorname{rat}}$ is a monoidal subcategory of $^{R}_{R}\mathcal{YD}(\mathcal{C})$}.
-The category ${{}_{R}\mathcal{C}}_{\operatorname{rat}}$ is closed under arbitrary direct sums, subobjects and quotient objects.
+The category ${{}_{R}\mathcal{C}}_{\operatorname{rat}}$ is obviously closed under arbitrary direct sums, subobjects and quotient objects.
 (2) The forgetful functor ${{}^{R}_{R}\mathcal{YD}(\mathcal{C})}\rightarrow{{}_{R}\mathcal{C}}$ is monoidal.
@@
 We now turn to the study of pairings in the category $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$, which will later facilitate the construction of dualities between comodules and rational modules.
-Let $A$ and $B$ be objects in $\mathcal{C}$, and $\omega:A\otimes B\rightarrow\mathbbm{k}$ is a morphism in $\mathcal{C}$, named pairing. For subsets $X\subseteq A$ and $Y\subseteq B$, we define
+Let $A$ and $B$ be objects in $\mathcal{C}$, and $\omega:A\otimes B\rightarrow\mathbbm{k}$ is a morphism in $\mathcal{C}$, called a pairing. For subsets $X\subseteq A$ and $Y\subseteq B$, we define
 $X^{\perp}$ $:=\{b\in B\mid\omega(x,b)=0,\ \text{for all}\ x\in X\},$
 $Y^{\perp}$ $:=\{a\in A\mid\omega(a,y)=0,\ \text{for all}\ y\in Y\}.$
-We say $\omega$ is non-degenerate, if $A^{\perp}=0$ and $B^{\perp}=0$.
+We say $\omega$ is non-degenerate if $A^{\perp}=0$ and $B^{\perp}=0$.
 Lemma 3.4.
@@
 Let $E$ be a subobject of $A$, we need to show that $E^{\perp}$ is an object in $\mathcal{C}$.
-The proof of $E^{\perp}$ being a $H$-comodule is the same as in the Hopf algebra case, see [Heckenberger and Schneider, 2013a, Lemma 2.3].
-To show $H\rhd E^{\perp}\subseteq E^{\perp}$. One need to verify that for any $b\in E^{\perp}$, $h\in H$ and $e\in E$,
+The proof of $E^{\perp}$ being an $H$-comodule is the same as in the Hopf algebra case, see [Heckenberger and Schneider, 2013a, Lemma 2.3].
+To show $H\rhd E^{\perp}\subseteq E^{\perp}$. One needs to verify that for any $b\in E^{\perp}$, $h\in H$ and $e\in E$,
 $\omega(e,h\rhd b)=0.$
@@
 $=\sum_{1\leq i\leq n}\Phi(h^{i}_{1},e^{i}_{-1},b_{-1})(h^{i}_{2}\rhd(e_{0}^{i}\otimes b_{0}))\otimes h^{i}_{3}.$
-Here $b_{0}\in E^{\perp}$, because we have shown $E^{\perp}$ is a $H$-comodule.
+Here $b_{0}\in E^{\perp}$, because we have shown $E^{\perp}$ is an $H$-comodule.
 Also, $e_{0}^{i}\in E$ for all $i$.
@@
 Let $f\in\operatorname{Hom}_{C}(V,B\otimes W)$, and $g=G_{V}(f)$.
-For each $v\in V$, there is a finite-dimensional subobject $F\subseteq B$ with $f(v)=\sum_{i\in I}b^{i}\otimes w^{i}\in F\otimes W$.
+By assumption, for each $v\in V$, there is a finite-dimensional subobject $F\subseteq B$ with $f(v)=\sum_{i\in I}b^{i}\otimes w^{i}\in F\otimes W$.
 Now for any $a\in F^{\perp}$, by definition of $G_{V}$ and Lemma 3.4, we have
@@
 $g(F^{\perp}\otimes(h\rhd V^{\prime}))=0.$
-By construction of $\langle V^{\prime}\rangle$, if we denote $V^{\prime}$ as $X_{0}$, this shows $g(F^{\perp}\otimes Y_{0})=0$.
-This property implies that for all $h\in H$, $v\in V^{\prime}$, we have $f(h\rhd v)\in F\otimes W$.
+By construction of $\langle V^{\prime}\rangle$ in Lemma 3.5, if we denote $V^{\prime}$ as $X_{0}$, this shows $g(F^{\perp}\otimes Y_{0})=0$.
 Recall that $X_{1}=\sum_{y\in Y_{0}}\operatorname{span}\{y_{0}\mid\delta_{V}(y)=y_{-1}\otimes y_{0}\},$ We are going to show $g(F^{\perp}\otimes X_{1})=0$.
-This is equivalent to verify $f(X_{1})\in F\otimes W$.
-Since $f$ is a morphism in $\mathcal{C}$, we have
-$(\operatorname{id}\otimes f)\delta(h\rhd v)$ $=\delta(f(h\rhd v)).$
-Recall $F$ is an object in $^{H}_{H}\mathcal{YD}$}, we know that $\delta(f(h\rhd v))\in\delta(F\otimes W)\subseteq H\otimes(F\otimes W)$.
-This implies $f((h\rhd v)_{0})\in F\otimes W.$ Hence $g(F^{\perp}\otimes X_{1})=0$.
-By iterated statements, we can prove that
-$g(F^{\perp}\otimes\langle V\rangle)=0.$
+Since $g$ is a morphism in $\mathcal{C}$, we have
+$(\operatorname{id}_{H}\otimes g)\circ\delta(F^{\perp}\otimes Y_{0})=\delta(g(F^{\perp}\otimes Y_{0})))=0.$
+By definition of $X_{1}$,the set $\operatorname{span}\{z_{0}\mid\delta(z)=z_{-1}\otimes z_{0},\ z\in F^{\perp}\otimes Y_{0}\}=F^{\perp}\otimes X_{1}$.
+This implies $g(F^{\perp}\otimes X_{1})=0$.
+By iterating this argument, we can prove that
+$g(F^{\perp}\otimes\langle V^{\prime}\rangle)=0.$
 Note that we have such a non-degenerate pairing
@@
 For $U_{1},U_{2}\in\mathcal{V}$, let $U_{1}\subseteq U_{2}$, we have $f_{U_{2}}\mid_{U_{1}}=f_{U_{1}}$, since $G_{V}$ is injective.
-Hence the maps $f_{U}$ deine a linear map $f:V\rightarrow B\otimes W$ by $f(v)=f_{U}(v)$, where $U$ is an element in $\mathcal{V}$ containing $v$.
+Hence the maps $f_{U}$ define a linear map $f:V\rightarrow B\otimes W$ by $f(v)=f_{U}(v)$, where $U$ is an element in $\mathcal{V}$ containing $v$.
 This construction yields $G_{V}(f)=g$.
@@
 We begin by introducing the notion of Hopf pairings.
-For our purpose, we assume that $A$ and $B$ are locally finite $\mathbb{N}_{0}$-graded Hopf algebras in $\mathcal{C}$ in this section.
+For our purposes, we assume that $A$ and $B$ are locally finite $\mathbb{N}_{0}$-graded Hopf algebras in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$ in this section.
 Since this does not affect our study of Nichols algebras, we assume that $A$ and $B$ have bijective antipodes in $\mathcal{C}$.
@@
   • (2)
-    If $\omega:A\otimes B\rightarrow\textbf{1}$ is a dual pair, then also
+    If $\omega:A\otimes B\rightarrow\textbf{1}$ is a dual pair, then
     $\omega^{+}:=\omega\circ c_{B,A}\circ(\mathcal{S}_{B}\otimes\mathcal{S}_{A}):B\otimes A\rightarrow\mathbbm{k},$ (3.10)
-    $\omega^{+\operatorname{cop}}:=\omega^{+}\circ(\operatorname{id}_{B}\otimes\mathcal{S}^{-1}_{A}):B^{\operatorname{cop}}\otimes A^{\operatorname{cop}}\rightarrow\mathbbm{k}.$ (3.11)
-The following lemma holds in general boards of monoidal category, here we cite this lemma for simplicity.
+    $\omega^{+\operatorname{cop}}:=\omega^{+}\circ(\operatorname{id}_{B}\otimes\mathcal{S}^{-1}_{A}):B^{\operatorname{cop}}\otimes A^{\operatorname{cop}}\rightarrow\mathbbm{k}$ (3.11)
+    are dual pairs too.
+Recall that $\overline{C}$ means the reverse braided monoidal category. The following lemma holds in general braided monoidal categories; here we cite this lemma for simplicity.
 Lemma 3.9.
@@
 Proposition 3.10.
-Let $(A,B,\omega)$ be a dual pair of locally finite $\mathbb{N}_{0}$-graded Hopf algberas in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$. The functor
+Let $(A,B,\omega)$ be a dual pair of locally finite $\mathbb{N}_{0}$-graded Hopf algebras in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$. The functor
 $D_{1}:\$${{}^{B}\mathcal{C}}\longrightarrow{{}_{A^{\operatorname{cop}}}\overline{\mathcal{C}}_{\operatorname{rat}}},\ \ (V,\delta)\mapsto(V,\lambda),$
 with $\lambda=(A\otimes V\xrightarrow{\operatorname{id}_{A}\otimes\delta}A\otimes(B\otimes V)\xrightarrow{a^{-1}_{A,B,V}}(A\otimes B)\otimes V\xrightarrow{\omega\otimes\operatorname{id}_{V}}V)$
-is a strict monoidal isomorphism of categories.
+is a strict monoidal isomorphism.
 Proof.
@@
 Therefore $\lambda(A(n)\otimes v)=0$ for all $n>n_{0}$.
-Thus if $(V,\lambda)\in{{}_{A}\mathcal{C}}$, then $(V,\lambda)\in{{}_{A}\mathcal{C}_{\operatorname{rat}}}$ automatically.
+Thus if $(V,\lambda)\in{{}_{A}\mathcal{C}}$, where $\lambda\in\operatorname{Hom}_{C,\text{rat}}(A\otimes V,V)$, then $(V,\lambda)\in{{}_{A}\mathcal{C}_{\operatorname{rat}}}$ automatically.
 Now we are going to prove $(V,\lambda)\in{{}_{A}\mathcal{C}}_{\operatorname{rat}}$ if and only if $(V,\delta)\in{{}^{B}\mathcal{C}}$.
@@
 $\circ(\operatorname{id}_{A}\circ a^{-1}_{A,B,B\otimes V})\circ(\operatorname{id}_{A}\otimes(\operatorname{id}_{A}\otimes(\operatorname{id}_{B}\otimes\delta)\circ\delta)).$
-Now we write down explicit formula of $\lambda\circ(\mu_{A}\otimes\operatorname{id}_{V})\circ a_{A,A,V}^{-1}$.
+Now we write down explicit formula of $\lambda\circ(\mu_{A}\otimes\operatorname{id}_{V})\circ a_{A,A,V}^{-1}$
 $\lambda\circ(\mu_{A}\otimes\operatorname{id}_{V})\circ a_{A,A,V}^{-1}=(\omega\otimes\operatorname{id}_{V})\circ a^{-1}_{A,B,V}\circ(\operatorname{id}\otimes\delta)\circ(\mu_{A}\otimes\operatorname{id})\circ a^{-1}_{A,A,V}.$
@@
 $\circ a^{-1}_{A\otimes A,B\otimes B,V}\circ a^{-1}_{A,A,(B\otimes B)\otimes V}\circ(\operatorname{id}_{A}\otimes(\operatorname{id}_{A}\otimes a_{B,B,V}^{-1}))\circ(\operatorname{id}_{A}\otimes(\operatorname{id}_{A}\otimes(a_{B,B,V}\circ((\Delta_{B}\otimes\operatorname{id}_{V})\circ\delta)))).$
-For all $a,a^{\prime}\in A$, $v\in V$, we have $\lambda\circ(\operatorname{id}_{A}\otimes\lambda)(a\otimes(a^{\prime}\otimes v))=\lambda\circ(\mu_{A}\otimes\operatorname{id}_{V})\circ a_{A,A,V}^{-1}(a\otimes(a^{\prime}\otimes v)).$ By non-degeneracy of $\omega$ and Mac-Lane’s coherence theorem, this is equivalent to
+Since we have assumed $(V,\lambda)\in{{}_{A}\mathcal{C}}$, for all $a,a^{\prime}\in A$, $v\in V$, we have $\lambda\circ(\operatorname{id}_{A}\otimes\lambda)(a\otimes(a^{\prime}\otimes v))=\lambda\circ(\mu_{A}\otimes\operatorname{id}_{V})\circ a_{A,A,V}^{-1}(a\otimes(a^{\prime}\otimes v)).$ By non-degeneracy of $\omega$ and Mac-Lane’s coherence theorem, this is equivalent to
 $(\operatorname{id}\otimes\delta)\circ\delta(v)=a_{B,B,V}\circ(\Delta_{B}\otimes\operatorname{id}_{V})\circ\delta(v).$
@@
 3.4 Monoidal equivalence between Yetter-Drinfeld module categories related by a dual pair
-Let $(A,B,\omega)$ be a dual pair of locally finite $\mathbb{N}_{0}$-graded Hopf algberas with bijective antipodes in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$.
+Let $(A,B,\omega)$ be a dual pair of locally finite $\mathbb{N}_{0}$-graded Hopf algebras with bijective antipodes in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$.
 The main result in this subsection is to show there is a braided monoidal equivalence $(\Omega,\beta):{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}\rightarrow{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$.
@@
 $(X,\lambda_{X})=D_{2}(X,\overline{\delta}_{X}),\ (V,\overline{\lambda})=D_{1}(V,\delta).$
-where $D_{1},D_{2}$ is given by Lemma 3.9, then we have
+where $D_{1},D_{2}$ is given by Lemma 3.9, then
 $\overline{c}^{\mathcal{YD}(\mathcal{C})}_{(X,\lambda_{X}),(V,\delta)}=c^{\mathcal{YD}(\overline{\mathcal{C}})}_{(X,\overline{\delta}_{X}),(V,\overline{\lambda})}.$ (3.20)
@@
 $=((p\otimes\operatorname{id}_{V})\circ(\operatorname{id}_{A}\otimes\delta)\otimes\operatorname{id}_{X})\circ(\operatorname{id}_{A}\otimes\overline{c}_{X,V})\circ(\overline{\delta}_{X}\otimes\operatorname{id}_{V})=c^{\mathcal{YD}(\overline{\mathcal{C}})}_{(X,\overline{\delta}_{X}),(V,\overline{\lambda})}.$
-Here the second equality uses (3.9) and (3.11), the third and the sixth equalities use naturality of braiding, the fourth and fifth equalities use braiding axiom. ∎
+Here the second equality uses (3.9) and (3.11), the third and the sixth equalities use the naturality of the braiding, the fourth and the fifth equalities use the braiding axiom. ∎
 Now we return to the setting of coquasi-Hopf algebras.
 We want to restrict the Yetter-Drinfeld module criterion to rational modules.
-Let $B$ be a locally finite $\mathbb{N}_{0}$-graded Hopf algbera in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$ with bijective antipode.
+Let $B$ be a locally finite $\mathbb{N}_{0}$-graded Hopf algebra in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$ with bijective antipode.
 Let $(V,\lambda)\in{{}_{B}\mathcal{C}_{\operatorname{rat}}}$, and $(V,\delta)\in{{}^{B}\mathcal{C}}$.
@@
 $S_{1}=\{(\lambda,\delta)\mid(V,\lambda)\in{{}_{B}\mathcal{C}_{\operatorname{rat}}},(V,\delta)\in{{}^{B}\mathcal{C}}\}$
-$S_{2}=\{(\overline{\lambda},\overline{\delta})\mid(V,\overline{\lambda})\in{{}_{A^{\operatorname{cop}}}\overline{\mathcal{C}}_{\operatorname{rat}}},(V,\overline{\delta})\in{{}^{A^{\operatorname{cop}}}\overline{\mathcal{C}}}\}$
+$S_{2}=\{(\overline{\lambda},\overline{\delta})\mid(V,\overline{\lambda})\in{{}_{A^{\operatorname{cop}}}\overline{\mathcal{C}}_{\operatorname{rat}}},(V,\overline{\delta})\in{{}^{A^{\operatorname{cop}}}\overline{\mathcal{C}}}\}.$
 We define a map $\phi:S_{1}\rightarrow S_{2}$, by $(\lambda,\delta)\mapsto(\overline{\lambda},\overline{\delta})$, where $\overline{\lambda},\overline{\delta}$ are given by
@@
 The construction of the functor $\Omega$ relies on the interplay between the equivalences established in the previous subsections, as illustrated in the following diagram:
-$\overline{{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}$${{}^{A^{\operatorname{cop}}}_{A^{\operatorname{cop}}}\mathcal{YD}(\overline{\mathcal{C}})}_{\operatorname{rat}}$$\overline{{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}$$\overline{{\mathcal{YD}(\mathcal{C})^{A}_{A}}_{,\operatorname{rat}}}$$\scriptstyle{\Gamma}\scriptstyle{\cong}$$\scriptstyle{\Omega}$$\scriptstyle{F_{lr}}$$\scriptstyle{(F_{rl},\alpha)}$
+$\overline{{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}$${{}^{A^{\operatorname{cop}}}_{A^{\operatorname{cop}}}\mathcal{YD}(\overline{\mathcal{C}})}_{\operatorname{rat}}$$\overline{{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}$$\overline{{\mathcal{YD}(\mathcal{C})^{A}_{A}}_{,\operatorname{rat}}}$$\scriptstyle{\Gamma}\scriptstyle{\cong}$$\scriptstyle{\Omega}$$\scriptstyle{F_{lr}}$$\scriptstyle{\overline{F_{rl}}}$
 Theorem 3.16.
@@
 Proof.
-s depicted in the diagram above, we define the functor $\Omega$ as the composition:
+As depicted in the diagram above, we define the functor $\Omega$ as the composition:
 $\Omega:=F_{rl}\circ F_{lr}\circ\Gamma.$
-We now compute the explicit structure of this compositionLet $(V,\lambda_{B},\delta_{B})\in{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$. Then
+We now compute the explicit structure of this composition. Let $(V,\lambda_{B},\delta_{B})\in{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$. Then
 $F_{rl}F_{lr}\Gamma(V,\lambda_{B},\delta_{B})=$ $F_{rl}F_{lr}(V,\overline{\lambda_{A}},\overline{\delta_{A}})$
@@
 Thus $(\Omega,\overline{\beta}):\overline{{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}\rightarrow\overline{{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}$ is a braided monoidal equivalence.
-Now we take reverse braiding, and it is direct that
-$(\Omega,\beta):{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}\rightarrow{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}},\ \beta_{X,Y}=c_{Y,X}^{{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}}\circ\overline{c}_{X,Y}.$
+Now we take the reverse braiding, and it is direct that
+$(\Omega,\beta):{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}\rightarrow{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}},\ \beta_{X,Y}=c_{Y,X}^{{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}}\circ\overline{c}_{X,Y}$
 is a braided monoidal equivalence.
@@
 We first recall the definition of Nichols algebras $\mathcal{B}(V)$ in $\mathcal{C}$, where $V\in\mathcal{C}$ is an arbitrary object.
-Then we are going to show that there is a non-degenerate Hopf pairing between $\mathcal{B}(V^{*})$ and $\mathcal{B}(V)$ when $V$ has a left dual $V^{*}$.
+Then we will show that there is a non-degenerate Hopf pairing between $\mathcal{B}(V^{*})$ and $\mathcal{B}(V)$ when $V$ has a left dual $V^{*}$.
 We denote by $T(V)$ the tensor algebra in $\mathcal{C}$ generated freely by $V$. Here we denote that
@@
 $\operatorname{Wor}(c)=\bigoplus_{n\geq 1}i_{n}[n]!_{c},$
-where $n]!_{c}\in\operatorname{End}(V^{\otimes n})$ is the braided symmertriser. For explicit definition of $n]!_{c}$, one may refer to [Bazlov and Berenstein, 2013, Section 5.6].
+where $n]!_{c}\in\operatorname{End}(V^{\otimes n})$ is the braided symmertriser. For the explicit definition of $n]!_{c}$, one may refer to [Bazlov and Berenstein, 2013, Section 5.6].
 Definition 4.1.
@@
 There is a unique Hopf pairing $\omega:T(V^{*})\otimes T(V)\to\textbf{1}$.
-It is shown in [Majid, 1993] that $\omega$ is given by
+It is shown in [Bazlov and Berenstein, 2013] that $\omega$ is given by
 $\omega\mid_{V^{*}\otimes V}=\operatorname{ev}_{V}.$
 and
-$\omega_{n}:=\omega|_{V^{*\otimes n}\otimes V^{\otimes n}}=\operatorname{ev}_{V^{\otimes n}}\circ(\operatorname{id}_{V^{*\otimes n}}\otimes[n]!_{c})=\operatorname{ev}_{V^{\otimes n}}\circ([n]!_{c^{*}}\otimes\operatorname{id}_{V^{\otimes n}})$
-The Hopf pairing between $V^{\otimes n}$ and $V^{*\otimes m}$ is 0 unless $n=m$. Here $c^{*}$ denotes the the braided symmertriser of $V^{*}$.
-The Hopf pairing between $T(V)$ and $T(V^{*})$ inducing a Hopf pairing between $\mathcal{B}(V^{*})$ and $\mathcal{B}(V)$.
+$\omega_{n}:=\omega|_{V^{*\otimes n}\otimes V^{\otimes n}}=\operatorname{ev}_{V^{\otimes n}}\circ(\operatorname{id}_{V^{*\otimes n}}\otimes[n]!_{c})=\operatorname{ev}_{V^{\otimes n}}\circ([n]!_{c^{*}}\otimes\operatorname{id}_{V^{\otimes n}}).$
+The Hopf pairing between $V^{\otimes n}$ and $V^{*\otimes m}$ is 0 unless $n=m$. Here $c^{*}$ denotes the braided symmetriser of $V^{*}$.
+The Hopf pairing between $T(V)$ and $T(V^{*})$ induces a Hopf pairing between $\mathcal{B}(V^{*})$ and $\mathcal{B}(V)$.
 Note that the Hopf pairing $\omega$ vanishes on the kernels of the symmetrisers, and for all positive integer $n$, we have
 $\omega_{n}({\operatorname{ker}[n]!_{c}^{*},\ V^{\otimes n}})=0,\ \omega_{n}({V^{*\otimes n},\ \operatorname{ker}[n]!_{c}})=0.$
-Now we return to the case of coquasi-Hopf algebra $H$ with bijective antipode.
+Now we return to the case of a coquasi-Hopf algebra $H$ with bijective antipode.
 Let $V$ be a finite-dimensional object in $^{H}_{H}\mathcal{YD}$}.
@@
 This follows immediately from the existence of a dual pair $\omega:\mathcal{B}(V^{*})\otimes\mathcal{B}(V)\rightarrow\mathbbm{k}$ and Theorem 3.16
-Let $A$, $B$ be $\mathbb{N}_{0}$-graded locally finite Hopf algebra in $\mathcal{C}$ with bijective antipodes and $\omega:A\otimes B\rightarrow\mathbbm{k}$ a dual pair.
+Let $A$, $B$ be $\mathbb{N}_{0}$-graded locally finite Hopf algebra in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$ with bijective antipodes and $\omega:A\otimes B\rightarrow\mathbbm{k}$ a dual pair.
 Recall the braided monoidal equivalence in Theorem 3.16:
@@
 If $V$ is rational, then for any $n\in\mathbb{N}_{\geq 1}$, $V^{\otimes n}$ is rational in $^{B}_{B}\mathcal{YD}(\mathcal{C})$}.
-Since ${{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$ is closed under direct sum and quotient by Lemma 3.3 both objects $T(V)=\bigoplus_{n\geq 0}V^{\otimes n}$ and $\mathcal{B}(V)$ are rational.
+Since ${{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$ is closed under direct sum and quotient by Lemma 3.3.
+Both objects $T(V)=\bigoplus_{n\geq 0}V^{\otimes n}$ and $\mathcal{B}(V)$ are rational.
 ∎
@@
 4.2 The functor $\Omega$ under $\mathbb{Z}$-gradings
-[…]
+Having established the duality for Nichols algebras, we now investigate how the functor $\Omega$ interacts with gradings.
+Let $\mathbb{N}_{0}\text{-Gr}\mathcal{M}_{\mathbbm{k}}$(resp., -$\mathbb{Z}\text{-Gr}\mathcal{M}_{\mathbbm{k}}$) be the category of $\mathbb{N}_{0}$-graded vector spaces(resp., $\mathbb{Z}$-graded vector spaces).
+The functor $\mathbb{N}_{0}\text{-Gr}\mathcal{M}_{k}\rightarrow\mathbb{Z}\text{-Gr}\mathcal{M}_{k}$, which extends the $\mathbb{N}_{0}$-grading of an object $V$ in $\mathbb{N}_{0}\text{-Gr}\mathcal{M}_{k}$ to a $\mathbb{Z}$-grading by setting $V(n)=0$ for all $n<0$, is monoidal.
+This functor allows us to view $\mathbb{N}_{0}$-graded coalgebras and coquasi-Hopf algebras as $\mathbb{Z}$-graded coalgebras and $\mathbb{Z}$-graded coquasi-Hopf algebras, respectively.
 Furthermore, suppose $H$ is a $\mathbb{N}_{0}$-graded coquasi-Hopf algebra.
 A $\mathbb{Z}$-graded Yetter-Drinfeld module $V$ over $H$ is by definition an object $V$ in ${{}^{H}_{H}\mathcal{YD}}(\mathbb{Z}\text{-Gr}\mathcal{M}_{k})$.
-In other words, $V$ is a $\mathbb{Z}$-graded vector space such that the map $H\otimes V\rightarrow V$ and comodule structure maps $V\rightarrow H\otimes V$ are $\mathbb{Z}$-graded.
+In other words, $V$ is a $\mathbb{Z}$-graded vector space such that the map $H\otimes V\rightarrow V$ and the comodule structure maps $V\rightarrow H\otimes V$ are $\mathbb{Z}$-graded.
 Lemma 4.6.
@@
 Here, the grading of $\mathbbm{k}$ is given by $\mathbbm{k}(n)=0$, if $n\neq 0$, and $\mathbbm{k}(0)=\mathbbm{k}$.
-Now we return to the purpose of this subsection. We first introduce the following notations for further use.
+Now we return to the purpose of this subsection. We first introduce the following notation for further use.
 Definition 4.7.
@@
 (1) Recall
-$\lambda_{\mathcal{B}(N^{*})}=(\mathcal{B}(N^{*})\otimes V\xrightarrow{\operatorname{id}\otimes\delta_{\mathcal{B}(N)}}\mathcal{B}(N^{*})\otimes(\mathcal{B}(N)\otimes V)\xrightarrow{a^{-1}_{\mathcal{B}(N^{*}),\mathcal{B}(N),V}}(\mathcal{B}(N^{*})\otimes\mathcal{B}(N))\otimes V\xrightarrow{\omega\otimes\operatorname{id}_{V}}\mathcal{B}(N^{*})),$
+$\lambda_{\mathcal{B}(N^{*})}=(\mathcal{B}(N^{*})\otimes V\xrightarrow{\operatorname{id}\otimes\delta_{\mathcal{B}(N)}}\mathcal{B}(N^{*})\otimes(\mathcal{B}(N)\otimes V)\xrightarrow{a^{-1}_{\mathcal{B}(N^{*}),\mathcal{B}(N),V}}(\mathcal{B}(N^{*})\otimes\mathcal{B}(N))\otimes V\xrightarrow{\omega\otimes\operatorname{id}_{V}}V),$
 Now for fixed $n\geq 0$, the kernel of the induced map:
@@
 On the other hand, if $v\in\mathcal{F}_{n}(V,\lambda_{\mathcal{B}(N^{*})})$, this will imply $\delta(v)\in\bigoplus_{i=0}^{n}\mathcal{B}(N)(i)\otimes V$.
-which completes the proof of (1): $\mathcal{F}_{n}\Omega(V)=\mathcal{F}_{n}(V,\lambda_{\mathcal{B}(N^{*})})=\mathcal{F}^{n}V$.
+This completes the proof of (1): $\mathcal{F}_{n}\Omega(V)=\mathcal{F}_{n}(V,\lambda_{\mathcal{B}(N^{*})})=\mathcal{F}^{n}V$.
 (2) For the second statement, recall $\delta_{\mathcal{B}(N^{*})}$ is given by
@@
 The above two equations imply that $\Omega(V)$ is a well-defined $\mathbb{Z}$-graded object in $^{\mathcal{B}(N^{*})}_{\mathcal{B}(N^{*})}\mathcal{YD}(\mathcal{C})$}. ∎
-5 Reflection of Nichols algebras over coquasi-Hopf algebra with bijective antipode
+5 Reflection of Nichols algebras over coquasi-Hopf algebras with bijective antipode
 Let $H$ be a coquasi-Hopf algebra with bijective antipode.
 In this section, we study projection of Nichols algebras in $^{H}_{H}\mathcal{YD}$}, which will play an important role when considering reflection of simple Yetter-Drinfeld modules.
-Some proofs in this section are parallel to those in the pointed coquasi-Hopf algebra setting Li and Liu (2025b), we will omit them for simplicity.
+Some proofs in this section are parallel to those in the pointed coquasi-Hopf algebra setting Li and Liu (2025b); therefore we will omit them for simplicity.
 5.1 The structure of the space of coinvariants $K$
@@
 There is a natural injection $\iota:\mathcal{A}(N)\rightarrow\mathcal{A}(M\oplus N)$ such that $\pi\circ\iota=\operatorname{id}_{\mathcal{A}(N)}$.
-Denote $\pi_{H}:\mathcal{A}(M\oplus N)\rightarrow H$, and $\pi^{\prime}_{H}:\mathcal{A}(N)\rightarrow H$ be the natural projection.
+Let $\pi_{H}:\mathcal{A}(M\oplus N)\rightarrow H$, and $\pi^{\prime}_{H}:\mathcal{A}(N)\rightarrow H$ be the natural projection.
 These maps satisfy a compatibility condition, namely, there exists a commutative diagram relating these projections:
@@
 $\operatorname{ad}:\mathcal{A}(N)\otimes K\rightarrow K,\ \ \ \ a\otimes x\mapsto\operatorname{ad}(a)(x)=\Phi(a_{1}x_{-1},\mathbb{S}(a_{3})_{1},a_{4})(a_{2}x_{0})\mathbb{S}(a_{3})_{2}.$ (5.1)
-Here, $\mathbb{S}$ denotes preantipode of $\mathcal{A}(N)$. We have the following two observations.
+Here, $\mathbb{S}$ denotes the preantipode of $\mathcal{A}(N)$. We have the following two observations.
 Lemma 5.1.
@@
 Moreover, we have $\operatorname{ad}(N)^{n}(M)\cap\operatorname{ad}(N)^{m}(M)=0$ for $n\neq m$.
-This is implies the direct sum decomposition of $L$.
+This implies the direct sum decomposition of $L$.
 Since $M\subseteq K$ and $K\in{{}^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}}$, we conclude that $L=\operatorname{ad}(\mathcal{B}(N))(M)\subseteq K$. Note that
@@
 Now we assume that for some fixed integer $n_{0}$, we have $\delta_{K}(\operatorname{ad}(a)(x))\in\mathcal{A}(N)\otimes L$ for all $x\in\operatorname{ad}(N)^{n^{\prime}}(M)$, where $n^{\prime}\leq n_{0}$.
-Then for $a\in N$ and $x\in\operatorname{ad}(N)^{n_{0}+1}(M)$.
-Since by definition, $\operatorname{ad}(N)^{n_{0}+1}(M)=\operatorname{ad}(N)(\operatorname{ad}(N)^{n_{0}}(M))$.
-We have $\delta_{K}(x)\in\mathcal{A}(N)\otimes L$ by inductive hypothesis.
+Then for $a\in N$ and $x\in\operatorname{ad}(N)^{n_{0}+1}(M)$, since by definition $\operatorname{ad}(N)^{n_{0}+1}(M)=\operatorname{ad}(N)(\operatorname{ad}(N)^{n_{0}}(M))$, we have $\delta_{K}(x)\in\mathcal{A}(N)\otimes L$ by the inductive hypothesis.
 Thus $\operatorname{ad}(a_{0})(x_{0})\in\operatorname{ad}(N)(L)\subseteq L$.
@@
 Since $\operatorname{ad}(a_{-2})(x_{0})\in\bigoplus_{i=0}^{n_{0}+1}\operatorname{ad}(N)^{i}(M)$, we have $(\operatorname{ad}(a_{-2})(x_{0}))_{0}\in L$ by inductive hypothesis.
-This completes the induction, and we shows $\delta_{K}:L\rightarrow\mathcal{A}(N)\otimes L$ is well-defined.
+This completes the induction, and we show $\delta_{K}:L\rightarrow\mathcal{A}(N)\otimes L$ is well-defined.
 The third axiom holds automatically since $L\subseteq K$.
@@
 The term $\operatorname{ad}(a)(x)\otimes 1$ will occur in $\Delta_{\mathcal{B}(M\oplus N)}(\operatorname{ad}(a)(x))$.
-Now the left terms are
+Now the remaining terms are
 $\sum_{x^{1}\otimes x^{2}\neq x\otimes 1}\Phi(a_{-1},x^{1}_{-1},x^{2}_{-1})a_{0}x^{1}_{0}\otimes x^{2}_{0}+\frac{\Phi(a_{-3},x^{1}_{-1},x^{2}_{-2})}{\Phi((\operatorname{ad}(a_{-2})(x^{1}_{0}))_{-1},a_{-1},x^{2}_{-1})}(\operatorname{ad}(a_{-2})(x^{1}_{0}))_{0}\otimes a_{0}x^{2}_{0}$
@@
 $\sum_{x^{1}\otimes x^{2}\neq x\otimes 1}\frac{\Phi(a_{-3},x^{1}_{-1},x^{2}_{-2})}{\Phi((\operatorname{ad}(a_{-2})(x^{1}_{0}))_{-1},a_{-1},x^{2}_{-1})}(\operatorname{ad}(a_{-2})(x^{1}_{0}))_{0}\otimes\operatorname{ad}(a_{-1})(x_{0}^{2})a_{0}.$
-Now we combine the second and third term together, yields
+Combining the second and third terms together yields
 $\sum_{x^{1}\otimes x^{2}\neq x\otimes 1}\frac{\Phi(a_{-3},x^{1}_{-1},x^{2}_{-2})}{\Phi((\operatorname{ad}(a_{-2})(x^{1}_{0}))_{-1},a_{-1},x^{2}_{-1})}(\operatorname{ad}(a_{-2})(x^{1}_{0}))_{0}\otimes\operatorname{ad}(a_{0})(x^{2}_{0}).$
@@
 Now we are going to show $K$ is a Nichols algebra by verifying the axiom in Definition 4.2.
-(1) It follows direct that $K$ inherits an $\mathbb{N}_{0}$-graded Hopf algebra in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} with $K(n)=K\cap(\mathcal{B}(M\oplus N)\#H)(n)$.
+(1) It follows directly that $K$ inherits an $\mathbb{N}_{0}$-graded Hopf algebra in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} with $K(n)=K\cap(\mathcal{B}(M\oplus N)\#H)(n)$.
 (2) We have $K(0)=K\cap(\mathcal{B}(M\oplus N)\#H)(0)=K\cap(\mathcal{B}(M\oplus N)(0)=\mathbbm{k}$.
-(3) Note that $K(1)=K\cap(\mathcal{B}(M\oplus N)\#H)(1)=K\cap(\mathcal{B}(M\oplus N)(1))=\operatorname{ad}(\mathcal{B}(N))(M)=L$.
-We are going to show $K$ is generated its degree one component $L$.
+(3) We are going to show $K$ is generated by $L$.
 Let $K^{\prime}$ be the subalgebra of $K$ generated by $L$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}, which is an object in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} with structures induced by tensor product of $L$.
@@
 Thus $W$ is a subalgebra of $\mathcal{B}(M\oplus N)$, which establishes that $K$ is generated by $L$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}.
+Note that $L\subseteq K(1)$, this forces $K(1)=L$.
+Hence $K$ is generated by $K(1)$.
 (4) To show $P(K)=K(1)$, that is, there is no primitive element in $K(n)$, $n\geq 2$.
@@
 By [Li and Liu, 2025b, Lemma 4.12], $M$ is an object in $^{H}_{H}\mathcal{YD}$} with decomposition into irreducible objects $M=\bigoplus_{i\in I}M_{i}:=\bigoplus_{i\in I}M\cap W_{i}$.
+Here $M_{i}$ was proven to be the $\mathbb{Z}$-homogeneous component of $W_{i}$ of minimal degree.
 Moreover, we have $W_{i}=\mathcal{B}(N)\rhd M_{i}=\bigoplus_{n\in\mathbb{N}_{0}}N^{\otimes n}\rhd M_{i}$ for each $i\in I$.
-Let $\operatorname{deg}(M_{i})=m_{i}$ for each $i\in I$. We endow $W_{i}$ a new grading by $\widetilde{W_{i}}=W_{i}$ with
+Denote $\operatorname{deg}(M_{i})=m_{i}$ for each $i\in I$. We endow $W_{i}$ with a new grading by $\widetilde{W_{i}}=W_{i}$ with
 $\widetilde{W_{i}}(n)=W(n+m_{i}-1)=N^{n-1}\rhd M_{i},\ \text{for all}\ n\in\mathbb{N}_{0}.$
@@
 $R:=\mathcal{B}(W)\#\mathcal{B}(N)=(\mathcal{B}(W)\#(\mathcal{B}(N)\#H))^{\operatorname{co}H}$
-is a $\mathbb{N}_{0}$-graded Hopf algebra in $^{H}_{H}\mathcal{YD}$}. Its degree 0’s part is $\mathbbm{k}$, and its degree $1$’s part is $M\oplus N$.
+is an $\mathbb{N}_{0}$-graded Hopf algebra in $^{H}_{H}\mathcal{YD}$}. Its degree 0 part is $\mathbbm{k}$, and its degree $1$ part is $M\oplus N$.
 We are going to show $R$ is a pre-Nichols algebra in $^{H}_{H}\mathcal{YD}$}, it remains to show $R$ is generated by $M\oplus N$ as an algebra in $^{H}_{H}\mathcal{YD}$}.
@@
 This implies terms like $(y\rhd x)_{0}$ are contained in the subalgebra generated by $\mathcal{B}(N)$ and $M$.
-Let us fixed $k\geq 0$, for all $z\in\mathcal{B}(N)(k^{\prime})$ and $x\in M$, where $k^{\prime}\geq k$.
+Let us fix $k\geq 0$, for all $z\in\mathcal{B}(N)(k^{\prime})$ and $x\in M$, where $k^{\prime}\geq k$.
 We assume the element $z\rhd x$ as well as $(z\rhd x)_{0}$ lie in the subalgebra generated by $\mathcal{B}(N)$ and $M$, then for homogeneous $y\in N$,
@@
 5.2 The semi-Cartan graph of a Nichols algebra
-In this subsection, we develop reflections of Nichols algebras over arbitrary coquasi-Hopf algebra with bijective antipode.
+In this subsection, we develop reflections of Nichols algebras over arbitrary coquasi-Hopf algebras with bijective antipode.
 The reflection allows us to systematically relate different Nichols algebra realizations through a well-defined transformation procedure.
@@
 Two tuples $M,M^{\prime}\in\mathcal{F}_{\theta}$ are called isomorphic, denoted $M\cong M^{\prime}$, if $M_{j}\cong M_{j}^{\prime}$ in $^{H}_{H}\mathcal{YD}$} for all j.
-The isomorphism class of $M\in\mathcal{F}_{\theta}$ is denote by $[M]$.
+The isomorphism class of $M\in\mathcal{F}_{\theta}$ is denoted by $[M]$.
 For $1\leq i\leq\theta$ and $M\in\mathcal{F}_{\theta}$, we say the tuple $M$ admits the $i$-th reflection $R_{i}(M)$ if for all $j\neq i$ there is a natural number $m_{ij}^{M}\geq 0$ such that $(\mathrm{ad}\,M_{i})^{m_{ij}^{M}}(M_{j})$ is a non-zero finite-dimensional subspace of $\mathcal{B}(M)$, and $(\mathrm{ad}\,M_{i})^{m_{ij}^{M}+1}(M_{j})=0$.
@@
 Then we set $R_{i}(M)=(V_{1},\ldots,V_{\theta})$, where
-$V_{j}=\begin{cases}M_{i}^{*},&\text{if }j=i,\\ (\mathrm{ad}\,M_{i})^{m_{ij}^{M}}(M_{j}),&\text{if }j\neq i.\end{cases}$
+$V_{j}=\begin{cases}M_{i}^{*},&\text{if }j=i,\\ \mathrm{ad}(M_{i})^{m_{ij}^{M}}(M_{j}),&\text{if }j\neq i.\end{cases}$
 Such a question arises naturally: how do the irreducibility properties behave under reflections? The following result provides a reassuring answer.
@@
 The object $\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})$ belongs to $^{H}_{H}\mathcal{YD}$}.
-Now suppose $0\neq P\subset\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})$ is a Yetter-Drinfeld submodule over $^{H}_{H}\mathcal{YD}$}.
+Moreover, it is obvious that $\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})$ is generated by $\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})$ as a $\mathcal{A}(M_{i})$-comodule.
+Now suppose $0\neq P\subset\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})$ is a Yetter-Drinfeld submodule in $^{H}_{H}\mathcal{YD}$}.
 Let $\langle P\rangle$ be the $\mathcal{A}(M_{i})$-subcomodule of $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ generated by $P$, defined explicitly as
 $\langle P\rangle:=\{\langle f,x_{-1}\rangle x_{0}\mid x\in P,f\in\operatorname{Hom}(\mathcal{A}(M_{i}),\mathbbm{k})\}.$
-To establish this, we prove $\langle P\rangle$ is a Yetter-Drinfeld submodule of $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$}.
+We want to prove $\langle P\rangle$ is a Yetter-Drinfeld submodule of $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$}.
 To see this, for all $a\in\mathcal{A}(M_{i})$ and $x\in P$, by Lemma 2.6,
@@
 Hence $\operatorname{ad}(a)(\langle f,x_{-1}\rangle x_{0})\in\langle P\rangle.$ This ensures the first axiom of a Yetter-Drinfeld module, and the third axiom holds automatically.
-Since $\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})$ is generated by $\operatorname{ad}(M_{i})^{m_{ij}}(M_{j})$ as a Yetter-Drinfeld module and is irreducible in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$} by Lemma 5.4 (2).
-Then $\langle P\rangle=\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})$ since $P\neq 0$.
-This forces $P=\operatorname{ad}(M_{i})^{m_{ij}}(M_{j}),$ establishing the irreducibility.
+Therefore $\langle P\rangle$ is a Yetter-Drinfeld submodule of $\operatorname{ad}\mathcal{B}(M_{i})(M_{j})$ in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$}.
+Since $\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})$ is irreducible in $^{A(M_{i})}_{A(M_{i})}\mathcal{YD}$} by Lemma 5.4 (2), and $P\neq 0$, we have $\langle P\rangle=\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})$.
+It is direct to see that $\langle P\rangle=\bigoplus_{i=0}^{m_{ij}}\langle P\rangle\cap\mathcal{B}^{i}(M_{i}\oplus M_{j})$.
+Then
+$P=\langle P\rangle\cap\mathcal{B}^{m_{ij}}(M_{i}\oplus M_{j})=\operatorname{ad}(\mathcal{B}(M_{i}))(M_{j})\cap\mathcal{B}^{m_{ij}}(M_{i}\oplus M_{j})=\operatorname{ad}(M_{i})^{m_{ij}}(M_{j}).$
+This establishes the irreducibility. ∎
 The following lemma is standard, we list here for completeness.
@@
 Then $(a_{ij}^{M})_{i,j\in\mathbb{I}}$ is a generalized Cartan matrix.
-(2) Suppose $M\cong N$ in $\mathcal{F}_{\theta}$, if $M$ admits the $i$-th reflection for some $i\in\mathbb{I}$, so is $N$.
+(2) Suppose $M\cong N$ in $\mathcal{F}_{\theta}$, if $M$ admits the $i$-th reflection for some $i\in\mathbb{I}$, so does $N$.
 Furthermore, $R_{i}(M)\cong R_{i}(N)$ and $a_{ij}^{M}=a_{ij}^{N}$ for each $j\in\mathbb{I}$.
-Recall by Remark 2.12, let $A$ be a Hopf algebra in $^{H}_{H}\mathcal{YD}$} with bijective antipode, we have such a braided monoidal isomorphism
+Recall from Remark 2.12, let $A$ be a Hopf algebra in $^{H}_{H}\mathcal{YD}$} with bijective antipode, we have such a braided monoidal isomorphism
 $F:{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}\cong{{}^{A\#H}_{A\#H}\mathcal{YD}}.$
 Now restricting to rational Yetter-Drinfeld modules, we denote the image of $F$ by ${{}^{A\#H}_{A\#H}\mathcal{YD}}_{\operatorname{rat}}$, which is a monoidal full subcategory of $^{A\#H}_{A\#H}\mathcal{YD}$.
-From now on, we always fix a tuple $M=(M_{1},\ldots,M_{\theta})\in\mathcal{F}_{\theta}$, with each component $M_{i}$ is irreducible for $i\in\mathbb{I}$.
+From now on, we always fix a tuple $M=(M_{1},\ldots,M_{\theta})\in\mathcal{F}_{\theta}$, with each component $M_{i}$ being irreducible for $i\in\mathbb{I}$.
 By Corollary 4.3, we have such a braided tensor equivalence for each $i\in\mathbb{I}$.
@@
 Lemma 5.11.
-With above assumptions on $M$, and suppose $M$ admits the $i$-th reflection with $R_{i}(M)=(V_{1},\ldots,M_{i}^{*},\ldots,V_{\theta})$, then the following equalities hold.
+Under the above assumptions on $M$, and suppose $M$ admits the $i$-th reflection with $R_{i}(M)=(V_{1},\ldots,M_{i}^{*},\ldots,V_{\theta})$, then the following equalities hold.
 $\operatorname{ad}(\mathcal{B}(M_{i})(M_{j}))=\bigoplus_{n=0}^{m_{ij}}\operatorname{ad}(M_{i})^{n}(M_{j})$ $,\ \ \ \ V_{j}=\mathcal{F}_{0}(\operatorname{ad}\mathcal{B}(M_{i})(M_{j})),$
@@
 The next theorem gives a natural explanation of reflections of tuples of Yetter-Drinfeld modules.
-Although the proof strategy parallels the group case [Li and Liu, 2025b, Theorem 5.7], we include the details here for the sake of completeness and the reader’s convenience.
+Although the proof strategy parallels that of group case [Li and Liu, 2025b, Theorem 5.7], we include the details here for the sake of completeness and the reader’s convenience.
 Theorem 5.12.
-With above assumptions on $M$, and suppose $M$ admits the $i$-th reflection, then there is an isomorphism of Hopf algebras in $^{H}_{H}\mathcal{YD}$}:
+Under the above assumptions on $M$, and suppose $M$ admits the $i$-th reflection, then there is an isomorphism of Hopf algebras in $^{H}_{H}\mathcal{YD}$}:
 $\Theta_{i}:\mathcal{B}(R_{i}(M))\cong\Omega_{i}\left(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\right)\#\mathcal{B}(M_{i}^{*}).$ (5.5)
@@
 Definition 5.13.
-Let $M\in\mathcal{F}_{\theta}$, with each $M_{i}$ is irreducible for $i\in I$. For $l\in\mathbb{N}_{0}$ and $i_{1},i_{2},...,i_{l}\in\mathbb{I}$.
-$(1)$ We say $M$ admits the reflection sequence $(i_{1},i_{2},...,i_{l})$ if $l=0$ or $M$ amdits the $i_{1}$-th reflection and $R_{i_{1}}(M)$ satisfies the reflection sequence $(i_{2},i_{3},...,i_{l})$.
-$(2)$ We say $M$ admits all reflection sequences if $M$ admits all reflection sequence $(i_{1},i_{2},...,i_{l})$ for all $l\in\mathbb{N}_{0}$ and $i_{1},i_{2},...,i_{l}\in\mathbb{I}$.
+Let $M\in\mathcal{F}_{\theta}$, with each $M_{i}$ is irreducible for $i\in\mathbb{I}$. For $l\in\mathbb{N}_{0}$ and $i_{1},i_{2},...,i_{l}\in\mathbb{I}$.
+$(1)$ We say $M$ admits the reflection sequence $(i_{1},i_{2},...,i_{l})$ if $l=0$ or $M$ admits the $i_{1}$-th reflection and $R_{i_{1}}(M)$ satisfies the reflection sequence $(i_{2},i_{3},...,i_{l})$.
+$(2)$ We say $M$ admits all reflection sequences if $M$ admits reflection sequence $(i_{1},i_{2},...,i_{l})$ for all $l\in\mathbb{N}_{0}$ and $i_{1},i_{2},...,i_{l}\in\mathbb{I}$.
 For $M\in\mathcal{F}_{\theta}$ admitting all reflections, we denote
@@
 The quadruple $\mathcal{G}=\mathcal{G}(\mathbb{I},\mathcal{X},(r_{i})_{i\in I},(A^{X})_{X\in\mathcal{X}})$ is called a semi-Cartan graph if for all $X\in\mathcal{X}$, the matrix $A^{X}$ is a generalized Cartan matrix, and the following axioms hold.
-$(CG1)$ For all $i\in I$, the map $r_{i}$ satisfies $r_{i}^{2}=\mathrm{id}_{\mathcal{X}}$.
-$(CG2)$ For all $i\in I$, $X\in\mathcal{X}$, $A^{X}$ and $A^{r_{i}(X)}$ have the same $i$-th row.
+$(CG1)$ For all $i\in\mathbb{I}$, the map $r_{i}$ satisfies $r_{i}^{2}=\mathrm{id}_{\mathcal{X}}$.
+$(CG2)$ For all $i\in\mathbb{I}$, $X\in\mathcal{X}$, $A^{X}$ and $A^{r_{i}(X)}$ have the same $i$-th row.
 As a consequence of Theorem 5.12, we obtain the following corollary.
@@
 Let $M=(M_{1},\ldots,M_{\theta})\in\mathcal{F}_{\theta}$ with each component $M_{i}$ is irreducible.
-(1) For each $i\in\mathbb{I}$, suppose $M$ admits $i$-th reflection, then $R_{i}(M)$ admits $i$-th reflection. Furthermore, we have:
+(1) For each $i\in\mathbb{I}$, suppose $M$ admits $i$-th reflection, then $R_{i}(M)$ admits the $i$-th reflection. Furthermore, we have:
 $R_{i}^{2}(M)\cong M,$ (5.6)
@@
 $r:\mathbb{I}\times\mathcal{X}\rightarrow\mathcal{X},\ i\times[X]\mapsto[R_{i}(X)].$
-Furthermore, we assume each $\mathcal{B}(P_{i})$ is finite-dimensional for $P\in\mathcal{X}$ and $i\in\mathbb{I}$, then
+Then
 $\mathcal{G}(M)=(\mathbb{I},\mathcal{X},r,(A^{X})_{X\in\mathcal{X}}),$
@@
 Further, $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}$ is generated by $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)$ by Theorem 5.5.
-Then we have $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)+M_{i}$ generates $\mathcal{B}(M)$ and $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)+M_{i}^{*}$ generates $\Omega_{i}(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})})\#\mathcal{B}(M_{i}^{*})$, and
+Then $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)+M_{i}$ generates $\mathcal{B}(M)$ and $\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)+M_{i}^{*}$ generates $\Omega_{i}(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})})\#\mathcal{B}(M_{i}^{*})$, and
 $(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)+M_{i})^{n}$ $=\bigoplus_{k=0}^{n}\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}(1)^{n-k}\#M_{i}^{k}\quad\text{in }\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})}\#\mathcal{B}(M_{i}),$
@@
 Hence the theorem holds since $\mathcal{B}(R_{i}(M))\simeq\Omega_{i}(\mathcal{B}(M)^{\operatorname{co}\mathcal{B}(M_{i})})\#\mathcal{B}(M_{i}^{*})$ as Hopf algebras in $^{H}_{H}\mathcal{YD}$}. ∎
-6 An example of affine Nichols algebra
-In this section, we consider a rank three Nichols algebra $\mathcal{B}(M)$ of non-diagonal type in [Li and Liu, 2025b], which plays a central role in classification of finite-dimensional Nichols algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$}, where $G$ is a finite abelian group and $\Phi$ is a $3$-cocycle on $G$.
-It was proved to be infinite-dimensional via different ways in [Huang et al., 2026,Li and Liu, 2025a,Li and Liu, 2025b].
+6 Example of an affine Nichols algebra
+In this section, we consider a rank three Nichols algebra $\mathcal{B}(M)$ of non-diagonal type in [Li and Liu, 2025b], which plays a central role in the classification of finite-dimensional Nichols algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$}, where $G$ is a finite abelian group and $\Phi$ is a $3$-cocycle on $G$.
+It was proved to be infinite-dimensional in different ways in [Huang et al., 2026,Li and Liu, 2025a,Li and Liu, 2025b].
 Moreover, it gives rise to a semi-Cartan graph $\mathcal{G}(M)$.
@@
 By [Huang et al., 2026, Lemma 3.5], we may give a complete list of irreducible Yetter-Drinfeld modules over $^{G}_{G}\mathcal{YD}^{\Phi}$}.
-In particular, restricting to those irreducible Yetter–Drinfeld modules that generate finite-dimensional Nichols algebras, the isomorphism classes are given by the set
+In particular, when restricting to those irreducible Yetter-Drinfeld modules that generate finite-dimensional Nichols algebras, we see that the isomorphism classes are given by the set
 $S=\{[^{h}M]\mid\ ^{h}M\ \text{is irreducible},\ h\in\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}-\{1,h_{1}h_{2}h_{3}\}\}.$ (6.1)
@@
 For $1\leq i\leq 6$, let $M_{i}\in S$ be pairwise non-isomorphic simple modules such that $\operatorname{dim}(M_{i})=2$, and $\operatorname{deg}(M_{1})=h_{1}$, $\operatorname{deg}(M_{2})=h_{2}$, $\operatorname{deg}(M_{3})=h_{3}$, $\operatorname{deg}(M_{4})=h_{1}h_{2}$, $\operatorname{deg}(M_{5})=h_{1}h_{3}$, $\operatorname{deg}(M_{6})=h_{2}h_{3}$.
-One may refer to Li and Liu (2025b) for explicit expression of Yetter-Drinfeld structure, we omit here for simplicity.
-The Nichols algebra $\mathcal{B}(M)$ which we concern is generated by $M_{1}\oplus M_{2}\oplus M_{3}$. Furthermore, we proved that:
+One may refer to Li and Liu (2025b) for the explicit expression of the Yetter-Drinfeld structure, we omit here for simplicity.
+The Nichols algebra $\mathcal{B}(M)$ which we are concerned with is generated by $M_{1}\oplus M_{2}\oplus M_{3}$. Furthermore, we proved that:
 Theorem 6.1.
@@
 This agreement is expected because the semi-Cartan graph is standard, and thus $\Delta^{M,\operatorname{re}}$ coincides with the set of real roots for the affine Lie algebra $A_{2}^{(1)}$.
-Unfortunately, $\Delta^{M,\operatorname{re}}$ is a infinite set, by definition $\mathcal{G}(M)$ is not a finite semi-Cartan graph.
+Unfortunately, $\Delta^{M,\operatorname{re}}$ is an infinite set, by definition $\mathcal{G}(M)$ is not a finite semi-Cartan graph.
 Definition 6.3.
@@
 (2) A semi-Cartan graph $\mathcal{G}$ is called simply connected if for any two points $X,Y$ of $\mathcal{G}$ there is at most one morphism from $X$ to $Y$ in $\mathcal{W}(\mathcal{G})$.
-(3) For any $X\in\mathcal{X}$ and $i,j\in I$ let
+(3) For any $X\in\mathcal{X}$ and $i,j\in\mathbb{I}$ let
 $m_{ij}^{X}=|\Delta^{X\,\text{re}}\cap(\mathbb{N}_{0}\alpha_{i}+\mathbb{N}_{0}\alpha_{j})|.$
@@
   3. (3)
-    For any $X\in\mathcal{X}$ and $i\in I$, $s_{i}^{X}(R^{X})=R^{r_{i}(X)}$.
+    For any $X\in\mathcal{X}$ and $i\in\mathbb{I}$, $s_{i}^{X}(R^{X})=R^{r_{i}(X)}$.
   4. (4)
@@
 A root system over $\mathcal{G}$ is said to be reduced if for all $X\in\mathcal{X}$ and $\alpha\in R^{X}$ the roots $\alpha$ and $\alpha$ are the only rational multiples of $\alpha$ in $R^{X}$.
-A root system is said to be reduced if for all $X\in\mathcal{X}$, $R^{X}$ is finite.
+A root system is said to be finite if for all $X\in\mathcal{X}$, $R^{X}$ is finite.
 If $\mathcal{G}$ is a Cartan graph, the pair $(\mathcal{G},(\Delta^{X,\operatorname{re}})_{X\in\mathcal{X}})$ is a reduced root system over $\mathcal{G}$ by [Heckenberger and Schneider, [2020] ©2020, Example 10.4.4].
@@
 6.2 The Tits cone induced by $(\mathcal{G}(M),(\Delta^{X,\operatorname{re}})_{X\in\mathcal{X}})$ is a half plane
-In [Cuntz, 2018,Cuntz et al., 2019], the authors showed that most of Nichols algebras over Hopf algebras can define a Tits cone, via its root system.
+In [Cuntz, 2018,Cuntz et al., 2019], the authors showed that most Nichols algebras over Hopf algebras can define a Tits cone, via their root systems.
 In particular, they call a Nichols algebra affine if the corresponding Tits cone is a half plane.
@@
 Under this correspondence, equivalent Cartan graphs correspond to combinatorially equivalent Tits arrangements and vice versa, giving rise to a one-to-one correspondence between the respective equivalence classes.
-A crystallographic Tits arrangements consists of a pair $(\mathcal{A},T)$, where $\mathcal{A}$ is a set of linear hyperplanes in $\mathbb{R}^{r}$, and $T$ is a Tits cone in $\mathbb{R}^{r}$, satisfying some additional conditions.
-For explicit definition of crystallographic Tits arrangements, one may refer to [Cuntz et al., 2019, Definition 3.1].
-We give a brief recall of calculating Tits cone when given a connected simply connected Cartan graph $\mathcal{G}=\mathcal{G}(\mathbb{I},\mathcal{X},(r_{i})_{i\in I},(A^{X})_{X\in\mathcal{X}})$ with a root system $R=(\Delta^{X,\operatorname{re}})_{X\in\mathcal{X}}$.
+A crystallographic Tits arrangement consists of a pair $(\mathcal{A},T)$, where $\mathcal{A}$ is a set of linear hyperplanes in $\mathbb{R}^{r}$, and $T$ is a Tits cone in $\mathbb{R}^{r}$, satisfying some additional conditions.
+For the explicit definition of crystallographic Tits arrangements, one may refer to [Cuntz et al., 2019, Definition 3.1].
+We briefly recall the calculation of the Tits cone when given a connected simply connected Cartan graph $\mathcal{G}=\mathcal{G}(\mathbb{I},\mathcal{X},(r_{i})_{i\in I},(A^{X})_{X\in\mathcal{X}})$ with a root system $R=(\Delta^{X,\operatorname{re}})_{X\in\mathcal{X}}$.
 Fix an object $A\in\mathcal{X}$.
@@
 Theorem 6.7 guarantees that $(\mathcal{A},T)$ is a crystallographic Tits arrangement.
-The next proposition seems to be new, although those who are familiarly with affine Nichols algebras may regard it as a fact.
+The next proposition seems to be new, although those who are familiar with affine Nichols algebras may regard it as a known fact.
 Proposition 6.8.
@@
 Proof.
-We denote $r=|\mathbb{I}|$. Since $A$ is affine, it is automatically symmetrizable, then there exists a positive vector $v=(v_{1},v_{2},..,v_{r})\in\mathbb{Z}_{\geq 0}$ such that $v^{T}A=0$, that is
+We denote $r=|\mathbb{I}|$.
+Since $A$ is affine, it is automatically symmetrizable; thus there exists a positive vector $v=(v_{1},v_{2},..,v_{r})\in\mathbb{Z}^{r}_{\geq 0}$ such that $vw3e2qA=0$, that is
 $\sum_{i=1}^{r}v_{i}c_{ij}=0\ \text{for all}\ j.$
@@
 Moreover, each chamber $K^{X}$ is a r‑dimensional open simplicial cone.
-The intersection $K^{X}$ with $H_{1}$ is a non‑empty (r1)(r-1)-dimensional simplex.
+The intersection of $K^{X}$ with $H_{1}$ is a non‑empty (r1)(r-1)-dimensional simplex.
 In the theory of affine reflection groups, $\Delta:=K^{X}\cap H_{1}$ is known as a fundamental alcove.
-A classical result [Humphreys, 1990, Chapter 4] states that the affine Weyl group $W=W(A)$ acts simply transitively on the set of alcoves and that the closures of the alcoves tile the space $H_{1}$, i.e.
+A classical result in [Humphreys, 1990, Chapter 4] states that the affine Weyl group $W=W(A)$ acts simply transitively on the set of alcoves and that the closures of the alcoves tile the space $H_{1}$, i.e.
 $H_{1}=\bigcup_{w\in W}w(\overline{\Delta}).$
@@
 The point $x/\delta(x)$ must lie in the closure of some simplex.
-Consequently, we can find finitely many vertices $p_{1},\dots,p_{n}$ of these triangles, where each vertex belongs to some chamber $K^{X_{i}}\cap H_{1}$ and non‑negative coefficients $t_{1},\dots,t_{n}$ with $\sum_{i=1}^{n}t_{i}=1$ such that
+Consequently, we can find finitely many vertices $p_{1},\dots,p_{n}$ of these simplices, where each vertex belongs to some chamber $K^{X_{i}}\cap H_{1}$ and non‑negative coefficients $t_{1},\dots,t_{n}$ with $\sum_{i=1}^{n}t_{i}=1$ such that
 $x=\sum_{i=1}^{n}\delta(x)t_{i}p_{i}=\sum_{i=1}^{n}t_{i}(\delta(x)p_{i}).$
@@
 Now we obtain $T=\delta^{+}$.
-Thus by philosophy of Theorem 6.7, the Tits cone corresponding to $(\mathcal{G},(\Delta^{X,\operatorname{re}})_{X\in\mathcal{X}})$ is a half plane.
+Thus, by the philosophy of Theorem 6.7, the Tits cone corresponding to $(\mathcal{G},(\Delta^{X,\operatorname{re}})_{X\in\mathcal{X}})$ is a half plane.
 ∎
@@
 Remark 6.10.
-(1) This means we realize an affine Nichols algebra over a coquasi-Hopf algebra, which broadens the realizing approaches of affine Nichols algebras.
-(2) We hope to prove that a semi-Cartan graph induced by Nichols algebra over $^{H}_{H}\mathcal{YD}$} is actually a Cartan graph, which will endow them more beautiful properties.
+(1) This means we realize an affine Nichols algebra over a coquasi-Hopf algebra, which broadens the approaches to realizing affine Nichols algebras.
+(2) We hope to prove that the semi-Cartan graph induced by a Nichols algebra over $^{H}_{H}\mathcal{YD}$} is actually a Cartan graph, which would endow it with more beautiful properties.
 Acknowledgment This work is supported by the National Key R&D Program of China